### v2

# Robust Instance Optimal Phase-Only Compressed Sensing

Abstract
Phase-only compressed sensing (PO-CS) concerns the recovery of sparse signals from the phases of complex measurements. Recent results show that sparse signals in the standard sphere $\mathbb{S}^{n-1}$ can be exactly recovered from complex Gaussian phases by a linearization procedure, which recasts PO-CS as linear compressed sensing and then applies (quadratically constrained) basis pursuit to obtain $\mathbf{x}^{\sharp}$. This paper focuses on the instance optimality and robustness of $\mathbf{x}^{\sharp}$. First, we strengthen the nonuniform instance optimality of Jacques and Feuillen (2021) to a uniform one over the entire signal space. We show the existence of some universal constant $C$ such that $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq Cs^{-1/2}\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})$ holds for all $\mathbf{x}$ in the unit Euclidean sphere, where $\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})$ is the $\ell_{1}$ distance of $\mathbf{x}$ to its closest $s$-sparse signal. This is achieved by showing the new sensing matrices corresponding to all approximately sparse signals simultaneously satisfy RIP. Second, we investigate the estimator’s robustness to noise and corruption. We show that dense noise with entries bounded by some small $\tau_{0}$, appearing either prior or posterior to retaining the phases, increments $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ by $O(\tau_{0})$. This is near-optimal (up to log factors) for any algorithm. On the other hand, adversarial corruption, which changes an arbitrary $\zeta_{0}$-fraction of the measurements to any phase-only values, increments $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ by $O(\sqrt{\zeta_{0}\log(1/\zeta_{0})})$. We demonstrate the tightness of this result via a partial analysis under suboptimal noise parameter and numerical evidence, while showing that the impact of sparse corruption can be eliminated by proposing an extended linearization approach that can exactly recover $\mathbf{x}$ from the corrupted phases. The developments are then combined to yield a robust instance optimal guarantee that resembles the standard one in linear compressed sensing.
Keywords: Compressed sensing, Nonlinear observations, Instance optimality, Robustness, Covering
1 Introduction
Compressed sensing has proven to be an effective method in acquiring and reconstructing high-dimensional signals [20, 17, 6, 5, 13]. Mathematically, the goal of linear compressed sensing is to reconstruct sparse signals $\mathbf{x}$ from a set of measurements $\mathbf{y}=\mathbf{A}\mathbf{x}+\bm{\epsilon}$, under the sensing matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ and noise vector $\bm{\epsilon}\in\mathbb{R}^{m}$. Restricted isometry property (RIP) lies at the center of linear compressed sensing theory, whose major finding is a set of efficient algorithms achieving instance optimality under RIP sensing matrices [20, 47, 51, 2, 14, 5]. Here, the instance optimality describes the capacity of an algorithm to achieve estimation error proportional to the signal’s distance to the cone of $s$-sparse vectors $\Sigma^{n}_{s}$. In the noiseless case, this translates into exact reconstruction of sparse signals and accurate estimate of approximately sparse signals. As an example, if $\mathbf{A}$ satisfies RIP over the cone of sparse vectors and $\varepsilon\geq\|\bm{\epsilon}\|_{2}$, then basis pursuit $\Delta(\mathbf{A};\mathbf{y};\varepsilon)$
$\hat{\mathbf{x}}=\mathrm{arg}\min\|\mathbf{u}\|_{1},\quad\text{subject to }\|\mathbf{A}\mathbf{u}-\mathbf{y}\|_{2}\leq\varepsilon$ (1.1)
achieves
$\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}\leq C_{1}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}+C_{2}\varepsilon,\qquad\forall\mathbf{x}\in\mathbb{R}^{n}$ (1.2)
for some absolute constants $C_{1},C_{2}$, where $\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s}):=\min_{\mathbf{u}\in\Sigma^{n}_{s}}\|\mathbf{x}-\mathbf{u}\|_{1}$ denotes the $\ell_{1}$ distance of $\mathbf{x}$ to $\Sigma^{n}_{s}.$ More details are given in Section 2.
The focus of the present paper is on the nonlinear compressed sensing model of phase-only compressed sensing (PO-CS), which concerns the reconstruction of sparse signals in the standard sphere $\mathbb{S}^{n-1}=\{\mathbf{x}\in\mathbb{R}^{n}:\|\mathbf{x}\|_{2}=1\}$ from
$\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})=[\operatorname{sign}(\bm{\Phi}_{1}^{*}\mathbf{x}),\cdots,\operatorname{sign}(\bm{\Phi}_{m}^{*}\mathbf{x})]^{\top}$ (1.3)
under a complex sensing matrix $\bm{\Phi}=[\bm{\Phi}_{1},\cdots,\bm{\Phi}_{m}]^{*}\in\mathbb{C}^{m\times n}$, with the phase function being given by $\operatorname{sign}(c)=\frac{c}{|c|}$ for $c\in\mathbb{C}\setminus\{0\}$ and $\operatorname{sign}(0)=1$ by convention. We assume that we observe $\breve{\mathbf{z}}\in\mathbb{C}^{m}$ that equaling $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ in the noiseless case or a perturbed version of $\mathbf{z}$ in the noisy case. We note that phase-only reconstruction of unstructured signals has been well studied [26, 38, 9, 25, 18, 32, 48, 33].
PO-CS was initially considered as a natural extension of 1-bit compressed sensing [^1] This concerns the recovery of $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\mathbf{y}=\operatorname{sign}(\mathbf{A}\mathbf{x})$ with real sensing matrix $\mathbf{A}\in\mathbb{R}^{m\times n}.$ [3, 4] and was recently revisited in [19, 10, 27, 8]. Exact reconstruction was observed experimentally in [4] and theoretically proved very recently by Jacques and Feuillen [27] who proposed to recast PO-CS as a linear compressed sensing problem. We proceed to introduce this linearization approach under the noiseless setting $\breve{\mathbf{z}}=\mathbf{z}$. Since $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ implies that the entries of $\operatorname{diag}(\breve{\mathbf{z}}^{*})\bm{\Phi}\mathbf{x}$ are non-negative real numbers, the phases give the linear measurements
$\frac{1}{\sqrt{m}}\Im(\operatorname{diag}(\breve{\mathbf{z}}^{*})\bm{\Phi})\mathbf{x}=0,$ (1.4)
where we use $\Re$ and $\Im$ to denote the real part and imaginary part, $\breve{\mathbf{z}}^{*}$ to denote the conjugate transpose of $\breve{\mathbf{z}}$, and $\operatorname{diag}(\mathbf{a})$ to denote the diagonal matrix with diagonal $\mathbf{a}$. Since (1.4) does not contain any information on $\|\mathbf{x}\|_{2}$, we note that $\Re(\breve{\mathbf{z}}^{*}\bm{\Phi}\mathbf{x})=\|\bm{\Phi}\mathbf{x}\|_{1}$ and further enforce an additional measurement
$\frac{1}{\kappa m}\Re(\breve{\mathbf{z}}^{*}\bm{\Phi})\mathbf{x}=1$ (1.5)
with $\kappa=\sqrt{\frac{\pi}{2}}$ to specify the norm of the desired signal. (The value of $\kappa$ here is non-essential but chosen to facilitate subsequent analysis.) Combining (1.4) and (1.5), we arrive at the linear compressed sensing problem
[…]
In the noisy case with $\breve{\mathbf{z}}\neq\mathbf{z}$, the linear equations in (1.6) become inexact and we instead encounter a noisy linear compressed sensing problem.
For convenience, for any $\mathbf{w}\in\mathbb{C}^{m}$, we introduce the notation
$\mathbf{A}_{\mathbf{w}}:=\begin{bmatrix}\frac{1}{\kappa m}\Re(\mathbf{w}^{*}\bm{\Phi})\\ \frac{1}{\sqrt{m}}\Im(\operatorname{diag}(\mathbf{w}^{*})\bm{\Phi})\end{bmatrix}.$ (1.7)
We will refer to $\mathbf{A}_{\breve{\mathbf{z}}}$ in (1.6) as the new sensing matrix in order to distinguish it with the original complex sensing matrix $\bm{\Phi}$. For a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and $\bm{\Phi}$ with i.i.d. $\mathcal{N}(0,1)+\mathcal{N}(0,1)\mathbf{i}$ entries, it was proved [27] that, with small enough $\|\breve{\mathbf{z}}-\mathbf{z}\|_{\infty}=\max_{i\in[m]}|\breve{z}_{i}-z_{i}|$, the matrix $\mathbf{A}_{\breve{\mathbf{z}}}\in\mathbb{R}^{(m+1)\times n}$ with $m=O(s\log(\frac{en}{s}))$ satisfies RIP over the cone of $2s$-sparse vectors with high probability. Hence, one may solve (1.6) by instance optimal algorithms from linear compressed sensing theory to exactly recover sparse signals and accurately recover approximately sparse signals. However, this guarantee from [27] is a nonuniform instance optimality result that only works for a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$. In fact, the new sensing matrix $\mathbf{A}_{\breve{\mathbf{z}}}$ depends on $\mathbf{x}$ through $\breve{\mathbf{z}}$, so proving the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$ for a fixed $\mathbf{x}$ only implies the nonuniform recovery of this specific $\mathbf{x}$.
In [10], the authors used a covering argument to show that the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}:=\Sigma^{n}_{s}\cap\mathbb{S}^{n-1}\}$ simultaneously obey RIP under a near-optimal number of measurements. This leads to a uniform exact reconstruction guarantee over $\Sigma^{n,*}_{s}$, but provides no guarantee for $\mathbf{x}\notin\Sigma^{n,*}_{s}$. Thus, their guarantee is not instance optimal.
The first contribution of this work is to show that the above linearization approach indeed achieves uniform instance optimality, which is stronger than the nonuniform result in [27]. For concreteness, we focus on quadratically constrained basis pursuit, that is to solve $\hat{\mathbf{x}}$ from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\varepsilon)$
$\min~{}\|\mathbf{u}\|_{1},\quad\text{subject to }~{}~{}\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}-\mathbf{e}_{1}\|_{2}\leq\varepsilon$ (1.8)
for some suitably chosen $\varepsilon$, and then use $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ as an estimator for $\mathbf{x}$. In the noiseless case, we show that when using a complex Gaussian matrix $\bm{\Phi}$ with $O(s\log(\frac{en}{s}))$ rows, with high probability $\mathbf{x}^{\sharp}$ satisfies
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{10\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}},\qquad\forall\mathbf{x}\in\mathbb{S}^{n-1}.$
The main ingredient is to show the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$, where $\mathbb{B}_{1}^{n}(\sqrt{2s})=\{\mathbf{u}\in\mathbb{R}^{n}:\|\mathbf{u}\|_{1}\leq\sqrt{2s}\}$ is the scaled-$\ell_{1}$ ball, simultaneously satisfy RIP. Note that [10] proved that the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy RIP through a covering argument, yet their arguments are not sufficient to prove the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$. The main issue is that $\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}$ is essentially larger than $\Sigma^{n,*}_{s}$ in terms of metric entropy (or covering number) under an $o(1)$ covering radius. To that end, we utilize a finer treatment to the perturbation of the complex phases to avoid a heavy-tailed random process. See Appendix A.1 for a summary. More generally, we establish the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ over some cone $\mathcal{U}$ for an arbitrary set $\mathcal{K}\subset\mathbb{S}^{n-1}$. Here is an informal version.
Theorem (Informal).
Given a cone $\mathcal{U}$ in $\mathbb{R}^{n}$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if $m\geq C_{1}\big{(}\omega^{2}(\mathcal{U}\cap\mathbb{S}^{n-1})+\omega^{2}(\mathcal{K})\big{)}$ with sufficiently large $C_{1}$, then with high probability on the complex Gaussian $\bm{\Phi}$, the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ satisfy RIP over $\mathcal{U}$ with small enough distortion.
Our second contribution is to understand the robustness of $\mathbf{x}^{\sharp}$ to different patterns of noise and corruption. [^2] As a convention, noise refers to perturbation with small magnitude, while corruption can change a measurement to an arbitrary phase-only value. Prior works [27, 10] only considered small dense noise appearing after applying the sign function (termed as post-sign noise), formulated as $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\tau}$ where $\bm{\tau}\in\mathbb{C}_{m}$ satisfies $\|\bm{\tau}\|_{\infty}=\max_{i}|\tau_{i}|\leq\tau_{0}$. Under small enough $\tau_{0}$, they showed a stability result that such $\bm{\tau}$ increments the estimation error of $\mathbf{x}^{\sharp}$ by $O(\tau_{0})$. However, many questions remain unaddressed: Is the $O(\tau_{0})$ bound for post-sign noise tight? Is $\mathbf{x}^{\sharp}$ robust to small dense noise appearing before $\operatorname{sign}(\cdot)$, which we call pre-sign noise? Is the estimator robust to malicious sparse phase corruption? If so, how do the pre-sign noise and the sparse corruption increment $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$? Are these increments tight or suboptimal, in some sense?
Our results provide answers to all of these questions. Let us consider the nonuniform recovery of a fixed sparse signal. We show that small dense noise $\bm{\tau}$ considered in [10, 27], even when appearing before taking the phases (i.e., pre-sign noise), increments $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ by $O(\tau_{0})$. Moreover, the $O(\tau_{0})$ bound achieved by $\mathbf{x}^{\sharp}$ for pre-sign/post-sign dense noise is nearly tight over all algorithms. We also investigate the impact of sparse corruption which adversarially moves $\zeta_{0}m$ measurements to arbitrary phase-only values. This increments $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ by $O(\sqrt{\zeta_{0}\log(1/\zeta_{0})})$ if $\zeta_{0}$ is small enough. We expect that $\tilde{O}(\sqrt{\zeta_{0}})$ is tight for the specific estimator $\mathbf{x}^{\sharp}$, which we support by providing a partial analysis and numerical evidence. However, for general estimators such dependence is suboptimal and can be improved to “zero,” as in this regime we can still exactly recover $\mathbf{x}$, akin to existing studies of corrupted linear systems and corrupted sensing [21, 7, 24]. In particular, we propose to reformulate PO-CS under sparse corruption to a noiseless linear compressed sensing problem with an extended new sensing matrix. This matrix is then shown to satisfy RIP, thus implying exact reconstruction. See Table 1 for a summary of these results.
Assumption Error Bound Tightness w.r.t. the estimator $\mathbf{x}^{\sharp}$ Tightness w.r.t. all estimators Simulation
Bounded Dense Noise $\bm{\tau}$ $\|\bm{\tau}\|_{\infty}\leq\tau_{0}$ $O(\tau_{0})$ Thms. 3.3 & 3.4 Yes, up to log Thm. 3.5 Yes, up to log Thm. 3.5 Figs. 12
Sparse Corruption $\bm{\zeta}$ $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ $O\Big{(}\sqrt{\zeta_{0}\log(1/\zeta_{0})}\Big{)}$ Thm. 3.6 Partially Yes Prop. 3.1 No, 00 is optimal Thm. 3.8 Fig. 3
Table 1: A summary of our results on robustness. The estimator $\mathbf{x}^{\sharp}$ is defined in (1.8). By tightness, we mean the question of whether the scaling of the error bound is the best possible (for estimator $x^{\sharp}$ in the fourth column, and general estimators in the fifth column).
Combining the two developments, we obtain the following result that closely resembles the instance optimal guarantee in linear compressed sensing (1.2); see Section 4.
Theorem (Informal).
Under noisy observations $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)}+\bm{\zeta}_{(1)})+\bm{\tau}_{(2)}+\bm{\zeta}_{(2)}$ with $\|\bm{\tau}_{(j)}\|_{\infty}\leq\tau_{0}$ and $\|\bm{\zeta}_{j}\|_{0}\leq\zeta_{0}m$ for $j=1,2$ and some small enough $\tau_{0},\zeta_{0}$, and $\|\bm{\zeta}_{(2)}\|_{\infty}\leq 2$, consider $\mathbf{x}^{\sharp}$ with properly tuned $\varepsilon$. Then with high probability over the complex Gaussian $\bm{\Phi}$, we have
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq C_{1}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}+C_{2}\tau_{0}+C_{3}\sqrt{\zeta_{0}\log(1/\zeta_{0})}+C_{4}\sqrt{\frac{s\log(\frac{en}{s})}{m}},\qquad\forall\mathbf{x}\in\mathbb{S}^{n-1}.$ (1.9)
To our knowledge, this type of result is novel in nonlinear sensing problems, for which most existing guarantees provide the same error rate to all signals of interest and hence are not instance optimal (e.g., [50, 11, 40, 41, 44, 43, 23]). [^3] For instance, the best known $\ell_{2}$ error rate for 1-bit compressed sensing of the approximately sparse signals in $\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{S}^{n-1}$ is $\tilde{O}((s/m)^{1/3})$, while the optimal rate for sparse signals in $\Sigma^{n,*}_{s}$ is $\tilde{O}(s/m)$. One exception is the instance optimality in sparse phase retrieval achieved by an intractable algorithm [22]. We also note a generic discussion [29] which does not lead to efficient algorithm or address specific model.
Organization. The remainder of this paper is arranged as follows. In Section 2 we give the preliminaries. We present our main results on instance optimality and (nonuniform) robustness in Section 3, along with a few simulation results. In Section 4 we combine the instance optimality and nonuniform robustness in the previous section to establish the above (1.9). We give concluding remarks in Section 5 to close the paper. Some lengthy and secondary proofs are relegated to the appendices. The proof of Theorem 3.1 is rather technical and modified from the arguments in [10] (with crucial improvements); it is hence presented in Appendix A. The proofs of two side results are postponed to Appendix B.
2 Preliminaries
We start with some notational conventions. We denote matrices and vectors by boldface letters, and scalars by regular letters. $|\mathcal{S}|$ denotes the cardinality of a finite set $\mathcal{S}$. We use $\log(\cdot)$ to denote the natural logarithm to the base of the mathematical constant $e$. The standard Euclidean sphere, the $\ell_{2}$-ball and the $\ell_{1}$-ball in $\mathbb{R}^{n}$ are denoted by $\mathbb{S}^{n-1}$, $\mathbb{B}_{2}^{n}$ and $\mathbb{B}_{1}^{n}$, respectively. Given $\mathcal{K},\mathcal{K}^{\prime}\subset\mathbb{R}^{n}$ and some $\lambda\in\mathbb{R}$, we let $\mathcal{K}+\lambda\mathcal{K}^{\prime}:=\{\mathbf{u}+\lambda\mathbf{v}:\mathbf{u}\in\mathcal{K},\mathbf{v}\in\mathcal{K}^{\prime}\}$, $\mathcal{K}_{(\lambda)}=(\mathcal{K}-\mathcal{K})\cap(\lambda\mathbb{B}_{2}^{n})$ and $\mathcal{K}^{\rm(N)}=\mathcal{K}\cap\mathbb{S}^{n-1}$. We write $\mathbb{B}_{2}^{n}(\mathbf{u};r):=\mathbf{u}+r\mathbb{B}_{2}^{n}$, $\mathbb{B}_{2}^{n}(r):=r\mathbb{B}_{2}^{n}$ and $\mathbb{B}_{1}^{n}(r):=r\mathbb{B}_{1}^{n}$. We also define $\operatorname{rad}(\mathcal{K})=\sup_{\mathbf{u}\in\mathcal{K}}\|\mathbf{u}\|_{2}$. Recall that $\Sigma^{n,*}_{s}:=(\Sigma^{n}_{s})^{\rm(N)}$ is the set of all $s$-sparse signals in $\mathbb{S}^{n-1}$.
We refer to complex numbers with absolute value $1$ as the phase-only values. For a vector $\mathbf{u}=[u_{i}]\in\mathbb{C}^{n}$, we work with the $\ell_{p}$-norm $\|\mathbf{u}\|_{p}=(\sum_{i}|u_{i}|^{p})^{1/p}$ ($p\geq 1$), max norm $\|\mathbf{u}\|_{\infty}=\max_{i}|u_{i}|$, and zero “norm” $\|\mathbf{u}\|_{0}$ that counts the number of non-zero entries. Further given $\mathbf{v}=[v_{i}]\in\mathbb{C}^{n}$, we work with the inner product $\langle\mathbf{u},\mathbf{v}\rangle=\mathbf{u}^{*}\mathbf{v}=\sum_{i=1}^{n}\overline{u_{i}}v_{i}$ and the Hadamard product $\mathbf{u}\odot\mathbf{v}=(u_{1}v_{1},u_{2}v_{2},\cdots,u_{n}v_{n})^{\top}$. For a complex matrix $\mathbf{A}=\mathbf{B}+\mathbf{C}\mathbf{i}$ with $\mathbf{i}$ reserved for the complex unit, we will use $\mathbf{A}^{\Re}$ or $\Re(\mathbf{A})$ to denote its real part $\mathbf{B}$, and $\mathbf{A}^{\Im}$ or $\Im(\mathbf{A})$ to denote its imaginary part $\mathbf{C}$. For a random variable $X$ we define the sub-Gaussian norm as $\|X\|_{\psi_{2}}=\inf\{t>0:\mathbbm{E}\exp(X^{2}/t^{2})<2\}$ and the sub-exponential norm as $\|X\|_{\psi_{1}}=\inf\{t>0:\mathbbm{E}\exp(|X|/t)<2\}$. For independent zero-mean sub-Gaussian variables $\{X_{i}\}_{i=1}^{N}$, there exists absolute constant $C$ such that
$\left\|\sum_{i=1}^{N}X_{i}\right\|_{\psi_{2}}^{2}\leq C\sum_{i=1}^{N}\|X_{i}\|_{\psi_{2}}^{2}.$ (2.1)
The sub-Gaussian norm of a random vector $\mathbf{X}\in\mathbb{R}^{n}$ is defined as $\|\mathbf{X}\|_{\psi_{2}}=\sup_{\mathbf{v}\in\mathbb{S}^{n-1}}\|\mathbf{v}^{\top}\mathbf{X}\|_{\psi_{2}}$. We refer readers to [49, Sec. 2] for details on these definitions and properties.
We use $\{C,C_{1},C_{2},\cdots\}$ and $\{c,c_{1},c_{2},\cdots\}$ to denote absolute constants whose values may vary from line to line. For two positive quantities $I_{1}$ and $I_{2}$, we write $I_{1}=O(I_{2})$ if $I_{1}\leq CI_{2}$ holds for some $C$, and write $I_{1}=\Omega(I_{2})$ if $I_{1}\geq cI_{2}$ for some $c>0$. We write $I_{1}=\Theta(I_{2})$ if $I_{1}=O(I_{2})$ and $I_{1}=\Omega(I_{2})$ simultaneously hold. We also use $\tilde{O}(\cdot),\tilde{\Omega}(\cdot),\tilde{\Theta}(\cdot)$ as the less precise versions of these that hide log factors. We use $o(1)$ to generically denote quantity that tends to zero when $m,n,s\to\infty$.
We say $I_{1}$ is small enough (or sufficiently small) if $I_{1}\leq c_{1}$ for some suitably small constant $c_{1}$. Conversely, it is large enough (or sufficiently large) if $I_{1}\geq C_{1}$ for some suitably large constant $C_{1}$. We say $I_{1}$ is bounded away from 00 if $I_{1}\geq c_{1}$ for some $c_{1}>0$.
Linear Compressed Sensing: Let $\mathcal{U}$ be a cone in $\mathbb{R}^{n}$, we say $\mathbf{A}$ satisfies RIP over $\mathcal{U}$ with distortion $\delta>0$, denoted by $\mathbf{A}\sim\operatorname{RIP}(\mathcal{U},\delta)$, if
$(1-\delta)\|\mathbf{u}\|_{2}^{2}\leq\|\mathbf{A}\mathbf{u}\|_{2}^{2}\leq(1+\delta)\|\mathbf{u}\|_{2}^{2},\qquad\forall\mathbf{u}\in\mathcal{U}.$
By homogeneity, this is equivalent to
$\sqrt{1-\delta}\leq\|\mathbf{A}\mathbf{u}\|_{2}\leq\sqrt{1+\delta},\qquad\forall\mathbf{u}\in\mathcal{U}^{\rm(N)}=\mathcal{U}\cap\mathbb{S}^{n-1}.$
To be specific, we will focus on sparse recovery and the corresponding program of $\ell_{1}$-norm minimization (1.1). Therefore, we typically utilize the RIP over $\Sigma^{n}_{ts}$ for certain $t>0$ to imply the instance optimality. We shall work with RIP over $\Sigma^{n}_{2s}$ and utilize the following. (Note that RIP$(\Sigma^{n}_{2s},\delta)$ with $\delta<\frac{\sqrt{2}}{2}$ works, and we set $\delta=\frac{1}{3}$ just for concreteness.)
Proposition 2.1 (see Thm. 2.1 in [5]).
Consider $\hat{\mathbf{x}}$ obtained by solving $\Delta(\mathbf{A};\mathbf{y};\varepsilon)$ in (1.1). If $\mathbf{A}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$ and $\|\mathbf{y}-\mathbf{A}\mathbf{x}\|_{2}\leq\varepsilon$, then we have
$\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}\leq 7\varepsilon+5\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}},\qquad\forall\mathbf{x}\in\mathbb{R}^{n}.$ (2.2)
Guarantees of this type are standard in linear compressed sensing theory (e.g., the block sparsity example below; see also [46, 20]). We note three important features of (2.2): instance optimality characterized by $O(s^{-1/2}\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s}))$, robustness captured by $O(\varepsilon)$, and uniformity over the entire signal space $\mathbb{R}^{n}$. Indeed, the central goal of this work is to prove an analog for an efficient PO-CS algorithm.
[…]
$\hat{\mathbf{x}}=(\hat{\mathbf{x}}_{1}^{\top},\hat{\mathbf{x}}_{2}^{\top})^{\top}=\mathrm{arg}\min_{\mathbf{u}=(\mathbf{u}_{1}^{\top},\mathbf{u}_{2}^{\top})^{\top}}~{}\frac{\|\mathbf{u}_{1}\|_{1}}{\sqrt{s_{1}}}+\frac{\|\mathbf{u}_{2}\|_{1}}{\sqrt{s_{2}}},\quad\text{subject to }\|\mathbf{A}\mathbf{u}-\mathbf{y}\|_{2}\leq\varepsilon.$ (2.3)
This weighted $\ell_{1}$ norm can better promote the above block sparsity, which is slightly more structured than the ordinary sparsity $\Sigma^{n}_{s_{1}+s_{2}}$. Similarly to Proposition 2.1, we have the following. (Note that $\operatorname{RIP}(\Sigma^{n_{1}}_{2s_{1}}\times\Sigma^{n_{2}}_{2s_{2}},\delta)$ with $\delta<\frac{1}{2}$ works, and we set $\delta=\frac{1}{3}$ just for concreteness.)
Proposition 2.2 (Thms. 4.3 & 4.6 in [46]).
Consider recovering $\mathbf{x}=(\mathbf{x}_{1}^{\top},\mathbf{x}_{2}^{\top})^{\top}$ by solving $\hat{\mathbf{x}}$ from (2.3). If $\mathbf{A}\sim\operatorname{RIP}(\Sigma^{n}_{2s_{1}}\times\Sigma^{n}_{2s_{2}},\frac{1}{3})$ and $\|\mathbf{y}-\mathbf{A}\mathbf{x}\|_{2}\leq\varepsilon$, then for some absolute constants $C_{1},C_{2}$, we have
$\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}\leq C_{1}\varepsilon+C_{2}\left(\frac{\sigma_{\ell_{1}}(\mathbf{x}_{1},\Sigma^{n_{1}}_{s_{1}})}{\sqrt{s_{1}}}+\frac{\sigma_{\ell_{1}}(\mathbf{x}_{2},\Sigma^{n_{2}}_{s_{2}})}{\sqrt{s_{2}}}\right),\qquad\forall\mathbf{x}\in\mathbb{R}^{n}.$
Gaussian Width & Metric Entropy: We need to work with two natural quantities that characterize the complexity of a set $\mathcal{K}$. The first one is the Gaussian width $\omega(\mathcal{K}):=\mathbbm{E}\sup_{\mathbf{u}\in\mathcal{K}}~{}\mathbf{g}^{\top}\mathbf{u}$, where $\mathbf{g}\sim\mathcal{N}(0,\mathbf{I}_{n})$. The second one is the metric entropy $\mathscr{H}(\mathcal{K},r)=\log\mathscr{N}(\mathcal{K},r)$ where $\mathscr{N}(\mathcal{K},r)$ denotes the covering number of $\mathcal{K}$ under radius $r$, defined as the minimal number of radius-$r$ $\ell_{2}$-balls needed to cover $\mathcal{K}$. Metric entropy can be bounded in terms of the Gaussian width via Sudakov’s inequality [49, Coro. 7.4.3],
$\mathscr{H}(\mathcal{K},r)\leq\frac{C\cdot\omega^{2}(\mathcal{K})}{r^{2}}$ (2.4)
for some absolute constant $C$. We also have Dudley’s inequality [49, Sec. 8.1] for the converse purpose. A notable difference is that the Gaussian width remains invariant after taking the convex hull, while the metric entropy under $o(1)$ covering radius could change significantly. For instance, the set $\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{B}_{2}^{n}$ (whose elements are known as the approximately $s$-sparse signals in $\mathbb{B}_{2}^{n}$) can be essentially viewed as the convex hull of $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}$ [41, Lem. 3.1]. Their Gaussian widths are of the same order [40, Sec. 2],
$c_{1}\sqrt{s\log\big{(}\frac{en}{s}\big{)}}\leq\omega(\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n})\leq\omega(\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{B}_{2}^{n})\leq C_{2}\sqrt{s\log\big{(}\frac{en}{s}\big{)}}$ (2.5)
for some absolute constants $c_{1},C_{2}$. However, while we have
$\mathscr{H}(\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n},r)\leq C_{1}s\log\Big{(}\frac{en}{rs}\Big{)},$ (2.6)
after convexification we only have
$\mathscr{H}(\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{B}_{2}^{n},r)\leq C_{2}r^{-2}s\log\Big{(}\frac{en}{s}\Big{)}.$ (2.7)
The dependence on $r$ in (2.7) is tight in some regime; see [41, Sec. 3] and [44, Sec. 4.3.3]. In particular, the cardinality of an $r$-net for $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}$ logarithmically increases with $r^{-1}$, while that of $\sqrt{s}\mathbb{B}_{1}^{n}$ increases quadratically with $r^{-1}$.
Next, we introduce some useful sub-Gaussian concentration bounds that capture the Gaussian width of the relevant set. The following has proven highly effective in dealing with sparse corruption and yielding uniformity [11, 15, 28, 16], and we will rely on it to achieve similar goals.
Lemma 2.1 (e.g., Thm. 2.10 in [15]).
Let $\mathbf{a}_{1},...,\mathbf{a}_{m}$ be independent isotropic random vectors with $\max_{i}\|\mathbf{a}_{i}\|_{\psi_{2}}\leq L$, and consider some given $\mathcal{T}\subset\mathbbm{R}^{n}$. If $1\leq k\leq m$, then the event
$\sup_{\mathbf{u}\in\mathcal{T}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq k\end{subarray}}\left(\frac{1}{k}\sum_{i\in I}|\langle\mathbf{a}_{i},\mathbf{u}\rangle|^{2}\right)^{1/2}\leq C_{1}\left(\frac{\omega(\mathcal{T})}{\sqrt{k}}+\operatorname{rad}(\mathcal{T})\sqrt{\log\big{(}\frac{em}{k}\big{)}}\right)$ (2.8)
holds with probability at least $1-2\exp(-C_{2}k\log(\frac{em}{k}))$, where $C_{1}$ and $C_{2}$ are absolute constants only depending on $L$.
The following upper bound is a simple consequence of the matrix deviation inequality [49, Sec. 9.1] and will be of recurring use: if $\mathbf{A}$ has i.i.d. $\mathcal{N}(0,1)$ entries and $\mathcal{T}\subset\mathbb{R}^{n}$, then for any $t\geq 0$,
$\mathbbm{P}\left(\sup_{\mathbf{u}\in\mathcal{T}}\frac{\|\mathbf{A}\mathbf{u}\|_{2}}{\sqrt{m}}\leq\operatorname{rad}(\mathcal{T})+\frac{C_{1}\omega(\mathcal{T})+C_{2}t\cdot\operatorname{rad}(\mathcal{T})}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2}).$ (2.9)
A simple consequence of (2.9) is that, for $\bm{\Phi}$ with i.i.d. $\mathcal{N}(0,1)+\mathcal{N}(0,1)\mathbf{i}$ entries and some $\mathcal{T}\subset\mathbb{S}^{n-1}$, if $m=\Omega(\omega^{2}(\mathcal{T}))$, then with probability at least $1-4\exp(-c_{1}m)$ we have
$\sup_{\mathbf{u}\in\mathcal{T}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}\leq\sup_{\mathbf{u}\in\mathcal{T}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\mathcal{T}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\leq C_{2}.$ (2.10)
Perturbation of Complex Phase: Under the convention $\frac{x}{0}=\infty$ for any $x\geq 0$, it holds for any $a,b\in\mathbb{C}$ that (e.g., [10, Lem. 8])
$|\operatorname{sign}(a)-\operatorname{sign}(b)|\leq\min\left\{\frac{2|a-b|}{\max\{|a|,|b|\}},2\right\}.$ (2.11)
Note that $a\in\mathbb{C}$ can be identified with $(\Re(a),\Im(a))^{\top}\in\mathbb{R}^{2}$, and we can indeed generalize (2.11) to any $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$,
$\left\|\frac{\mathbf{a}}{\|\mathbf{a}\|_{2}}-\frac{\mathbf{b}}{\|\mathbf{b}\|_{2}}\right\|_{2}\leq\min\left\{\frac{2\|\mathbf{a}-\mathbf{b}\|_{2}}{\max\{\|\mathbf{a}\|_{2},\|\mathbf{b}\|_{2}\}},2\right\},$ (2.12)
by a proof identical to [10, Lem. 8]. In general, $|\operatorname{sign}(b+\delta)-\operatorname{sign}(b)|$ is harder to control under smaller $|b|$. This inspires us to introduce the index set
$\mathcal{J}_{\mathbf{x},\eta}=\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\}$
for some $\mathbf{x}\in\mathbb{S}^{n-1}$ and $\eta>0$. Intuitively, the measurements in $\mathcal{J}_{\mathbf{x},\eta}$ are possibly problematic in PO-CS in terms of the sensitivity to pre-sign perturbation.
3 Main Results
We present our results for $\mathbf{x}^{\sharp}$ obtained by normalizing $\hat{\mathbf{x}}=\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\epsilon)$ in (1.8). Throughout the paper, we assume complex Gaussian $\bm{\Phi}$ with i.i.d. $\mathcal{N}(0,1)+\mathcal{N}(0,1)\mathbf{i}$ entries (whose real part and imaginary part are independent) without always explicitly stating it. We will first study the performance of $\hat{\mathbf{x}}$ and then transfer this to $\mathbf{x}^{\sharp}$ via (2.12); thus, it is useful to identify the ground truth that $\hat{\mathbf{x}}$ approximates. This is a scaled version of $\mathbf{x}$ given by [10, 27]
$\mathbf{x}^{\star}:=\frac{\kappa m\cdot\mathbf{x}}{\|\bm{\Phi}\mathbf{x}\|_{1}},$ (3.1)
as it is easy to check $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$. (This means that $\mathbf{x}^{\star}$ is the point that satisfies the linear measurements in (1.8) in a noiseless case.) With $\kappa=\sqrt{\frac{\pi}{2}}$, $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ sharply concentrates about $1$ due to sub-Gaussian tail bounds, as we have $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=O(\frac{1}{\sqrt{m}})$ by (2.1). [^4] By (2.1), we have $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=\|\frac{1}{m}\sum_{i=1}^{m}(\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1)\|_{\psi_{2}}\leq\frac{C}{m}(\sum_{i=1}^{m}\|\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1\|_{\psi_{2}}^{2})^{1/2}\leq\frac{C_{1}}{\sqrt{m}}$. As a consequence, we have the sub-Gaussian tail bound $\mathbbm{P}(|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|\geq t)\leq 2\exp(-c_{2}mt)$ for any $t\geq 0$, hence $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ is small enough with probability at least $1-2\exp(-c_{3}m)$. We will use this observation in subsequent analysis. Hence, $\mathbf{x}^{\star}$ is in general very close to $\mathbf{x}$.
3.1 Instance Optimality
Our first result concerns the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ for arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$ over a general cone $\mathcal{U}$.
Theorem 3.1 (RIP of a set of $\mathbf{A}_{\mathbf{z}}$).
Given a set $\mathcal{K}$ contained in $\mathbb{S}^{n-1}$, a cone $\mathcal{U}\subset\mathbb{R}^{n}$, and any small enough $\eta\in(0,1)$, we let $r=\eta^{2}\log^{1/2}(\eta^{-1})$ and consider drawing a complex Gaussian $\bm{\Phi}\in\mathbb{C}^{m\times n}$. If
$m\geq C_{1}\left(\frac{\omega^{2}(\mathcal{U}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right)$ (3.2)
for some large enough $C_{1}$, then for some $C_{2}$ the event
$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U},C_{2}\eta\log^{1/2}(\eta^{-1})),\quad\forall\mathbf{x}\in\mathcal{K}$
holds with probability at least $1-C_{3}\exp(-c_{4}\omega^{2}(\mathcal{U}^{\rm(N)})-c_{4}\mathscr{H}(\mathcal{K},r))-C_{5}\exp(-c_{6}\eta^{2}m)$.
Remark 3.1 (Recovering $\mathcal{K}=\Sigma^{n,*}_{s}$ in [10]).
Setting $(\mathcal{K},\mathcal{U})=(\Sigma^{n,*}_{s},\Sigma^{n}_{2s})$, noticing $\mathcal{K}_{(t)}=(\mathcal{K}-\mathcal{K})\cap(t\mathbb{B}_{2}^{n})\subset t(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})$ and using (2.5)–(2.6), we have
[…]
Further, setting $\eta=\frac{c\delta}{\sqrt{\log(\delta^{-1})}}$ with small enough $c$, Theorem 3.1 yields the following: The matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy $\operatorname{RIP}(\Sigma^{n}_{2s},\delta)$ (w.h.p.) as long as $m=\Omega(\delta^{-2}\log^{2}(\delta^{-1})s\log(\frac{en}{s}))$. This improves on [10, Thm. 1], which requires $m=\Omega(\delta^{-4}s\log(\frac{n\log(mn)}{\delta s}))$ for the same purpose, in terms of log factors and the dependence on $\delta$. Indeed, the dependence on $\delta$ matches that of achieving RIP via a Gaussian matrix up to log factors (e.g., see [31]).
Remark 3.2 (Arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$).
More importantly, Theorem 3.1 applies to arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$ with a number of measurements proportional to $\omega^{2}(\mathcal{K})$. To see this, by Sudakov’s inequality (2.4) and $\omega^{2}(\mathcal{K}_{(t)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})=4\omega^{2}(\mathcal{K})$ [49, Sec. 7.5.1], we find that (3.2) can be implied by the following based only on the Gaussian width,
$m\geq C_{1}^{\prime}\left(\frac{\omega^{2}(\mathcal{U}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K})}{\eta^{8}}\right)\quad\text{with large enough }C_{1}^{\prime}.$ (3.3)
The first informal theorem in introduction thus follows.
We specialize $\mathcal{K}$ to the set of approximately sparse signals $\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}$ and choose sufficiently small $\eta$, along with the sufficiency of (3.3) and (2.5), to obtain the following.
Corollary 1 (RIP of $\mathbf{A}_{\mathbf{z}}$ over approximately sparse signals).
If $m\geq C_{1}s\log(\frac{en}{s})$, then with probability at least $1-C_{2}\exp(-c_{3}s\log(\frac{en}{s}))$ over the complex Gaussian $\bm{\Phi}$, we have
$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3),\qquad\forall\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}.$
The distortion $1/3$ can be replaced by any given positive constant $\delta$, up to changes in the values of $C_{1},C_{2},c_{3}$.
The proof of Theorem 3.1 is based on covering and analogous to [10]. Nonetheless, [10] is restricted to $\mathcal{K}=\Sigma^{n,*}_{s}$ or at most other $\mathcal{K}$ with metric entropy logarithmically depending on the covering radius, and the techniques therein do not suffice for proving Corollary 1. We make a number of nontrivial modifications, with the most notable one being to introduce an additional index set when controlling the orthogonal term (see Appendix A.1). To preserve the presentation flow, we postpone the proof of Theorem 3.1 and the detailed discussions to Appendix A.
Our first main reconstruction guarantee immediately follows from Corollary 1.
Theorem 3.2 (Uniform instance optimality).
Consider the noiseless case where $\breve{\mathbf{z}}=\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ and the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ with $\hat{\mathbf{x}}$ obtained by solving $\Delta(\mathbf{A}_{\mathbf{z}};\mathbf{e}_{1};0)$ in (1.8). If $m\geq C_{1}s\log(\frac{en}{s})$ for some sufficiently large absolute constant $C_{1}$, then
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{10\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}},\qquad\forall\mathbf{x}\in\mathbb{S}^{n-1}$ (3.4)
holds with probability at least $1-C_{2}\exp(-c_{3}s\log(\frac{en}{s}))$ over the complex Gaussian $\bm{\Phi}$.
Proof.
We assume the event in Corollary 1 holds and consider any $\mathbf{x}\in\mathbb{S}^{n-1}$. If $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})$, then the event in Corollary 1 gives $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3)$, and further Proposition 2.1 implies
$\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 5\frac{\sigma_{\ell_{1}}(\mathbf{x}^{\star},\Sigma^{n}_{s})}{\sqrt{s}}=5\|\mathbf{x}^{\star}\|_{2}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}.$
Finally, we use (2.12) to obtain
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}=\left\|\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}-\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}\right\|_{2}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\|\mathbf{x}^{\star}\|_{2}}\leq 10\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}},$
as claimed. If $\mathbf{x}\notin\mathbb{B}_{1}^{n}(\sqrt{2s})$, meaning that $\|\mathbf{x}\|_{1}>\sqrt{2s}$, then we let $\mathbf{x}_{[s]}=\mathrm{arg}\min_{\mathbf{u}\in\Sigma^{n}_{s}}\|\mathbf{u}-\mathbf{x}\|_{2}$ and notice that $\|\mathbf{x}_{[s]}\|_{1}\leq\sqrt{s}$, and we have $\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})=\|\mathbf{x}-\mathbf{x}_{[s]}\|_{1}\geq\|\mathbf{x}\|_{1}-\|\mathbf{x}_{[s]}\|_{1}\geq(\sqrt{2}-1)\sqrt{s}$. Therefore,
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 2\leq\frac{2}{\sqrt{2}-1}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}\leq 10\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}.$
The proof is now complete. ∎
Remark 3.3.
While we focus on sparse recovery via basis pursuit, the generality of Theorem 3.1 in terms of $(\mathcal{K},\mathcal{U})$ allows for a straightforward generalization to other signal structures, such as PO-CS of low-rank matrices. Our subsequent technical results (Theorem 3.7, Lemmas 4.1, 4.2) are also presented in a similar manner. Also note that Theorem 3.2 directly generalizes to other instance optimal algorithms such as iterative hard thresholding (see [20, Sec. 6]).
3.2 Bounded Dense Noise
Next, we proceed to explore the robustness to noise and corruption. We consider the nonuniform robustness concerning the reconstruction of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$, which restricts our attention to robustness without being distructed by the concerns of uniformity or instance optimality. We will discuss in Section 4 the work needed to improve these forthcoming nonuniform robustness results to uniform ones with instance optimality.
We first consider pre-sign small dense noise $\bm{\tau}\in\mathbb{C}^{m}$ obeying $\|\bm{\tau}\|_{\infty}\leq\tau_{0}$, that is, observations given by $\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}$ [^5] We emphasize that the only constraint on $\bm{\tau}$ is a small enough max norm; under this constraint, it can be generated by an adversary having full knowledge of $(\bm{\Phi},\mathbf{x})$. This has been treated in [27, Sec. IV] and [10, Thm. 3]. We reproduce the result and proof here to demonstrate the two-step analysis for noisy PO-CS from [27]:
  • (1)
    Show the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$—based on the RIP of $\mathbf{A}_{\mathbf{z}}$ (Corollary 1) we only need to control the impact of noise;
  • (2)
    Estimate $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$, which is the noise level in the resulting linear compressed sensing problem, to indicate the choice of $\epsilon$ in (1.8).
Theorem 3.3 (Post-sign noise).
Consider PO-CS of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\tau}$ with $\bm{\tau}$ obeying $\|\bm{\tau}\|_{\infty}\leq\tau_{0}\leq\frac{1}{36}$. If $m\geq C_{1}s\log(\frac{en}{s})$ with large enough $C_{1}$, then the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$, with $\hat{\mathbf{x}}$ being solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\frac{5\tau_{0}}{2})$ in (1.8), satisfies
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 36\tau_{0}$
with probability at least $1-C_{2}\exp(-c_{3}s\log(\frac{en}{s}))$.
Proof.
By the linearity of $\mathbf{A}_{\mathbf{u}}$ on $\mathbf{u}$ and $\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}$, we have $\mathbf{A}_{\breve{\mathbf{z}}}=\mathbf{A}_{\mathbf{z}}+\mathbf{A}_{\bm{\tau}}$. We proceed in several steps.
Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: We first show that $\mathbf{A}_{\bm{\tau}}$ cannot have great affect on the RIP of $\mathbf{A}_{\mathbf{z}}$ under small enough $\tau_{0}$. To that end, we need to bound $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}}\mathbf{u}\|_{2}$. By (1.7) and triangle inequality, we have
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}}\mathbf{u}\|_{2}$ $\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}$ (3.5)
$\leq\tau_{0}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}+\tau_{0}\left(\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\right).$
We use a concentration bound from prior works in the area: by [10, Lem. 6] (or [19, Thm. 6]), if $m\geq C_{1}s\log(\frac{en}{s})$ for large enough $C_{1}$, then we have
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}\leq 1+c_{1}$
for some $c_{1}$ that can be set sufficiently small with probability at least $1-2\exp(-c_{2}s\log(\frac{en}{s}))$. (This can also be achieved by Lemma 4.2 appearing later.) Noting that $\bm{\Phi}^{\Re}$ and $\bm{\Phi}^{\Im}$ have i.i.d. $\mathcal{N}(0,1)$ entries, then (2.9) gives that if $m\geq C_{1}s\log(\frac{en}{s})$ for large enough $C_{1}$, we have
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\leq 2+c_{3}$
for some $c_{3}$ that can be set sufficiently small, with probability at least $1-\exp(-s\log(\frac{en}{s}))$. Setting $c_{1}=c_{3}=\frac{1}{2}$, we obtain $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}}\mathbf{u}\|_{2}\leq 4\tau_{0}\leq\frac{1}{9}$ with the promised probability. We now invoke Corollary 1 to ensure small enough $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}|\|\mathbf{A}_{\mathbf{z}}\mathbf{u}\|_{2}-1|$ when $m\geq C_{1}s\log(\frac{en}{s})$ for large enough $C_{1}$, with the desired probability. Taken collectively, we find that $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: Next, we bound
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\|\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}+\mathbf{A}_{\bm{\tau}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\|\mathbf{A}_{\bm{\tau}}\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\bm{\tau}}\mathbf{x}\|_{2}$ (3.6)
for fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ and the corresponding $\mathbf{x}^{\star}=\frac{\kappa m\cdot\mathbf{x}}{\|\bm{\Phi}\mathbf{x}\|_{1}}$. Again using (1.7) we have
$\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\tau}\mathbf{x}\|_{2}$ $\leq\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\left[\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right]$
$\leq\tau_{0}+\tau_{0}\cdot\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\cdot\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\leq\frac{5\tau_{0}}{2},$ (3.7)
where in the last inequality we use sub-Gaussian tail bounds to ensure sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ (recall that $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=O(1/\sqrt{m})$ due to (2.1)), and Bernstein’s inequality [49, Thm. 2.8.1] to ensure sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{2}^{2}}{m}-2|$, with the promised probability.
Now by Proposition 2.1 we have $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 7\cdot\frac{5\tau_{0}}{2}=\frac{35}{2}\tau_{0}$. Then by (2.12) and the condition of $\|\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}$ being sufficiently close to $1$ we have
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}=\left\|\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}-\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}\right\|_{2}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\|\mathbf{x}^{\star}\|_{2}}\leq 36\tau_{0},$ (3.8)
as claimed. ∎
While explicit constants are provided in some of our results, no attempts have been made to optimize them.
We move to post-sign small dense noise. We again denote the dense noise by $\bm{\tau}$, but now the noisy observations are $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})$. The robustness in this regime is less straightforward than the post-sign noise. The reason is that a small enough pre-sign perturbation $\tau$ can still greatly affect $\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau)$ if $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ is small. This makes the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$ less evident, for which we have to separately treat a small fraction of measurements with small $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ and the majority with $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ bounded away from 00. On the other hand, it comes a bit surprising that the error horizon remains at $O(\tau_{0})$, since some algebra with (2.12) finds $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=O(\tau_{0})$.
Theorem 3.4 (Pre-sign noise).
Consider PO-CS of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})$ with $\bm{\tau}$ obeying $\|\bm{\tau}\|_{\infty}\leq\tau_{0}\leq c_{0}$ for some small enough $c_{0}$. If $m\geq C_{1}s\log(\frac{en}{s})$ with sufficiently large $C_{1}$, then the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$, with $\hat{\mathbf{x}}$ being solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};4\tau_{0})$ in (1.8), satisfies
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 57\tau_{0}$
with probability at least $1-C_{2}\exp(-c_{3}s\log(\frac{en}{s}))$.
Proof.
We first transfer $\bm{\tau}$ to post-sign noise by writing $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\tilde{\bm{\tau}}$, where
$\tilde{\bm{\tau}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})-\operatorname{sign}(\bm{\Phi}\mathbf{x}).$
The entries of $\tilde{\bm{\tau}}$ may not be uniformly small, but we can establish a decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ where $\|\tilde{\bm{\tau}}_{1}\|_{\infty}$ is small and $\tilde{\bm{\tau}}_{2}$ is sparse. For the fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ and for some $\eta_{0}>0$ to be chosen, we have
$\mathbbm{P}(|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta_{0})\leq\mathbbm{P}\big{(}|\Re(\bm{\Phi}_{i}^{*}\mathbf{x})|\leq\eta_{0}\big{)}\leq\sqrt{\frac{2}{\pi}}\eta_{0}.$
Letting $\mathcal{J}_{\mathbf{x},\eta_{0}}=\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta_{0}\}$, the Chernoff bound gives
$\mathbbm{P}\big{(}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\big{)}\geq\mathbbm{P}\Big{(}{\rm Bin}(m,\sqrt{2/\pi}\eta_{0})\leq\eta_{0}m\Big{)}\geq 1-\exp(-c_{1}\eta_{0}m)$ (3.9)
for some $c_{1}>0$. We proceed on this event and define $(\tilde{\bm{\tau}}_{1},\tilde{\bm{\tau}}_{2})$ such that the support of $\tilde{\bm{\tau}}_{2}$ is contained in $\mathcal{J}_{\mathbf{x},\eta_{0}}$, and the support of $\tilde{\bm{\tau}}_{1}$ is contained in $\mathcal{J}_{\mathbf{x},\eta_{0}}^{c}=[m]\setminus\mathcal{J}_{\mathbf{x},\eta_{0}}$. These two requirements uniquely determine the decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$. We note that $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$ and $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq\|\tilde{\bm{\tau}}\|_{\infty}\leq 2$. Moreover, the entries of $\tilde{\bm{\tau}}_{1}$ take the form
$[\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})]\mathbbm{1}(|\bm{\Phi}_{i}^{*}\mathbf{x}|>\eta_{0}),$
and hence by (2.11) we obtain $\|\tilde{\bm{\tau}}_{1}\|_{\infty}\leq\frac{2\tau_{0}}{\eta_{0}}$. The observations can now be expressed as $\breve{\mathbf{z}}=\mathbf{z}+\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}.$
Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: We now want to show $\mathbf{A}_{\breve{\mathbf{z}}}=\mathbf{A}_{\mathbf{z}}+\mathbf{A}_{\tilde{\bm{\tau}}_{1}}+\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$. Similarly to the proof of Theorem 3.3, we can use Corollary 1 to obtain $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\bm{\Sigma}^{n}_{2s},c_{2})$ with sufficiently small $c_{2}$, and all that remains is to show $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{1}}\mathbf{u}\|_{2}$ and $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}$ are both small enough. Now let us fix $\eta_{0}$ to be some small enough positive constant. With the promised probability, the arguments from the proof of Theorem 3.3 imply
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{1}}\mathbf{u}\|_{2}\leq 4\|\tilde{\bm{\tau}}_{1}\|_{\infty}\leq\frac{8\tau_{0}}{\eta_{0}},$
which is small enough because of sufficiently small $\tau_{0}$ (and a specified $\eta_{0}>0$).
Next, we need to make sure $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}$ is sufficiently small. We let $\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\in\{0,1\}^{m}$ be the vector whose $1$’s indicate the support of $\tilde{\bm{\tau}}_{2}$. Starting as in (3.5) while proceeding with $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$, $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$, and Cauchy-Schwarz inequality, we obtain
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\tilde{\bm{\tau}}_{2}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}_{2}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}$
$\leq\|\tilde{\bm{\tau}}_{2}\|_{2}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\kappa m}+\|\tilde{\bm{\tau}}_{2}\|_{\infty}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\sqrt{m}}$
$\leq\Big{(}2+\frac{\sqrt{\eta_{0}}}{\kappa}\Big{)}\sqrt{\eta_{0}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\sqrt{\eta_{0}m}}.$ (3.10)
This can be controlled by Lemma 2.1,
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\sqrt{\eta_{0}m}}$ $\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq\eta_{0}m\end{subarray}}\left(\frac{1}{\eta_{0}m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
$\leq C_{3}\sqrt{\frac{s\log(\frac{en}{s})}{\eta_{0}m}}+C_{3}\sqrt{\log(\frac{e}{\eta_{0}})}$ (3.11)
for some absolute constant $C_{3}$ with the probability $2\exp(-c_{3}\eta_{0}\log(\frac{e}{\eta_{0}})m)$. Under $m=\Omega(s\log(\frac{en}{s}))$, substituting (3.11) into (3.10) yields
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\eta_{0}\log(\frac{e}{\eta_{0}})}\right),$ (3.12)
which is small enough due to the scaling of $m$ and sufficiently small $\eta_{0}$. Overall, we have arrived at the desired $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: Next, as with (3.6), we bound the $\ell_{2}$ measurement error at $\mathbf{x}^{\star}$,
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}\|_{2}\leq(1+c_{4})\left[\frac{|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right],$ (3.13)
where $c_{4}>0$ can be small enough due to the sub-Gaussian concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$. Letting $\tilde{\tau}_{i}=\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})$ be the $i$-th entry of $\tilde{\bm{\tau}}$, we have
$|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|=\left|\sum_{i=1}^{m}\Re\big{(}\tilde{\tau}_{i}^{*}\bm{\Phi}_{i}^{*}\mathbf{x}\big{)}\right|\leq\sum_{i=1}^{m}\big{|}\tilde{\tau}_{i}^{*}\bm{\Phi}_{i}^{*}\mathbf{x}\big{|},$ (3.14)
$\big{\|}\Im(\operatorname{diag}(\tilde{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\big{\|}_{2}=\left(\sum_{i=1}^{m}\big{[}\Im(\tilde{\tau}_{i}^{*}\bm{\Phi}_{i}^{*}\mathbf{x})\big{]}^{2}\right)^{1/2}\leq\left(\sum_{i=1}^{m}|\tilde{\tau}_{i}^{*}\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\right)^{1/2}.$ (3.15)
The key observation is that $|\tilde{\tau}_{i}^{*}\bm{\Phi}_{i}^{*}\mathbf{x}|$ is actually well bounded for any $i\in[m]$. Without loss of generality we can assume $|\bm{\Phi}_{i}^{*}\mathbf{x}|>0$; then, by (2.11) we obtain
$|\tilde{\tau}_{i}^{*}\bm{\Phi}_{i}^{*}\mathbf{x}|=|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})||\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\frac{2|\tau_{i}|}{|\bm{\Phi}_{i}^{*}\mathbf{x}|}|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq 2|\tau_{i}|\leq 2\tau_{0}.$ (3.16)
Thus, we obtain $|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|\leq 2m\tau_{0}$ and $\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}\leq 2\sqrt{m}\tau_{0}$, and hence by (3.13) with small enough $c_{4}$ we obtain $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}\leq 4\tau_{0}$. Therefore, Proposition 2.1 implies $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 28\tau_{0}$. By using an argument analogous to (3.8), we arrive at $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 57\tau_{0}$. ∎
The next theorem provides converse bounds that indicate the sharpness of the $O(\tau_{0})$ bounds in Theorems 3.33.4. Using the complex Gaussian $\bm{\Phi}$ and observations $\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})$ or $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\tau}$, we show that no algorithm can reconstruct $\mathbf{x}\in\Sigma^{n,*}_{s}$ to an error substantially smaller than $O(\tau_{0})$. The idea is to identify another signal $\mathbf{x}^{\prime}\in\Sigma^{n,*}_{s}$ that is indistinguishable from $\mathbf{x}$ and satisfies $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}=\tilde{\Omega}(\tau_{0})$. We leave the removal of log factors to future work.
Theorem 3.5 ($O(\tau_{0})$ is nearly sharp).
For the reconstruction of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ ($s\geq 4$) in PO-CS with a complex Gaussian design, we have the following:
  • In the setting of Theorem 3.4, no algorithm can guarantee an $\ell_{2}$ error smaller than $\frac{\tau_{0}}{12\sqrt{\log m}}$ with probability at least $1-\frac{4}{m}$;
  • In the setting of Theorem 3.3, assume $m=C_{0}s\log(\frac{en}{s})$ for some absolute constant $C_{0}$, then no algorithm can achieve an $\ell_{2}$ error smaller than $\frac{\tau_{0}}{48C_{0}\log(\frac{en}{s})\sqrt{\log m}}$ with probability at least $1-\frac{4}{m}-\exp(-c_{1}s)$.
Proof.
To obtain the first statement that complements Theorem 3.4, we pick $\bm{\delta}\in\Sigma^{n}_{s}$ such that
$|\operatorname{supp}(\bm{\delta})\cup\operatorname{supp}(\mathbf{x})|\leq s,\qquad\bm{\delta}^{\top}\mathbf{x}=0,\qquad\|\bm{\delta}\|_{2}=\tau_{*}$ (3.17)
for some $\tau_{*}>0$ to be chosen. Then we let $\mathbf{x}^{\prime}=\frac{\mathbf{x}+\bm{\delta}}{\|\mathbf{x}+\bm{\delta}\|_{2}}\in\Sigma^{n,*}_{s}$. When $\tau_{*}$ is small enough, we have
$\frac{\tau_{*}}{2}\leq\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}=\sqrt{\frac{\tau_{*}^{2}}{1+\tau_{*}^{2}}+\Big{(}1-\frac{1}{\sqrt{1+\tau_{*}^{2}}}\Big{)}^{2}}\leq\frac{3\tau_{*}}{2}.$ (3.18)
Then by the Gaussian tail bounds $\mathbbm{P}_{g\sim\mathcal{N}(0,1)}(|g|\geq t)\leq\exp(-\frac{t^{2}}{2})$, $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq\|\bm{\Phi}^{\Re}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}+\|\bm{\Phi}^{\Im}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}$ and $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}\leq\frac{3\tau_{*}}{2}$, a standard union bound shows that $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ with probability at least $1-\frac{4}{m}$. Letting $\tau_{*}=\frac{\tau_{0}}{6\sqrt{\log m}}$, we have identified $\mathbf{x}^{\prime}\in\Sigma^{n,*}_{s}$ satisfying
$\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}\geq\frac{\tau_{0}}{12\sqrt{\log m}}\qquad\text{and}\qquad\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq\tau_{0}.$
In the regime of Theorem 3.4, the observations $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}^{\prime})=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x}))$ can be generated through the following two indistinguishable cases:
  • The underlying signal is $\mathbf{x}$ and $\bm{\tau}=\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})$ is added by an adversary as pre-sign noise;
  • The underlying signal is $\mathbf{x}^{\prime}$ and the adversary adds nothing.
Therefore, no algorithm can distinguish $\mathbf{x}$ and $\mathbf{x}^{\prime}$, and hence no algorithm can achieve estimation error smaller than $\frac{\tau_{0}}{12\sqrt{\log m}}$.
We now move on to the second statement that complements Theorem 3.3. We consider $\mathcal{J}_{\mathbf{x},s/4m}=\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\frac{s}{4m}\}$. Then by re-iterating the arguments for (3.9), we obtain that with probability $1-\exp(-c_{1}s)$, we have $|\mathcal{J}_{\mathbf{x},s/4m}|\leq\frac{s}{4}$. Now we choose $\tilde{\bm{\delta}}\in\Sigma^{n}_{s}$ satisfying the conditions in (3.17): $|\operatorname{supp}(\tilde{\bm{\delta}})\cup\operatorname{supp}(\mathbf{x})|\leq s$, $\tilde{\bm{\delta}}^{\top}\mathbf{x}=0$ and $\|\tilde{\bm{\delta}}\|_{2}=\tau_{*}$. Additionally, we require
$\bm{\Phi}_{i}^{*}\tilde{\bm{\delta}}=0,\qquad\forall i\in\mathcal{J}_{\mathbf{x},s/4m}.$ (3.19)
On the event $\{|\mathcal{J}_{\mathbf{x},s/4m}|\leq\frac{s}{4}\}$, (3.19) translates into $\frac{s}{2}$ real linear equations, so such $\tilde{\bm{\delta}}$ exists when $s\geq 4$. Now we consider $\mathbf{x}^{\prime}=\frac{\mathbf{x}+\tilde{\bm{\delta}}}{\|\mathbf{x}+\tilde{\bm{\delta}}\|_{2}}\in\Sigma^{n,*}_{s}$, and similarly to (3.18) we have $\frac{\tau_{*}}{2}\leq\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}\leq\frac{3\tau_{*}}{2}$, and $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ holds with probability at least $1-\frac{4}{m}$. We further bound $\|\operatorname{sign}(\bm{\Phi}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}\mathbf{x})\|_{\infty}$. For $i\in\mathcal{J}_{\mathbf{x},s/4m}$, we have $\bm{\Phi}_{i}^{*}\tilde{\bm{\delta}}=0$ and hence $\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}^{\prime})=\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})$. For $i\notin\mathcal{J}_{\mathbf{x},s/4m}$, (2.11) gives
$|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})|\leq\frac{2|\bm{\Phi}_{i}^{*}(\mathbf{x}^{\prime}-\mathbf{x})|}{s/(4m)}\leq 48C_{0}\log\big{(}\frac{en}{s}\big{)}\sqrt{\log m}\cdot\tau_{*},$
where the last inequality we use $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ and substitute $m=C_{0}s\log(\frac{en}{s})$. It follows that
$\|\operatorname{sign}(\bm{\Phi}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}\mathbf{x})\|_{\infty}\leq 48C_{0}\log\big{(}\frac{en}{s}\big{)}\sqrt{\log m}\cdot\tau_{*}.$
Setting
$\tau_{*}=\frac{\tau_{0}}{48C_{0}\log(\frac{en}{s})\sqrt{\log m}},$
we have identified $\mathbf{x},\mathbf{x}^{\prime}\in\Sigma^{n,*}_{s}$ such that
$\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}\geq\frac{\tau_{0}}{96C_{0}\log(\frac{en}{s})\sqrt{\log m}}\qquad\text{and}\qquad\|\operatorname{sign}(\bm{\Phi}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}\mathbf{x})\|_{\infty}\leq\tau_{0}.$
Therefore, no algorithm can distinguish $\mathbf{x}$ and $\mathbf{x}^{\prime}$ in the regime of Theorem 3.3 where an adversary can add post-sign noise bounded by $\tau_{0}$. ∎
Simulation: [^6] The MATLAB codes for generating the figures in this paper are available in https://junrenchen58.github.io/. We pause to use experimental results to provide evidence of the achievability and tightness of $O(\tau_{0})$. In all of our experiments, the data points are averaged over $50$ independent trials, each of which concerns the recovery of $\mathbf{x}$ uniformly drawn from $\Sigma^{500,*}_{5}$ from $300$ phase-only measurements. We provide the optimally tuned $\varepsilon$ to basis pursuit (1.8), namely $\varepsilon=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$. For the post-sign noise, we test $\tau_{0}\in\{0.04,0.08,0.12,0.16,\cdots,0.36,0.40\}$ and adopt such corruption pattern: find $\theta_{0}\in[0,\frac{\pi}{2}]$ such that $|e^{\mathbf{i}\theta_{0}}-1|=\tau_{0}$ and then corrupt $\mathbf{z}$ to $\breve{\mathbf{z}}=e^{\mathbf{i}\theta_{0}}\mathbf{z}$. For the pre-sign noise, we test $\tau_{0}\in\{0.04,0.12,0.20,0.28,\cdots,0.76,0.84\}$ and generate the noisy observations through $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\tau_{0}\operatorname{sign}(\bm{\Phi}\mathbf{x})\mathbf{i})$. The results are given in Figures 12 and are consistent with our theorems.
Figure 1: Reconstruction errors under post-sign bounded by $\tau_{0}$
Figure 2: Reconstruction errors under pre-sign bounded by $\tau_{0}$
3.3 Sparse Phase Corruption
We consider a corruption $\bm{\zeta}$ that can affect a small fraction of the observations arbitrarily over the complex phases. We suppose that there is an adversary (with full knowledge of $\bm{\Phi}$ and $\mathbf{x}$) that can change any $\zeta_{0}m$ measurements to arbitrary phase-only values. [^7] Note that our formulation (3.20) offers slightly more generality. This setting resembles the adversarial bit flips widely considered in the 1-bit compressed sensing literature [40, 11, 34, 15] where $\zeta_{0}m$ signs can be flipped. The mathematical formulation is given by
$\breve{\mathbf{z}}=\mathbf{z}+\bm{\zeta}$ (3.20)
for some $\bm{\zeta}\in\mathbb{C}^{m}$ satisfying $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ and $\|\bm{\zeta}\|_{\infty}\leq 2$.
Remark 3.4.
We consider (3.20) only, as the case of pre-sign corruption $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\zeta})$ can be written as $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\tilde{\bm{\zeta}}$ where $\tilde{\bm{\zeta}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\zeta})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$ satisfies $\|\tilde{\bm{\zeta}}\|\leq\zeta_{0}m$ and $\|\tilde{\bm{\zeta}}\|_{\infty}\leq 2$.
We show that $\mathbf{x}^{\sharp}$ is robust to sparse corruption, in that the $\zeta_{0}m$ adversarial attacks can only increment the estimation error by $O(\sqrt{\zeta_{0}\log(1/\zeta_{0})})$.
Theorem 3.6 (Sparse corruption).
Consider PO-CS of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\zeta}$ with $\bm{\zeta}$ obeying $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ for some small enough $\zeta_{0}$ and $\|\bm{\zeta}\|_{\infty}\leq 2$. If $m\geq C_{1}s\log(\frac{en}{s})$ with sufficiently large $C_{1}$, then the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$, with $\hat{\mathbf{x}}$ being solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};11\zeta_{0}\log(\frac{e}{\zeta_{0}}))$ in (1.8), satisfies
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 71\sqrt{\zeta_{0}\log(e/{\zeta_{0}})}$
with probability at least $1-C_{2}\exp(-c_{3}s\log(\frac{en}{s}))-\exp(-\zeta_{0}m\log(\frac{e}{\zeta_{0}}))$ for some absolute constants $C_{2},c_{3}$.
Proof.
Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: The first step is to show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$. As we can use Corollary 1 to show $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},c_{1})$ for sufficiently small $c_{1}$ with the promised probability, we only need to ensure small enough $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$. In the proof of Theorem 3.4 we show (3.12) for $\tilde{\bm{\tau}}_{2}$ satisfying $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$ and $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$. An identical argument yields
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\zeta_{0}\log(e/\zeta_{0})}\right)$
with the probability we promise. Thus, $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ is small enough by the scaling of $m$ and $\zeta_{0}$ being small enough.
Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: The second step is to bound the $\ell_{2}$ measurement error at $\mathbf{x}^{\star}=\frac{\kappa m\cdot\mathbf{x}}{\|\bm{\Phi}\mathbf{x}\|_{1}}$,
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\leq(1+c_{2})\left[\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right]$ (3.21)
where $c_{2}>0$ is small enough due to the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$. We let $\mathbf{1}_{\operatorname{supp}(\bm{\zeta})}\in\{-1,1\}^{m}$ whose $1$’s indicate the support set of $\bm{\zeta}$. Then we have
$\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}$
$\leq\frac{\|\bm{\zeta}\|_{2}\|\bm{\Phi}\mathbf{x}\odot\mathbf{1}_{\operatorname{supp}(\bm{\zeta})}\|_{2}}{\kappa m}+\frac{2\|\bm{\Phi}\mathbf{x}\odot\mathbf{1}_{\operatorname{supp}(\bm{\zeta})}\|_{2}}{\sqrt{m}}$
$\leq\frac{2}{\sqrt{m}}\Big{(}1+\frac{\sqrt{\zeta_{0}}}{\kappa}\Big{)}\|\bm{\Phi}\mathbf{x}\odot\mathbf{1}_{\operatorname{supp}(\bm{\zeta})}\|_{2}$
$\leq\frac{2}{\sqrt{m}}\Big{(}1+\frac{\sqrt{\zeta_{0}}}{\kappa}\Big{)}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=\zeta_{0}m\end{subarray}}\left(\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\right)^{1/2}.$ (3.22)
Without loss of generality, we assume $\zeta_{0}m$ is a positive integer. For fixed $I$ with cardinality $\zeta_{0}m$, $\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}$ follows Chi-squared distribution with $2\zeta_{0}m$ degrees of freedom. Then by a standard concentration bound [30, Lem. 1], we obtain that for any $t\geq 0$ and any $I\subset[m]$ with $|I|=\zeta_{0}m$,
$\mathbbm{P}\left(\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\leq 2\zeta_{0}m+2\sqrt{2\zeta_{0}mt}+2t\right)\geq 1-\exp(-t).$
Taking a union bound over $\binom{m}{\zeta_{0}m}$ possible $I$, it yields
$\mathbbm{P}\left(\max_{\begin{subarray}{c}I\subset[m]\\ |I|=\zeta_{0}m\end{subarray}}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\leq 2\zeta_{0}m+2\sqrt{2\zeta_{0}mt}+2t\right)\geq 1-\exp\Big{(}\zeta_{0}m\log\big{(}\frac{e}{\zeta_{0}}\big{)}-t\Big{)}.$
Setting $t=2\zeta_{0}m\log(\frac{e}{\zeta_{0}})$ and using the small enough $\zeta_{0}$, we arrive at
$\max_{\begin{subarray}{c}I\subset[m]\\ |I|=\zeta_{0}m\end{subarray}}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\leq 5\zeta_{0}m\log\big{(}\frac{e}{\zeta_{0}}\big{)}$ (3.23)
with probability at least $1-\exp(-\zeta_{0}m\log(\frac{e}{\zeta_{0}}))$. Combining (3.21), (3.22) and (3.23), using small enough $\zeta_{0},c_{2}$, we obtain $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}\leq 5\sqrt{\zeta_{0}\log(e/\zeta_{0})}$.
Combining the two prior steps, we invoke Proposition 2.1 to obtain $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 37\sqrt{\zeta_{0}\log(e/{\zeta_{0}})}$. This leads to $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 71\sqrt{\zeta_{0}\log(e/\zeta_{0})}$ by re-iterating (3.8). ∎
We expect that the error increment $\tilde{O}(\sqrt{\zeta_{0}})$ is tight for the specific estimator $\mathbf{x}^{\sharp}$. To support this, without considering the normalization $\mathbf{x}^{\sharp}=\hat{\mathbf{x}}/\|\hat{\mathbf{x}}\|_{2}$, we show $\Omega(\delta_{1}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$ is a lower bound on $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}$ under a suboptimal noise parameter $\varepsilon\geq(1+\delta_{1})\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$ for basis pursuit (1.8). In practice, this assumption can often be satisfied, and the near-optimal choice $\varepsilon=(1+o(1))\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$ could be unrealistic.
Proposition 3.1.
In the problem setting of Theorem 3.6, consider $\hat{\mathbf{x}}$ solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\varepsilon)$ in (1.8). If $\varepsilon\geq(1+\delta_{1})\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$ for some $\delta_{1}>0$, then under sparse phase corruption $\bm{\zeta}$ that changes the $\zeta_{0}m$ measurements with the largest $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ from $z_{i}$ to $\breve{z}_{i}=\mathbf{i}\cdot z_{i}$, for some absolute constant $c_{1}$ we have
$\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\geq c_{1}\delta_{1}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$
with probability at least $1-C_{2}\exp(-c_{3}s\log(\frac{en}{s}))-C_{4}\exp(-c_{5}\zeta_{0}\log(\frac{e}{\zeta_{0}})m)$.
The key idea is to show all $s$-sparse signals within the ball $\mathbb{B}_{2}^{n}(\mathbf{x}^{\star};\Theta(\delta_{1}\sqrt{\zeta_{0}\log(e/\zeta_{0})}))$ satisfy the constraint $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}-\mathbf{e}_{1}\|_{2}\leq\varepsilon$. Then, we argue that some signal living on the boundary of this ball is favored over $\mathbf{x}^{\star}$ by the decoder in (1.8). Since this statement is positioned as a secondary result, its proof is postponed to Appendix B.1.
Simulation: We pause to provide numerical evidence on the sharpness of $\tilde{O}(\sqrt{\zeta_{0}})$ for $\mathbf{x}^{\sharp}$, even under the optimally tuned noise level $\varepsilon=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$. We adopt the same settings as in earlier simulations but replace the dense noise $\bm{\tau}$ by $\zeta_{0}m$ adversarial phase corruptions. We test $\zeta_{0}m=\{1,2,3,5,7,9,11,13\}$ and corrupt the measurements through the mechanism described in Proposition 3.1. The log-log curve in Figure 3(Left) roughly has a slope of $\frac{1}{2}$ over small $\zeta_{0}$. This seems to suggest the tightness of $\tilde{O}(\sqrt{\zeta_{0}})$ for the specific estimator $\mathbf{x}^{\sharp}$.
Figure 3: The impact of sparse phase corruption on $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$
As related context, in 1-bit compressed sensing, $\zeta_{0}m$ adversarial bit flips increment the $\ell_{2}$ error of the convex relaxation approach [40] by $\tilde{O}(\sqrt{\zeta_{0}})$, which was then improved to $\tilde{O}(\zeta_{0})$ using different algorithms [12, 11, 34, 1], and this (almost) linear increment is near-optimal under Gaussian designs (e.g., see [39, Thm. 2.4]).
It is thus natural to investigate the tightness of $\tilde{O}(\sqrt{\zeta_{0}})$ in Theorem 3.6 without constraining the algorithm. We show that $\tilde{O}(\sqrt{\zeta_{0}})$ is indeed suboptimal and the impact of the sparse corruption can indeed be eliminated, meaning that there is an algorithm being capable of perfectly recovering $\mathbf{x}$ in this regime. As linear system or compressed sensing with sparse corruption [21, 24, 35, 37], the intuition is that the uncorrupted measurements remain numerous enough to uniquely identify the signal. More specifically, we achieve this through an efficient algorithm, which is an extension of the linearization approach. It reformulates corrupted PO-CS as linear compressed sensing with sparse corruption [37, 21, 7], which can also be simply viewed as a noiseless extended linear compressed sensing problem.
We consider the setting in Theorem 3.6. Combining $\mathbf{z}=\breve{\mathbf{z}}-\bm{\zeta}$ and $\Im(\frac{1}{\sqrt{m}}\operatorname{diag}(\mathbf{z}^{*})\bm{\Phi})\mathbf{x}=0$ as in (1.4), we arrive at
$\frac{1}{\sqrt{m}}\Im(\operatorname{diag}(\breve{\mathbf{z}}^{*})\bm{\Phi})\mathbf{x}+\mathbf{x}_{\bm{\zeta}}=0,$ (3.24)
where $\mathbf{x}_{\bm{\zeta}}:=\frac{1}{\sqrt{m}}\Im(\operatorname{diag}(\bm{\zeta})\bm{\Phi})\mathbf{x}$ is $(\zeta_{0}m)$-sparse. Since (3.24) does not contain any norm information on $(\mathbf{x},\mathbf{x}_{\bm{\zeta}})$, as was done in (1.5), we further introduce
$\frac{1}{\kappa m}\Re(\breve{\mathbf{z}}^{*}\bm{\Phi})\mathbf{x}=1$ (3.25)
to address the scaling issue. We are faced with a noiseless linear compressed sensing problem with extended signal space, whose goal is to find $(\mathbf{x},\mathbf{x}_{\bm{\zeta}})\in\Sigma^{n}_{s}\times\Sigma^{m}_{\zeta_{0}m}$ that satisfies (3.24) and (3.25). We define an extended new sensing matrix for any $\mathbf{w}\in\mathbb{C}^{m}$ as
$\widetilde{\mathbf{A}}_{\mathbf{w}}:=\begin{bmatrix}\frac{1}{\kappa m}\Re(\mathbf{w}^{*}\bm{\Phi})&0~{}\\ \frac{1}{\sqrt{m}}\Im(\operatorname{diag}(\mathbf{w})^{*}\bm{\Phi})&\mathbf{I}_{m}\end{bmatrix},$ (3.26)
and then the linear constraints (3.24) and (3.25) can be concisely expressed as
$\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}\begin{bmatrix}\mathbf{x}\\ \mathbf{x}_{\bm{\zeta}}\end{bmatrix}=\mathbf{e}_{1}.$ (3.27)
Remark 3.5.
Like $\mathbf{x}^{\star}$ in (3.1) such that $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$, the ground truth satisfying (3.27) is given by
$\mathbf{x}^{\star\star}=\frac{\kappa m\cdot\mathbf{x}}{\Re(\breve{\mathbf{z}}^{*}\bm{\Phi}\mathbf{x})}\quad\text{and}\quad\mathbf{x}_{\zeta}^{\star\star}=\frac{\Im(\operatorname{diag}(\bm{\zeta})\bm{\Phi}\mathbf{x}^{\star\star})}{\sqrt{m}}.$
We propose to find $(\mathbf{x},\mathbf{x}_{\bm{\zeta}})$ by solving weighted $\ell_{1}$-norm minimization
$(\hat{\mathbf{x}}_{e},\hat{\mathbf{x}}_{\bm{\zeta}})=\arg\min~{}\frac{\|\mathbf{u}\|_{1}}{\sqrt{s}}+\frac{\|\mathbf{w}\|_{1}}{\sqrt{\zeta_{0}m}},~{}~{}\text{subject to~{}}\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}\begin{bmatrix}\mathbf{u}\\ \mathbf{w}\end{bmatrix}=\mathbf{e}_{1}$ (3.28)
and then use $\mathbf{x}^{\sharp}_{e}=\hat{\mathbf{x}}_{e}/\|\hat{\mathbf{x}}_{e}\|_{2}$ as our final estimate. By establishing the RIP of $\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}$ over $\Sigma^{n}_{2s}\times\Sigma^{m}_{2\zeta_{0}m}$, we obtain the perfect reconstruction $\mathbf{x}^{\sharp}_{e}=\mathbf{x}$. Let us present the RIP of $\widetilde{\mathbf{A}}_{\mathbf{z}}$ for a fixed $\mathbf{x}$ and then the exact reconstruction guarantee.
Theorem 3.7 (RIP of $\widetilde{\mathbf{A}}_{\mathbf{z}}$ for fixed $\mathbf{x}$).
Consider $\mathcal{U}_{c}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ for some cones $\mathcal{U}_{1}\subset\mathbb{R}^{n}$ and $\mathcal{U}_{2}\subset\mathbb{R}^{m}$, fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and given $\delta\in(0,1)$. For some absolute constants $C_{1}$ and $c_{2}$, if $m\geq C_{1}\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\rm(N)})$, then $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},\delta)$ with probability at least $1-\exp(-c_{2}\delta^{2}m)$.
The proof of Theorem 3.7 can be found in Appendix A.4. We can improve it to be a uniform statement over $\mathbf{x}\in\mathcal{K}$ for some $\mathcal{K}\subset\mathbb{S}^{n-1}$; see Lemma A.1.
Theorem 3.8 (Perfect recovery under corruption).
Consider the same signal reconstruction problem as in Theorem 3.6 while using a different estimator $\mathbf{x}^{\sharp}_{e}=\frac{\hat{\mathbf{x}}_{e}}{\|\hat{\mathbf{x}}_{e}\|_{2}}$, where $\hat{\mathbf{x}}_{e}$ is obtained by solving (3.28). If $m\geq C_{1}s\log(\frac{en}{s})$ for some sufficiently large absolute constant $C_{1}$, $\zeta_{0}$ is small enough, then we have $\hat{\mathbf{x}}_{e}=\mathbf{x}$ with probability at least $1-\exp(-c_{2}m)-2\exp(-c_{3}\zeta_{0}\log(e/\zeta_{0})m)$.
Proof.
We first show $\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\mathcal{U}_{c},\frac{1}{3})$ where $\mathcal{U}_{c}=\Sigma^{n}_{2s}\times\Sigma^{m}_{2\zeta_{0}m}$. By $\mathcal{U}_{c}^{\rm(N)}\subset(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})\times(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{n})$, we have $\omega(\mathcal{U}_{c}^{{\rm(N)}})\leq\omega(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+\omega(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})$ and hence
$\omega^{2}(\mathcal{U}_{c}^{\rm(N)})\leq 2\omega^{2}(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+2\omega^{2}(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})\leq C_{0}\Big{(}s\log\big{(}\frac{en}{s}\big{)}+\zeta_{0}\log\big{(}\frac{e}{\zeta_{0}}\big{)}m\Big{)}$
for some absolute constant $C_{0}$. Therefore, when $\zeta_{0}$ is sufficiently small, $m\geq C_{1}s\log(\frac{en}{s})$ with large enough $C_{1}$ implies $m\geq C_{2}\omega^{2}(\mathcal{U}_{c}^{\rm(N)})$ with large enough $C_{2}$. Then, by Theorem 3.7 we have $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},c_{3})$ with some small enough $c_{3}\leq\frac{1}{12}$ with probability at least $1-\exp(-c_{4}m)$. All that remains is to ensure $\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\rm(N)}}\|(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\|_{2}$ to be sufficiently small. Comparing $\mathbf{A}_{\mathbf{w}}$ in (1.7) and $\widetilde{\mathbf{A}}_{\mathbf{w}}$ in (3.26), we have
$\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\rm(N)}}\big{\|}(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\big{\|}_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\big{\|}(\mathbf{A}_{\breve{\mathbf{z}}}-\mathbf{A}_{\mathbf{z}})\mathbf{u}\big{\|}_{2}=\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}.$
In the proofs of Theorems 3.4 and 3.6 we have shown the bound
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}=O\left(\sqrt{\frac{s\log(en/s)}{m}}+\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})}\right)$
with probability at least $1-2\exp(-c_{5}\zeta_{0}\log(\frac{e}{\zeta_{0}})m)$; see (3.10)–(3.12). Thus, $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ is small enough by the scaling of $m$ and small enough $\zeta_{0}.$
We have thus proved $\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\mathcal{U}_{c},\frac{1}{3})$. By Proposition 2.2 and the observation made in Remark 3.5, we have $\hat{\mathbf{x}}_{e}=\mathbf{x}^{\star\star}=\frac{\kappa m\cdot\mathbf{x}}{\Re(\breve{\mathbf{z}}^{*}\bm{\Phi}\mathbf{x})}$. To show $\mathbf{x}_{e}^{\sharp}=\mathbf{x}$, it remains to show $\frac{\Re(\breve{\mathbf{z}}^{*}\bm{\Phi}\mathbf{x})}{\kappa m}>0$. This can be seen from
$\frac{\Re(\breve{\mathbf{z}}^{*}\bm{\Phi}\mathbf{x})}{\kappa m}=\frac{\Re(\mathbf{z}^{*}\bm{\Phi}\mathbf{x})+\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})}{\kappa m}=\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{1}{2}-\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}\geq\frac{1}{4},$
where the last two inequalities hold with the promised probability due to the sub-Gaussian concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$ and the proven sufficiently small $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$. ∎
4 Robust Instance Optimality
The main aim of this section is to consolidate our prior results to show that $\mathbf{x}^{\sharp}$ is robust and instance optimal over the entire signal space $\mathbb{S}^{n-1}$. We consider the noisy phase-only observations
$\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)}+\bm{\zeta}_{(1)})+\bm{\tau}_{(2)}+\bm{\zeta}_{(2)},$
where the bounded dense noise vectors $\bm{\tau}_{(1)},\bm{\tau}_{(2)}\in\mathbb{C}^{m}$ satisfy $\|\bm{\tau}_{(1)}\|_{\infty}\leq\tau_{0}$ and $\|\bm{\tau}_{(2)}\|_{\infty}\leq\tau_{0}$, $\bm{\zeta}_{(1)},\bm{\zeta}_{(2)}\in\mathbb{C}^{m}$ are $(\zeta_{0}m)$-sparse, and the post-sign corruption $\bm{\zeta}_{(2)}$ additionally satisfies $\|\bm{\zeta}_{(2)}\|_{\infty}\leq 2$. Here, $(\bm{\tau}_{(1)},\bm{\tau}_{(2)},\bm{\zeta}_{(1)},\bm{\zeta}_{(2)})$ may be generated by an adversary and can depend on $(\bm{\Phi},\mathbf{x})$.
We first announce our result, which is novel for nonlinear compressed sensing and closely resembles the standard guarantee in linear case (e.g., Proposition 2.1). It is also a formal version of the second informal statement provided in introduction.
Theorem 4.1 (Robust, instance optimal & uniform).
Consider the above setting with sufficiently small $\tau_{0}$ and $\zeta_{0}$, and the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ where $\hat{\mathbf{x}}$ is solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\varepsilon)$ in (1.8) with
$\varepsilon=C_{1}\tau_{0}+C_{2}\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})}+C_{3}\sqrt{\frac{s\log({en}/{s})}{m}}$ (4.1)
for some large enough absolute constants $C_{1},C_{2},C_{3}$. If $m\geq C_{4}s\log(\frac{en}{s})$ for a large enough absolute constant $C_{4}$, then with high probability over the complex Gaussian $\bm{\Phi}$, we have
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{10\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}+15\varepsilon,\qquad\forall\mathbf{x}\in\mathbb{S}^{n-1}.$ (4.2)
By similar techniques, along with slightly more work, one can prove that $\mathbf{x}^{\sharp}_{e}$ in Theorem 3.8 satisfies (4.2) with the error increments of sparse corruption eliminated and $\varepsilon=O(\tau_{0})$, up to changes of constants. We omit the details to avoid repetition.
4.1 Proof Strategy
To prove Theorem 4.1, we essentially need to run the arguments in Theorems 3.2, 3.3, 3.4 and 3.6 again, along with some additional steps to make certain arguments uniform. To avoid repetition, we will only outline the proof strategies and emphasize the additional techniques.
Without loss of generality, we can assume that the supports of $\bm{\zeta}_{1}$ and $\bm{\zeta}_{2}$ are disjoint. By Remark 3.4 we can consolidate $\bm{\zeta}_{1}$ and $\bm{\zeta}_{2}$ to rewrite $\breve{\mathbf{z}}$ as
$\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})+\bm{\tau}_{(2)}+\bm{\zeta},$
where $\bm{\zeta}\in\mathbb{C}^{m}$ satisfies $\|\bm{\zeta}\|_{0}\leq 2\zeta_{0}m$ and $\|\bm{\zeta}\|_{\infty}\leq 2$. Recall also that we can focus on $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}:=\mathcal{X}$ (see the proof of Theorem 3.2). Further, we let $\tilde{\bm{\tau}}:=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$ and write
$\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}_{(2)}+\tilde{\bm{\tau}}+\bm{\zeta}.$ (4.3)
We then perform the two-step analysis. As we shall see, extra care is needed to ensure that every piece of the prior arguments is uniform for all $\mathbf{x}\in\mathcal{X}$, and the major additional work is on a uniform upper bound on $|\mathcal{J}_{\mathbf{x},\eta_{0}}|=|\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta_{0}\}|$ and a uniform concentration bound on $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$. We proceed in several steps.
Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: Corollary 1 yields $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},c_{1})$ for some sufficiently small $c_{1}$, and one can check that the relevant arguments in the proofs of Theorems 3.3 and 3.6 can show $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{u}\|_{2}$ and $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ are sufficiently small. Therefore, it remains to show $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{u}\|_{2}$ to be small enough. In the proof of Theorem 3.4, the central idea is the decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ where $\tilde{\bm{\tau}}_{1}$ is near-dense and has entries bounded by $\frac{2\tau_{0}}{\eta_{0}}$, and $\tilde{\bm{\tau}}_{2}$ is $(\eta_{0}m)$-sparse with support set $\mathcal{J}_{\mathbf{x},\eta_{0}}$ and has entries bounded by $2$. We obtain the desired result by setting $\eta_{0}$ as a small absolute constant because we can then re-iterate the arguments for small dense noise to treat $\tilde{\bm{\tau}}_{1}$, and these for sparse phase corruption to treat $\tilde{\bm{\tau}}_{2}$. Such idea and most arguments remain valid, while the only notable issue is that the proof of Theorem 3.4 only shows $\|\tilde{\bm{\tau}}_{2}\|_{0}=|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m$ for a fixed $\mathbf{x}$ and hence only ensures the existence of the decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ for this $\mathbf{x}$. We need to strengthen this step and show that such a decomposition exists for all $\mathbf{x}\in\mathcal{X}$; to that end, we need to show $\sup_{\mathbf{x}\in\mathcal{X}}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m$. This is still true with high probability under $m=\Omega(s\log(\frac{en}{s}))$, as guaranteed by the following lemma. See a more refined statement and the proof in Appendix B.2.
Lemma 4.1.
Given some sufficiently small $\eta\in[\frac{C_{1}}{m},1)$ for some absolute constant $C_{1}$ and some $\mathcal{K}\subset\mathbb{S}^{n-1}$, if for sufficiently large $C_{2}$ we have $m\geq C_{2}\eta^{-3}\log(\eta^{-1})\omega^{2}(\mathcal{K})$, then we have
$\mathbbm{P}\left(\sup_{\mathbf{x}\in\mathcal{K}}~{}|\mathcal{J}_{\mathbf{x},\eta}|\leq\eta m\right)\geq 1-3\exp(-c_{3}\eta m).$
Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: By (4.3) and $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$, we seek to bound
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$ $\leq\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{x}^{\star}\|_{2}+\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}^{\star}\|_{2}+\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}$
$=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\Big{(}\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{x}\|_{2}+\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}\|_{2}+\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\Big{)}.$
Recall that we repeatedly use the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$; e.g., see (3.7), (3.13), (3.21)). While this follows from sub-Gaussian concentration for a fixed $\mathbf{x}$, we need to strengthen it to uniform concentration over all $\mathbf{x}\in\mathcal{X}$. The following lemma shows that $\sup_{\mathbf{x}\in\mathcal{X}}|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ is small enough under $m=\Omega(s\log(\frac{en}{s}))$. Up to some simple modifications (from $\mathbb{R}$ to $\mathbb{C}$), the proof is identical to that of [42, Lem. 2.1], and is hence omitted.
Lemma 4.2.
Suppose that the entries of $\bm{\Phi}\in\mathbb{C}^{m\times n}$ are drawn i.i.d. from $\mathcal{N}(0,1)+\mathcal{N}(0,1)\mathbf{i}$ and let $\kappa=\sqrt{\frac{\pi}{2}}$. Given $\mathcal{T}\subset\mathbb{S}^{n-1}$, for some absolute constants $C_{1},c_{2}$ the event
$\mathbbm{P}\left(\sup_{\mathbf{u}\in\mathcal{T}}\Big{|}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}-1\Big{|}\leq\frac{C_{1}(\omega(\mathcal{T})+t)}{\sqrt{m}}\right)\geq 1-2\exp(-c_{2}t).$
Now, for $\mathbf{w}=\bm{\tau}_{(2)},\tilde{\bm{\tau}},\bm{\zeta}$, it remains to bound
$\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\mathbf{w}}\mathbf{x}\|_{2}\leq\sup_{\mathbf{x}\in\mathcal{X}}\frac{|\Re(\mathbf{w}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\Im(\operatorname{diag}(\mathbf{w}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}.$
We have
$\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{x}\|_{2}\leq\tau_{0}\left[\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\right]=O(\tau_{0})$
due to (2.10). For $\tilde{\bm{\tau}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$, we also have $\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}\|_{2}=O(\tau_{0})$ as our original arguments in (3.14)–(3.16) are already uniform over $\mathbf{x}\in\mathcal{X}$.
The bound on $\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}$ is a bit more tricky. From (3.22) we have
$\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\leq C_{1}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=\zeta_{0}m\end{subarray}}\left(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\right)^{1/2},$
and for a fixed $\mathbf{x}$ we achieve the bound $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}=O(\sqrt{\zeta_{0}\log(1/\zeta_{0})})$ through (3.23). In contrast, here we need a uniform bound over $\mathbf{x}\in\mathcal{X}$, and we use Lemma 2.1 to obtain
$\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\leq C_{1}\sup_{\mathbf{x}\in\mathcal{X}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=\zeta_{0}m\end{subarray}}\left(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\right)^{1/2}=O\left(\sqrt{\zeta_{0}\log(e/\zeta_{0})}+\sqrt{\frac{s\log(en/s)}{m}}\right).$
Therefore, a new term $O(\sqrt{\frac{s}{m}\log(\frac{en}{s})})$ arises in the error horizon, and consequently it also appears in our choice of the noise level (4.1).
5 Conclusion
In this paper, we analyzed the instance optimality and robustness of the recently proposed linearization approach for PO-CS [27], in which one reformulates PO-CS as linear compressed sensing and then solves it via quadratically constrained basis pursuit. We improved the nonuniform instance optimality in [27] to a uniform one over the entire sphere. The new technical tool is the RIP for all the new sensing matrices corresponding to an arbitrary set of signals in the unit sphere, which we proved by making important improvements on the arguments in [10].
Beyond Theorem 3.3 known from [27], we provided a new set of robustness results. First, dense noise bounded by small enough $\tau_{0}$ (either before or after taking the phases) increments the estimation error by $O(\tau_{0})$, and no algorithm can do substantially better than this. Second, an adversarial $\zeta_{0}$-fraction of sparse corruption increments the error by $\tilde{O}(\sqrt{\zeta_{0}})$. We conjectured that this is tight for our specific estimator and provided some evidence. Yet we showed that it can be improved to 00 by proposing an extended linearization approach which perfectly recovers sparse signal under sparse corruption.
We believe the following questions are interesting for future study:
  • Non-Gaussian sensing matrix. All existing recovery guarantees (that are exact in noiseless case) are built upon complex Gaussian $\bm{\Phi}$. Can we develop similar results for sub-Gaussian matrices or structured sensing matrices?
  • New algorithms & RIPless analysis. Existing works are building on the same linearization approach and similar RIP analysis. Can we develop new algorithms with comparable theoretical guarantee for PO-CS? Without linearization, can we directly analyze the original nonlinear phase-only observations?
  • Instance optimality in nonlinear sensing. Are there similar instance optimal results in other nonlinear sensing problems?
Acknowledgement
The authors are thankful to Prof. Zhaoqiang Liu for his helps and discussions in the numerical simulations. Junren Chen is also thankful to Prof. Laurent Jacques for some stimulating discussions on instance optimality and PO-CS.
Appendix
Appendix A RIP for New Sensing Matrices (Theorems 3.1 & 3.7)
Here we present the proofs for the RIP of the new sensing matrix $\mathbf{A}_{\mathbf{z}}$ (Theorem 3.1) and the extended new sensing matrix $\widetilde{\mathbf{A}}_{\mathbf{z}}$ (Theorem 3.7). We prove the following more general statement that asserts the RIP of $\widetilde{\mathbf{A}}_{\mathbf{z}}$ defined in (3.26) over a subset of $\mathbb{S}^{n-1}$.
Lemma A.1.
Suppose $\mathcal{U}_{c}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ for some cones $\mathcal{U}_{1}\subset\mathbb{R}^{n}$ and $\mathcal{U}_{2}\subset\mathbb{R}^{m}$. There exist some absolute constants $c_{1},C_{1},C_{2},C_{3},c_{4},C_{5},c_{6}$ such that for any $\eta\in(0,c_{1})$, if we let $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$ and $\delta_{\eta}=C_{1}\eta(\log(\eta^{-1}))^{1/2}$, if
$m\geq C_{2}\left[\frac{\omega^{2}(\mathcal{U}_{c}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right],$ (A.1)
then the event
$\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},\delta_{\eta}),\quad\forall\mathbf{x}\in\mathcal{K}$ (A.2)
holds with probability at least $1-C_{3}\exp(-c_{4}\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-c_{4}\mathscr{H}(\mathcal{K},r))-C_{5}\exp(-c_{6}\eta^{2}m)$.
We first show that this statement immediately leads to Theorem 3.1.
Proof of Theorem 3.1.
Setting $\mathcal{U}_{1}=\mathcal{U}$ and $\mathcal{U}_{2}=0$ in Lemma A.1 yields Theorem 3.1. ∎
In the remainder of this subsection, we will first establish a number of intermediate bounds, and then combine them to prove Lemma A.1.
By homogeneity and some algebra, we find that (A.2) is equivalent to
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\underbrace{\frac{[\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}]^{2}}{\kappa^{2}m^{2}}+\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}^{2}-1}_{:=f(\mathbf{x},\mathbf{u},\mathbf{w})}\Big{|}\leq\delta_{\eta}.$ (A.3)
Given the underlying signal $\mathbf{x}\in\mathbb{S}^{n-1}$ and some $\mathbf{u}\in\mathbb{R}^{n}$, we have the decomposition
$\mathbf{u}=\underbrace{\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x}}_{:=\mathbf{u}_{\mathbf{x}}^{\|}}+\underbrace{(\mathbf{u}-\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x})}_{:=\mathbf{u}_{\mathbf{x}}^{\bot}}$ (A.4)
where $\|\mathbf{u}_{\mathbf{x}}^{\|}\|_{2}^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}=\|\mathbf{u}\|_{2}^{2}.$ We recall that $\kappa:=\mathbbm{E}|\Phi_{i,j}|=\sqrt{\frac{\pi}{2}}$ and $\mathbf{u}_{\mathbf{x}}^{\bot}=\mathbf{u}-\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x}$, and observe that $1=\|\mathbf{u}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}=|\langle\mathbf{u},\mathbf{x}\rangle|^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}$ for $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$. Hence, we decompose $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|$ into
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.5)
where
$f^{\|}(\mathbf{x},\mathbf{u})=\frac{[\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}]^{2}}{\kappa^{2}m^{2}}-|\langle\mathbf{u},\mathbf{x}\rangle|^{2}$ (A.6)
$f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})=\left\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\right\|_{2}^{2}-\big{(}\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}\big{)}.$ (A.7)
We refer to (A.6) as the parallel term and (A.7) as the orthogonal term. We first control $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}f^{\|}(\mathbf{x},\mathbf{u})$ in Lemma A.3 and $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}f^{\bot}(\mathbf{x},\mathbf{u})$ in Lemma A.5, and then strengthen them to uniform bounds over $\mathbf{x}\in\mathcal{K}$ by covering arguments in Lemma A.6 and Lemma A.7, respectively.
A.1 Technical Contributions
Since the proof is lengthy and builds on existing work [10], we first pause to discuss the key differences and innovations. Specifically, some improvement is necessary to deal with arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$.
Main Improvement—Finer Treatment to Phase Perturbation via Introducing $\mathcal{I}_{\mathbf{x}-\mathbf{x}_{r},\eta^{\prime}}$: The most notable improvement is made when seeking a uniform bound on the orthogonal term (see our Lemma A.7). In the analysis, we need to bound
$\frac{1}{\sqrt{m}}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{\|}\Im\big{[}\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big{]}\big{\|}_{2}$
$=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i=1}^{m}\Big{[}\Im\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{]}^{2}\right)^{1/2}$
where $\mathbf{x}_{r}:=\mathrm{arg}\min_{\mathbf{u}\in\mathcal{N}_{r}}\|\mathbf{u}-\mathbf{x}\|_{2}$ is the point in $\mathcal{N}_{r}$ (which is a minimal $r$-net of $\mathcal{K}$) closest to $\mathbf{x}$. To control this term to be sufficiently small, the idea is to separate the $m$ measurements into two parts—a small “problematic part” $\mathcal{E}_{\mathbf{x}}$ where $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|$ is hard to control, and the “major part” $\mathcal{E}_{\mathbf{x}}^{c}=[m]\setminus\varepsilon_{\mathbf{x}}$.
In [10], the authors simply let $\mathcal{E}_{\mathbf{x}}=\mathcal{J}_{\mathbf{x},\eta}$ (with some small enough $\eta$) and then apply
$|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|\leq 2\qquad\text{for }i\in\mathcal{E}_{\mathbf{x}},$ (A.8)
$|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|\leq\frac{2|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|}{\eta}\qquad\text{for }i\in\mathcal{E}_{\mathbf{x}}^{c}.$ (A.9)
Thus, with respect to (A.9), they have to show the following is small enough:
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i\notin\mathcal{J}_{\mathbf{x},\eta}}\Big{[}\Im\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{]}^{2}\right)^{1/2}$
$\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
$\leq\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{v}|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}.$ (A.10)
This is a heavy-tailed random process that is in general hard to control.
The argument in [10] is to use extremely small $r$ with $o(1)$ scaling to ensure this term to be small enough. Note that their equations (III.63)–(III.64) essentially bound (A.10) as
$O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\|\bm{\Phi}\mathbf{v}\|_{\infty}\cdot\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\eta\sqrt{m}}\right)=O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\frac{\|\bm{\Phi}\mathbf{v}\|_{\infty}}{\eta}\right).$
By Gaussian concentration, it is easy to show that, with high probability,
$\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\frac{\|\bm{\Phi}\mathbf{v}\|_{\infty}}{\eta}\geq\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\frac{\|\bm{\Phi}^{\Re}\mathbf{v}\|_{1}}{\eta m}=\Omega\left(\frac{\omega(\mathcal{K}_{(r)})}{\eta}\right),$ (A.11)
and this is also an upper bound on $\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\eta^{-1}\|\bm{\Phi}\mathbf{v}\|_{\infty}$, up to log factors, due to Lemma 2.1. Therefore, for $\mathcal{K}=\Sigma^{n,*}_{s}$, the authors of [10] take $r=\tilde{O}(\frac{\eta}{\sqrt{s}})$ to guarantee sufficiently small
$\frac{\omega(\mathcal{K}_{(r)})}{\eta}=\Theta\left(\frac{r\sqrt{s\log(\frac{en}{s})}}{\eta}\right).$
Since $\mathscr{H}(\Sigma^{n,*}_{s},r)$ logarithmically depends on $r$ (2.6), such choice of $r$ only adds to log factors in the sample complexity. However, such small $r$ will significantly worsens the sample complexity (from $\tilde{O}(s\log(\frac{en}{s}))$ to $\Omega(s^{2}\log(\frac{en}{s}))$) for $\mathcal{K}=\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{S}^{n-1}$ whose metric entropy quadratically depends on the covering radius; see (2.7).
To make an improvement, we first notice that their choice of $\mathcal{E}_{\mathbf{x}}$ and (A.8)–(A.9) are suboptimal: for $i\in\mathcal{E}_{\mathbf{x}}^{c}$, the bound $\frac{2|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|}{\eta}$ they used could be worse than $2$ when $|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|\geq\eta$, and indeed by (A.11), $|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|$ could reach $\Omega(\omega(\mathcal{K}_{(r)}))$ for some $i$. (We note that $\omega(\mathcal{K}_{(r)})=\Theta(r\sqrt{s\log(\frac{en}{s})})$ for $\mathcal{K}=\Sigma^{n}_{s}$, while it even scales as $\omega(\mathcal{K}_{(r)})=\Theta(r\sqrt{n})$ for $\mathcal{K}=\sqrt{s}\mathbb{B}_{1}^{n}$ under $r\leq\sqrt{s/n}$.) Moreover, the heavy-tailed random process arises from the contribution of $|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|$.
Our remedy is to use $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|\leq 2$ for the measurements with overly large $|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|$. To formalise this idea, for some small enough $\eta^{\prime}>0$ to be chosen, we introduce
$\mathcal{I}_{\mathbf{x}-\mathbf{x}_{r},\eta^{\prime}}:=\big{\{}i\in[m]:|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|>\eta^{\prime}\big{\}}$
and define the set of problematic measurements as $\mathcal{E}_{\mathbf{x}}:=\mathcal{J}_{\mathbf{x},\eta}\cup\mathcal{I}_{\mathbf{x}-\mathbf{x}_{r},\eta^{\prime}}$. We use $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|\leq 2$ for $i\in\mathcal{E}_{\mathbf{x}}$. This is valid as $|\mathcal{E}_{\mathbf{x}}|$ remains small—one can invoke Lemma 2.1 to uniformly control $|\mathcal{I}_{\mathbf{x}-\mathbf{x}_{r},\eta^{\prime}}|$ over $\mathbf{x}-\mathbf{x}_{r}\in\mathcal{K}_{(r)}$. On the other hand, for $i\in\mathcal{E}_{\mathbf{x}}^{c}$ we now have a bound $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|\leq\frac{2}{\eta}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|\leq\frac{2\eta^{\prime}}{\eta},$ which is better than $2$ when we use $\eta^{\prime}<\eta$, and in the proof we guarantee small enough $\frac{2\eta^{\prime}}{\eta}$ by setting $\eta^{\prime}\ll\eta$. This also avoids the heavy-tailed random process in (A.10).
Other Refinements: We also briefly note that we have refined or simplified some steps. As an example, in contrast to the covering approach taken in [10], we directly use known concentration bounds to establish the uniformity over $\mathcal{U}_{c}^{\rm(N)}$. Consequently, the sample complexity (A.1) is only based on the Gaussian width of $\mathcal{U}_{c}^{\rm(N)}$ and is free of its metric entropy. As another example, while [10, Lem. 9] seeks to uniformly bound $|\mathcal{J}_{\mathbf{x},\eta}|$ over $\mathbf{x}\in\mathcal{K}=\Sigma^{n,*}_{s}$, we find that bounding $|\mathcal{J}_{\mathbf{x},\eta}|$ over the $r$-net of $\mathcal{K}$ is sufficient.
A.2 Fixed-$\mathbf{x}$ Bound on the Parallel Term
Observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$ implies $\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}$, and that $f^{\|}(\mathbf{x},\mathbf{u})$ does not depend on $\mathbf{w}$. We start by utilizing
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}}|f^{\|}(\mathbf{x},\mathbf{u})|=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|,$ (A.12)
where the last equality follows from $|f^{\|}(\mathbf{x},t\mathbf{u})|=t^{2}|f^{\|}(\mathbf{x},\mathbf{u})|$ for $t>0$. By $a^{2}-b^{2}=(a-b)(a+b)$, we have
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|$ $\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{\rm(N)}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big{|}\right)\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big{|}\right)$
$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big{|},$ (A.13)
where (A.13) follows from
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big{|}\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\|\mathbf{z}\|_{2}\cdot\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa m}\Big{|}+1\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa\sqrt{m}}+1=O(1),$ (A.14)
which holds with probability at least $1-4\exp(-cm)$ provided that $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see (2.10). Therefore, we only need to bound the term in (A.13), which reads as
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}-\langle\mathbf{x},\mathbf{u}\rangle\Big{|}:=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ (A.15)
We note that $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ is zero-mean. To see this, we use the decomposition $\mathbf{u}=\mathbf{u}_{\mathbf{x}}^{\|}+\mathbf{u}_{\mathbf{x}}^{\bot}$ (A.4), and then by the independence between $(\bm{\Phi}_{i}^{*}\mathbf{x},\bm{\Phi}_{i}^{*}\mathbf{u}_{\mathbf{x}}^{\bot})$, we have
$\mathbbm{E}\big{[}\kappa^{-1}\Re\big{(}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\big{]}=\mathbbm{E}\big{[}\kappa^{-1}\Re\big{(}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}\mathbf{u}_{\mathbf{x}}^{\|}\big{)}\big{]}=\langle\mathbf{x},\mathbf{u}\rangle\kappa^{-1}\mathbbm{E}|\bm{\Phi}_{i}^{*}\mathbf{x}|=\langle\mathbf{x},\mathbf{u}\rangle.$ (A.16)
Because $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ is linear in $\mathbf{u}$, $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ can be treated as the supremum of a standard random process and hence be directly controlled by the following lemma.
Lemma A.2 (See Sec. 8.6 in [49]).
Let $(R_{\mathbf{u}})_{\mathbf{u}\in\mathcal{T}}$ be a random process (not necessarily zero-mean) on a subset $\mathcal{T}\subset\mathbb{R}^{n}$. Assume that $R_{0}=0$, and for all $\mathbf{u},\mathbf{v}\in\mathcal{T}\cup\{0\}$ we have $\|R_{\mathbf{u}}-R_{\mathbf{v}}\|_{\psi_{2}}\leq K\|\mathbf{u}-\mathbf{v}\|_{2}$. Then, for every $t\geq 0$, the event
$\sup_{\mathbf{u}\in\mathcal{T}}\big{|}R_{\mathbf{u}}\big{|}\leq CK\big{(}\omega(\mathcal{T})+t\cdot\operatorname{rad}(\mathcal{T})\big{)}$
holds with probability at least $1-2\exp(-t^{2})$.
We then use it to bound $\sup_{\mathbf{u}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ for a fixed $\mathbf{x}$.
Lemma A.3.
Consider $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}f_{1}^{\|}(\mathbf{x},\mathbf{u})$ as in (A.15) with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and $\mathcal{U}_{1}^{\rm(N)}\subset\mathbb{S}^{n-1}$. Then for some absolute constant $C$ and any $t\geq 0$, we have
$\mathbbm{P}\left(\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C[\omega(\mathcal{U}_{1}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2}).$ (A.17)
Proof.
For any $\mathbf{u}_{1},\mathbf{u}_{2}\in\mathcal{U}_{1}^{\rm(N)}\cup\{0\}$, we have
$\big{\|}f_{1}^{\|}(\mathbf{x},\mathbf{u}_{1})-f_{1}^{\|}(\mathbf{x},\mathbf{u}_{2})\big{\|}_{\psi_{2}}$
$=\frac{1}{\kappa}\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\Re\big{(}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}(\mathbf{u}_{1}-\mathbf{u}_{2})\big{)}-\kappa\langle\mathbf{x},\mathbf{u}_{1}-\mathbf{u}_{2}\rangle\Big{\|}_{\psi_{2}}$
$\leq\frac{C_{1}}{\sqrt{m}}\big{\|}\Re\big{(}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}(\mathbf{u}_{1}-\mathbf{u}_{2})\big{)}-\kappa\langle\mathbf{x},\mathbf{u}_{1}-\mathbf{u}_{2}\rangle\big{\|}_{\psi_{2}}$ (A.18)
$\leq\frac{C_{1}}{\sqrt{m}}\Big{(}\big{\|}\Re\big{(}\overline{z_{i}}\bm{\Phi}_{i}^{*}(\mathbf{u}_{1}-\mathbf{u}_{2})\big{)}\big{\|}_{\psi_{2}}+\big{\|}\kappa\langle\mathbf{x},\mathbf{u}_{1}-\mathbf{u}_{2}\rangle\big{\|}_{\psi_{2}}\Big{)},$
where in (A.18) we note $\mathbbm{E}[\Re\big{(}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}(\mathbf{u}_{1}-\mathbf{u}_{2})\big{)}]=\kappa\langle\mathbf{x},\mathbf{u}_{1}-\mathbf{u}_{2}\rangle$ (e.g., see (A.16)) and then apply (2.1). Moreover, we note that $\|\Re(\overline{z_{i}}\bm{\Phi}_{i}^{*}(\mathbf{u}_{1}-\mathbf{u}_{2}))\|_{\psi_{2}}\leq C_{2}\|\mathbf{u}_{1}-\mathbf{u}_{2}\|_{2}$ because
$\|\Re(z_{i}\bm{\Phi}_{i})\|_{\psi_{2}}\leq\|z_{i}^{\Re}\bm{\Phi}_{i}^{\Re}\|_{\psi_{2}}+\|z_{i}^{\Im}\bm{\Phi}_{i}^{\Im}\|_{\psi_{2}}\leq\|\bm{\Phi}_{i}^{\Re}\|_{\psi_{2}}+\|\bm{\Phi}_{i}^{\Im}\|_{\psi_{2}}=O(1).$
In addition, the simple upper bound $|\kappa\langle\mathbf{x},\mathbf{u}_{1}-\mathbf{u}_{2}\rangle|\leq\kappa\|\mathbf{u}_{1}-\mathbf{u}_{2}\|_{2}$ implies $\|\kappa\langle\mathbf{x},\mathbf{u}_{1}-\mathbf{u}_{2}\rangle\|_{\psi_{2}}\leq C_{3}\|\mathbf{u}_{1}-\mathbf{u}_{2}\|_{2}.$ Therefore, we have shown
$\|f_{1}^{\|}(\mathbf{x},\mathbf{u}_{1})-f_{1}^{\|}(\mathbf{x},\mathbf{u}_{2})\|_{\psi_{2}}\leq\frac{C_{4}}{\sqrt{m}}\|\mathbf{u}_{1}-\mathbf{u}_{2}\|_{2}$
for some absolute constant $C_{4}$. Now we invoke Lemma A.2 to obtain (A.17), as desired. ∎
A.3 Fixed-$\mathbf{x}$ Bound on the Orthogonal Term
To deal with the orthogonal term $f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ defined in (A.7), we begin with
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ $\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}\right)$
$\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}+\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}\right)$
$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\underbrace{\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}}_{:=f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})}\Big{|},$ (A.19)
where the last inequality follows from $\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}\leq 1$ and
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}$ $\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{*}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}+1=O(1),$
which holds with probability at least $1-4\exp(-cm)$ under $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see (2.10).
It remains to bound $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$, where $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ is defined in (A.19). We first bound $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ for a fixed $\mathbf{x}$ by using the following extended matrix deviation inequality, along with the rotational invariance of $\bm{\Phi}$.
Lemma A.4 (see [7]).
Let $\mathbf{A}$ be an $m\times n$ matrix whose rows $\mathbf{A}_{i}$ are independent centered isotropic sub-Gaussian vectors in $\mathbb{R}^{n}$. Given any bounded subset $\mathcal{T}\subset\mathbb{R}^{n}\times\mathbb{R}^{m}$ and $t\geq 0$, the event
$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{T}}\Big{|}\big{\|}\mathbf{A}\mathbf{u}+\sqrt{m}\mathbf{w}\big{\|}_{2}-\sqrt{m}\cdot\sqrt{\|\mathbf{u}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}\leq CK^{2}\big{(}\omega(\mathcal{T})+t\cdot\operatorname{rad}(\mathcal{T})\big{)}$
holds with probability at least $1-\exp(-t^{2})$, where $K=\max_{i}\|\mathbf{A}_{i}\|_{\psi_{2}}$.
Lemma A.5.
Let $\mathbf{x}\in\mathbb{S}^{n-1}$ be fixed and $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ be given in (A.19). Then, for any $t\geq 0$, we have
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C[\omega(\mathcal{U}_{c}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2}).$
Proof.
We find an orthogonal matrix $\mathbf{P}$ such that $\mathbf{P}\mathbf{x}=\mathbf{e}_{1}$ and consider $\tilde{\bm{\Phi}}=\bm{\Phi}\mathbf{P}^{\top}$, which has the same distribution as $\bm{\Phi}$. Due to $\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi}\mathbf{x})=0$, we have $\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}=\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}_{\mathbf{x}}^{\bot}$, and hence we can write
$f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})=\left\|\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}\mathbf{e}_{1}))^{*})\tilde{\bm{\Phi}})\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}}{\sqrt{m}}+\mathbf{w}\right\|_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}~{}.$
By $\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}\in\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathcal{U}^{\rm(N)}=\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathbf{P}^{\top}\mathbf{P}\mathcal{U}^{\rm(N)}=(\mathbf{I}_{n}-\mathbf{e}_{1}\mathbf{e}_{1}^{\top})\mathbf{P}\mathcal{U}^{\rm(N)}$, we know the first entry of $\tilde{\mathbf{u}}:=\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}$ is always 00. We further observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$ implies
$\begin{bmatrix}\tilde{\mathbf{u}}\\ \mathbf{w}\end{bmatrix}=\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ \mathbf{w}\end{bmatrix}\in\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\mathcal{U}_{c}^{\rm(N)}:=\tilde{\mathcal{U}}_{0}.$
We let $\tilde{\mathbf{u}}_{1}\in\mathbb{R}^{n-1}$ be the restriction of $\tilde{\mathbf{u}}\in\mathbb{R}^{n}$ to the last $n-1$ entries, and let $\tilde{\mathcal{U}}\subset\mathbb{R}^{m+n-1}$ be the restriction of $\tilde{\mathcal{U}}_{0}\subset\mathbb{R}^{m+n}$ to the last $m+n-1$ entries, so that we have $(\tilde{\mathbf{u}}_{1}^{\top},\mathbf{w}^{\top})^{\top}\in\tilde{\mathcal{U}}$. This can be more precisely formulated as
$\tilde{\mathcal{U}}=\mathbf{R}\tilde{\mathcal{U}}_{0},\qquad\text{where }\mathbf{R}=[0,\mathbf{I}_{m+n-1}]\in\mathbb{R}^{(m+n-1)\times(m+n)}.$
In light of [49, Exercise 7.5.4] we have $\omega(\tilde{\mathcal{U}})\leq\omega(\tilde{\mathcal{U}}_{0})\leq\omega(\mathcal{U}_{c}^{\rm(N)})$.
With these preparations, we let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}^{[1]},\tilde{\bm{\Phi}}^{[2:n]}]$ where $\tilde{\bm{\Phi}}^{[1]}$ is the first column of $\tilde{\bm{\Phi}}$ and $\tilde{\bm{\Phi}}^{[2:n]}\in\mathbb{R}^{m\times(n-1)}$ is composed by the last $n-1$ columns of $\tilde{\bm{\Phi}}$, and can proceed as follows:
$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.20)
$\qquad=\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}\mathbf{e}_{1}))^{*}\tilde{\bm{\Phi}})\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}$
$\qquad\leq\sup_{(\tilde{\mathbf{u}},\mathbf{w})\in\tilde{\mathcal{U}}_{0}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}\mathbf{e}_{1}))^{*}\tilde{\bm{\Phi}})\tilde{\mathbf{u}}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\tilde{\mathbf{u}}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}$
$\qquad=\sup_{(\tilde{\mathbf{u}}_{1},\mathbf{w})\in\tilde{\mathcal{U}}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}^{[1]}))^{*}\tilde{\bm{\Phi}}^{[2:n]})\tilde{\mathbf{u}}_{1}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\tilde{\mathbf{u}}_{1}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}$ (A.21)
$\qquad\leq\sup_{(\tilde{\mathbf{u}}_{1},\mathbf{w})\in\tilde{\mathcal{U}}}\Big{|}\Big{\|}\frac{\hat{\bm{\Phi}}\tilde{\mathbf{u}}_{1}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\tilde{\mathbf{u}}_{1}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|},$
where $\hat{\bm{\Phi}}:=\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}^{[1]}))^{*}\tilde{\bm{\Phi}}^{[2:n]})$ has the same distribution as a matrix with i.i.d. $\mathcal{N}(0,1)$ entries almost surely. Recalling $\omega(\tilde{\mathcal{U}})\leq\omega(\mathcal{U}_{c}^{\rm(N)})$ and observing $\operatorname{rad}(\tilde{\mathcal{U}})\leq\operatorname{rad}(\tilde{\mathcal{U}}_{0})\leq\operatorname{rad}(\mathcal{U}_{c}^{\rm(N)})\leq 1$, a straightforward application of Lemma A.4 yields the desired claim. ∎
A.4 Proof of Theorem 3.7 (RIP of $\widetilde{\mathbf{A}}_{\mathbf{z}}$ for Fixed $\mathbf{x}$)
Proof.
We let $\mathcal{K}=\{\mathbf{x}\}$ for a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$. Substituting the fixed-$\mathbf{x}$ bound in Lemma A.3 into (A.12)–(A.13) yields
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C_{1}[\omega(\mathcal{U}_{1}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2})-4\exp(-c_{2}m).$ (A.22)
Similarly, we substitute the fixed-$\mathbf{x}$ bound in Lemma A.5 into (A.19) to obtain
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{3}[\omega(\mathcal{U}_{c}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2})-4\exp(-c_{4}m).$ (A.23)
By (A.5), $m=\Omega(\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\rm(N)}))$, and letting $t=c_{5}\sqrt{m}\delta$ with sufficiently small $c_{5}$ in (A.22)–(A.23), we obtain
$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\delta\right)\geq 1-11\exp(-c_{6}\delta^{2}m).$
Note that this event is the same as $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},\delta)$, so the proof is complete. ∎
A.5 Uniform (All-$\mathbf{x}$) Bound on the Parallel Term
We further extend Lemma A.3 to a uniform bound for all $\mathbf{x}\in\mathcal{K}$ by a covering argument.
Lemma A.6.
Under the setting of Lemma A.3, let $\mathcal{K}$ be an arbitrary subset of $\mathbb{S}^{n-1}$. There exist absolute constants $c_{1},C_{2},C_{3},C_{4}$ such that for any $\eta\in(0,c_{1})$ and $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$, if
$m\geq\frac{C_{2}}{\log(\eta^{-1})}\left(\frac{\omega^{2}(\mathcal{U}_{1}^{\rm(N)})}{\eta^{2}}+\frac{\mathscr{H}(\mathcal{K},r)}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}}\right),$ (A.24)
then with probability at least $1-C_{3}\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$, we have
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq C_{4}\eta\sqrt{\log(\eta^{-1})}.$
Proof.
We first extend the bound in Lemma A.3 to an $r$-net $\mathcal{N}_{r}$ of $\mathcal{K}$ and then bound the approximation error induced by approximating $\mathbf{x}\in\mathcal{K}$ by $\mathbf{x}\in\mathcal{N}_{r}$. For clarity we break down the proof into several pieces.
Uniform Bound on an $r$-Net: For some $r>0$ that will be chosen later, we let $\mathcal{N}_{r}$ be a $r$-net of $\mathcal{K}$ that is minimal in that $\log|\mathcal{N}_{r}|=\mathscr{H}(\mathcal{K},r)$. Then by Lemma A.3 and a union bound, we obtain
$\mathbbm{P}\left(\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq\frac{C(\omega(\mathcal{U}_{1}^{\rm(N)})+t)}{\sqrt{m}}\right)\geq 1-2\exp\big{(}\mathscr{H}(\mathcal{K},r)-t^{2}\big{)}$
for any $t\geq 0$. Therefore, setting $t=\Theta(\sqrt{\mathscr{H}(\mathcal{K},r)}+\omega(\mathcal{U}_{1}^{\rm(N)}))$ yields that the event
$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq\frac{C_{1}[\omega(\mathcal{U}_{1}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},r)}]}{\sqrt{m}}$ (A.25)
holds with probability at least $1-2\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$.
Bounding the Number of Small Measurements: For small enough $\eta>0$, we recall $\mathcal{J}_{\mathbf{x},\eta}:=\big{\{}i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\big{\}}.$ We now bound $|\mathcal{J}_{\mathbf{x},\eta}|$ over $\mathbf{x}\in\mathcal{N}_{r}$. For a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$, by $\Re(\bm{\Phi}_{i}^{*}\mathbf{x})\sim\mathcal{N}(0,1)$, we have $p_{0}:=\mathbbm{P}\big{(}|\bm{\Phi}_{i}\mathbf{x}|\leq\eta\big{)}\leq\mathbbm{P}\big{(}|\Re(\bm{\Phi}_{i}^{*}\mathbf{x})|\leq\eta\big{)}\leq\sqrt{\frac{2}{\pi}}\eta.$ Note that $|\mathcal{J}_{\mathbf{x},\eta}|\sim\text{Bin}(m,p_{0})$, so the Chernoff bound (e.g., [36, Sec. 4.1]) gives $\mathbbm{P}\big{(}|\mathcal{J}_{\mathbf{x},\eta}|\geq\eta m\big{)}\leq\exp(-c_{1}\eta m)$ for some absolute constant $c_{1}$. Thus, a union bound over $\mathbf{x}\in\mathcal{N}_{r}$ gives
$\mathbbm{P}\left(\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|<\eta m\right)\geq 1-\exp\Big{(}\mathscr{H}(\mathcal{K},r)-c_{1}\eta m\Big{)},$ (A.26)
which holds with probability at least $1-\exp(-c_{2}\eta m)$ as long as $m\geq\frac{C_{3}\mathscr{H}(\mathcal{K},r)}{\eta}$ for large enough $C_{3}$. The remainder of the proof proceeds on the events (A.25) and (A.26).
Bounding the Approximation Error: We seek to understand the gap between $\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ and $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ For any $\mathbf{x}\in\mathcal{K}$ we let $\mathbf{x}_{r}=\mathrm{arg}\min_{\mathbf{u}\in\mathcal{N}_{r}}\|\mathbf{u}-\mathbf{x}\|_{2}$. Here, $\mathbf{x}_{r}$ depends on $\mathbf{x}$, but we drop such dependence to avoid cumbersome notation. Note that $\|\mathbf{x}-\mathbf{x}_{r}\|_{2}\leq r$, and indeed we have $\mathbf{x}-\mathbf{x}_{r}\in\mathcal{K}_{(r)}=(\mathcal{K}-\mathcal{K})\cap\mathbb{B}_{2}^{n}(r).$ For clarity we consider a given $\mathbf{x}\in\mathcal{K}$, while we note beforehand that the final bounds in (A.30) and (A.33) hold uniformly for all $\mathbf{x}\in\mathcal{K}$. Now by $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ defined in (A.15), we calculate that
$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}-\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big{|}\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})-f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big{|}$ (A.27)
$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}\langle\mathbf{x}_{0}-\mathbf{x}_{r},\mathbf{u}\rangle\big{|}$
$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}+r.$ (A.28)
To bound the first term in (A.28), we first divide it into two terms according to $\mathcal{J}_{\mathbf{x}_{r},\eta}$:
$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}$
$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}$
$+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}:=I_{1}+I_{2}.$ (A.29)
Bounding $I_{1}$: On the event of (A.26) we have $|\mathcal{J}_{\mathbf{x}_{r},\eta}|<\eta m$. Thus, we simply use the universal bound $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|\leq 2$ to get
$I_{1}$ $\leq\frac{1}{\kappa m}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\big{|}\bm{\Phi}_{i}^{*}\mathbf{u}\big{|}^{2}\right)^{1/2}$
$\leq\frac{2\eta}{\kappa}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq\eta m\end{subarray}}\left(\frac{1}{\eta m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}\leq C_{4}\eta\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{\eta m}}+\sqrt{\log(\eta^{-1})}\right),$ (A.30)
where the last inequality follows from Lemma 2.1 and holds with probability at least $1-2\exp(-c_{5}\eta m\log(\eta^{-1}))$.
Bounding $I_{2}$: When $i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}$ we have $|\bm{\Phi}_{i}^{*}\mathbf{x}_{r}|>\eta$, and hence (2.11) gives $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|\leq\frac{2}{\eta}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|$. This allows us to proceed as follows:
$I_{2}\leq\frac{2}{\kappa\eta}\frac{1}{m}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\big{|}\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})\big{|}\big{|}\bm{\Phi}_{i}^{*}\mathbf{u}\big{|}$ (A.31)
$\leq\frac{2}{\kappa\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$ (A.32)
$\leq\frac{C_{5}}{\eta}\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{v}|^{2}\right)^{1/2}\leq\frac{C_{6}}{\eta}\left(\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{m}}+r\right),$ (A.33)
where in (A.33) we use $\mathbf{x}-\mathbf{x}_{r}\in\mathcal{K}_{(r)}$ and (2.10) which holds with probability at least $1-4\exp(-cm)$, and in the second inequality we use (2.9) with $t=\sqrt{m}$ to obtain
$\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{v}|^{2}\right)^{1/2}\leq\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\frac{\|\bm{\Phi}^{\Re}\mathbf{v}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\frac{\|\bm{\Phi}^{\Im}\mathbf{v}\|_{2}}{\sqrt{m}}=O\left(\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{m}}+r\right)$
with probability at least $1-4\exp(-m)$.
Substituting (A.30) and (A.33) into (A.29) we can bound the first term in (A.28) as
$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\right|\leq C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.34)
Completing the Proof: Note that $\mathbf{x}$ in (A.27) can be arbitrary point in $\mathcal{K}$, and our bound (A.34) is uniform for all $\mathbf{x}\in\mathcal{K}$. Substituting (A.34) into (A.27)–(A.28) and taking the supremum over $\mathcal{K}$, we obtain
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}$ $\leq\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}+r$
$+C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right),$
and then we further apply (A.25), as well as $\eta<1$, to arrive at
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq C_{8}\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{\frac{\mathscr{H}(\mathcal{K},r)}{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.35)
In summary, this bound holds with probability at least $1-C_{9}\exp(-c_{10}\eta m)-C_{11}\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathcal{H}(\mathcal{K},r))$ under the sample complexity $m=\Omega(\frac{\mathscr{H}(\mathcal{K},r)}{\eta}+\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see such sample complexity and probability term from (A.26), (A.25), (A.30) and (A.29). We now set $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$ as in our statement, and the bound in (A.35) reads as
$O\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{\frac{\mathscr{H}(\mathcal{K},r)}{m}}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}\right).$ (A.36)
Under the assumed sample complexity (A.24), the bound (A.36) scales as $O(\eta\sqrt{\log(\eta^{-1})})$, and we can promise a probability at least $1-C_{12}\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$ for the claim to hold. This completes the proof. ∎
A.6 Uniform (All-$\mathbf{x}$) Bound on the Orthogonal Term
Similarly to Section A.5, we strengthen Lemma A.5 to a universal bound over $\mathbf{x}\in\mathcal{K}$. Our major technical refinement over [10, Lem. 14] lies in the introduction of $\mathcal{I}_{\mathbf{u},\eta^{\prime}}$.
Lemma A.7.
In the setting of Lemma A.5, there exist some absolute constants $c_{1},C_{2},C_{3},c_{4},C_{5},c_{6},C_{7}$ such that given any $\eta\in(0,c_{1})$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if
$m\geq C_{2}\left(\frac{\omega^{2}(\mathcal{U}_{c}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}\right),$ (A.37)
then with probability at least $1-C_{3}\exp(-c_{4}[\omega^{2}(\mathcal{U}_{c}^{\rm(N)})+\mathscr{H}(\mathcal{K},\eta^{3})])-C_{5}\exp(-c_{6}\eta^{2}m)$, we have
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}\leq C_{7}\eta\sqrt{\log(\eta^{-1})}.$
Proof.
The proof will be presented in several steps.
Uniform Bound on an $r$-Net: For some $r>0$ that will be chosen later, we let $\mathcal{N}_{r}$ be an $r$-net of $\mathcal{K}$ that is minimal in that $\log|\mathcal{N}_{r}|=\mathscr{H}(\mathcal{K},r)$. We apply Lemma A.5 to every $\mathbf{x}\in\mathcal{N}_{r}$, along with a union bound, to obtain that for any $t\geq 0$, the event
$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{1}(\omega(\mathcal{U}_{c}^{\rm(N)})+t)}{\sqrt{m}}$
holds with probability at least $1-2\exp(\mathscr{H}(\mathcal{K},r)-t^{2})$. Therefore, setting $t^{2}=\Theta(\omega^{2}(\mathcal{U}_{c}^{\rm(N)})+\mathscr{H}(\mathcal{K},r))$ yields that the event
$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}\leq\frac{C_{2}(\omega(\mathcal{U}_{c}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},r)})}{\sqrt{m}}$ (A.38)
holds with probability at least $1-\exp(-\omega^{2}(\mathcal{U}_{c}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$.
Bounding the Number of Small Measurements: As shown in the proof of Lemma A.6, under the sample complexity $m\geq\frac{C_{3}\mathscr{H}(\mathcal{K},r)}{\eta}$ we have $\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|<\eta m$ with probability at least $1-\exp(-c_{4}\eta m)$, where $\mathcal{J}_{\mathbf{x},\eta}=\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\}$. We will utilize this event.
Bounding the Approximation Error: For any $\mathbf{x}\in\mathcal{K}$, we let $\mathbf{x}_{r}=\mathrm{arg}\min_{\mathbf{u}\in\mathcal{N}_{r}}\|\mathbf{u}-\mathbf{x}\|_{2}$. Note that $\|\mathbf{x}-\mathbf{x}_{r}\|_{2}\leq r$ and we have $\mathbf{x}-\mathbf{x}_{r}\in\mathcal{K}_{(r)}$. For clarity we consider a given $\mathbf{x}\in\mathcal{K}$, but we note that the forthcoming arguments hold uniformly for all $\mathbf{x}\in\mathcal{K}.$ By $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ defined in (A.19) and $\|\mathbf{u}-\mathbf{v}\|_{2}\leq|\|\mathbf{u}\|_{2}-\|\mathbf{v}\|_{2}|$, we can utilize the decomposition
$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}-\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big{|}$ (A.39)
$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})-f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big{|}$
$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}\Big{|}$ (A.40)
$\quad\quad\quad\quad+\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\begin{bmatrix}\mathbf{u}_{\mathbf{x}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big{\|}_{2}-\Big{\|}\begin{bmatrix}\mathbf{u}_{\mathbf{x}_{r}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big{\|}_{2}\Big{|}.$ (A.41)
For any $\mathbf{u}\in\mathbb{S}^{n-1}$, $\mathbf{x}\in\mathcal{K}$ and its associated $\mathbf{x}_{r}\in\mathcal{N}_{r}$, by the triangle inequality we have
(the term in (A.41)) $\leq\big{|}\|\mathbf{u}^{\bot}_{\mathbf{x}}\|_{2}-\|\mathbf{u}_{\mathbf{x}_{r}}^{\bot}\|_{2}\big{|}\leq\big{\|}\mathbf{u}_{\mathbf{x}}^{\bot}-\mathbf{u}_{\mathbf{x}_{r}}^{\bot}\big{\|}_{2}=\big{\|}[\mathbf{u}-\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x}]-[\mathbf{u}-\langle\mathbf{u},\mathbf{x}_{r}\rangle\mathbf{x}_{r}]\big{\|}_{2}$
$\leq\big{\|}\langle\mathbf{u},\mathbf{x}-\mathbf{x}_{r}\rangle\mathbf{x}\big{\|}_{2}+\big{\|}\langle\mathbf{u},\mathbf{x}_{r}\rangle(\mathbf{x}-\mathbf{x}_{r})\big{\|}_{2}\leq 2\|\mathbf{x}-\mathbf{x}_{r}\|_{2}\leq 2r,$
so the term in (A.41) is bounded by $2r$ uniformly for all $\mathbf{x}\in\mathcal{K}$. It remains to bound the term in (A.40).
By the triangle inequality and the observation that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$ gives $\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}$, the term in (A.40) is bounded by
(the term in (A.40)) $\leq\frac{1}{\sqrt{m}}\sup_{\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}}\big{\|}\Im\big{[}\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big{]}\big{\|}_{2}$
$=\frac{1}{\sqrt{m}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{\|}\Im\big{[}\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big{]}\big{\|}_{2}.$ (A.42)
We divide the $m$ measurements into two parts according to certain index sets. For some small enough $\eta^{\prime}>0$ to be chosen and $\mathbf{u}\in\mathbbm{R}^{n}$, we further introduce the index set
$\mathcal{I}_{\mathbf{u},\eta^{\prime}}=\big{\{}i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{u}|>\eta^{\prime}\big{\}}.$ (A.43)
We pause to establish a uniform bound on $|\mathcal{I}_{\mathbf{u},\eta^{\prime}}|$ for $\mathbf{u}\in\mathcal{K}_{(r)}$.
Bounding $|\mathcal{I}_{\mathbf{u},\eta^{\prime}}|$ uniformly over $\mathbf{u}\in\mathcal{K}_{(r)}$: For $\beta\in(0,\frac{1}{m})$ to be chosen, by Lemma 2.1, the event
$\sup_{\mathbf{u}\in\mathcal{K}_{(r)}}\big{|}\mathcal{I}_{\mathbf{u},\eta^{\prime}}\big{|}\leq\beta m$ (A.44)
holds with probability at least $1-4\exp(-c_{5}\beta m\log(\beta^{-1}))$, as long as
$\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\beta m}}+r\sqrt{\log(\beta^{-1})}\leq c_{6}\eta^{\prime}$ (A.45)
holds for some sufficiently small $c_{6}$. To see why this is sufficient, note that with the promised probability (A.45) implies
$\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq\beta m\end{subarray}}\left(\frac{1}{\beta m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}\leq\frac{\eta^{\prime}}{2},$
and this further implies (A.44). Our subsequent analysis is built upon the bound in (A.44).
For a specific $(\mathbf{x},\mathbf{x}_{r})\in\mathcal{K}\times\mathcal{N}_{r}$ we can define the index set for the “problematic measurements” as
$\mathcal{E}_{\mathbf{x}}:=\mathcal{J}_{\mathbf{x}_{r},\eta}\cup\mathcal{I}_{\mathbf{x}-\mathbf{x}_{r},\eta^{\prime}}$ (A.46)
and then bound the term in (A.42) as $I_{3}+I_{4}$, where
$I_{3}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}\Big{[}\Im\Big{(}\big{[}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big{]}\bm{\Phi}_{i}^{*}\mathbf{u}\Big{)}\Big{]}^{2}\right)^{1/2},$
$I_{4}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i\notin\mathcal{E}_{\mathbf{x}}}\Big{[}\Im\Big{(}\big{[}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big{]}\bm{\Phi}_{i}^{*}\mathbf{u}\Big{)}\Big{]}^{2}\right)^{1/2}.$
Bounding $I_{3}$: The issue for measurements in $\mathcal{E}_{\mathbf{x}}$ is the lack of a good bound on $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|$. Fortunately, these measurements are quite few: combining $\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|<\eta m$ and (A.44) gives
$|\mathcal{E}_{\mathbf{x}}|\leq|\mathcal{J}_{\mathbf{x}_{r},\eta}|+|\mathcal{I}_{\mathbf{x}-\mathbf{x}_{r},\eta^{\prime}}|\leq\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|+\sup_{\mathbf{u}\in\mathcal{K}_{(r)}}|\mathcal{I}_{\mathbf{u},\eta^{\prime}}|<(\eta+\beta)m.$
Combining with $|\Im([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u})|\leq 2|\bm{\Phi}_{i}^{*}\mathbf{u}|$, we proceed as follows:
$I_{3}$ $\leq 2\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{(}\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\Big{)}^{1/2}$
$\leq 2\sqrt{\eta+\beta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq(\eta+\beta)m\end{subarray}}\left(\frac{1}{(\eta+\beta)m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
$\leq C_{7}\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\frac{e}{\eta+\beta}}\right),$ (A.47)
where (A.47) follows from Lemma 2.1 and holds with probability at least $1-4\exp(-c_{8}(\eta+\beta)m\log(\frac{e}{\eta+\beta}))$.
Bounding $I_{4}$: For $i\notin\mathcal{E}_{\mathbf{x}}$ we have $|\bm{\Phi}_{i}^{*}\mathbf{x}_{r}|\geq\eta$ and $|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|<\eta^{\prime}$, and hence (2.11) implies
$\big{|}\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})\big{|}\leq\frac{2|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|}{\eta}\leq\frac{2\eta^{\prime}}{\eta}.$ (A.48)
Therefore, by $|\Im([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u})|\leq\frac{2\eta^{\prime}}{\eta}|\bm{\Phi}_{i}^{*}\mathbf{u}|$ we can bound $I_{4}$ as
$I_{4}\leq\frac{2\eta^{\prime}}{\eta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}\leq\frac{C_{9}\eta^{\prime}}{\eta},$ (A.49)
where the second inequality holds with probability at least $1-4\exp(-c_{10}m)$ if $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see (2.10). Combining (A.47) and (A.49) and recalling (A.42), we arrive at
[…]
Completing the Proof: Note that the terms in (A.40) and (A.41) are respectively bounded by (A.50) and $2r$, uniformly for all $\mathbf{x}\in\mathcal{K}$. Substituting them into (A.39)–(A.41), along with a supremum over $\mathbf{x}\in\mathcal{K}$, yields
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big{|}$
$\qquad+C_{11}\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\Big{(}\frac{e}{\eta+\beta}\Big{)}}+\frac{\eta^{\prime}}{\eta}\right)+2r.$
Combining with the bound in (A.38), taking $\beta=\Theta(\eta)$, and also summarizing the sample complexity and probability terms, we arrive at the following conclusion: Suppose
$m\geq C_{12}\left[\omega^{2}(\mathcal{U}_{1}^{\rm(N)})+\frac{\mathscr{H}(\mathcal{K},r)}{\eta}\right]\qquad\text{with large enough }C_{12},$ (A.51)
and
$\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\leq c_{13}\eta^{\prime}\qquad\text{with small enough }c_{13},$ (A.52)
the event
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{14}\left(\frac{\omega(\mathcal{U}^{\rm(N)}_{c})+\sqrt{\mathscr{H}(\mathcal{K},r)}}{\sqrt{m}}+\sqrt{\eta\log(\eta^{-1})}+\frac{\eta^{\prime}}{\eta}+r\right)$
holds with probability at least $1-C_{15}\exp(-c_{16}(\omega^{2}(\mathcal{U}^{\rm(N)}_{c})+\mathscr{H}(\mathcal{K},r)))-C_{17}\exp(-c_{18}\eta m)$. We mention that the condition (A.51) is needed for ensuring $\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|<\eta m$ and the second inequality of (A.49), and the condition (A.52) is needed in (A.44).
Further Simplification: We now take the tightest choice for $\eta^{\prime}$ that satisfies the required (A.52), namely $\eta^{\prime}=\Theta\big{(}\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\big{)}.$ We further set $\eta=\hat{\eta}^{2}$ and $r=\eta^{3/2}=\hat{\eta}^{3}$, and then the above statement simplifies to the following: if
$m\geq C_{19}\left(\omega^{2}(\mathcal{U}^{\rm(N)}_{c})+\frac{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}{\hat{\eta}^{2}}\right),$
then the event
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{20}\left(\frac{\omega(\mathcal{U}_{c}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}+\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}+\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})}\right)$ (A.53)
holds with probability at least $1-C_{21}\exp(-c_{22}(\omega^{2}(\mathcal{U}_{c}^{\rm(N)})+\mathscr{H}(\mathcal{K},\hat{\eta}^{3})))-C_{23}\exp(-c_{24}\hat{\eta}^{2}m)$. Under the sample complexity in (A.37) but with $\eta$ replaced by $\hat{\eta}$, it is easy to see
$\frac{\omega(\mathcal{U}_{c}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}+\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}=O(\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})}).$
Further renaming $\hat{\eta}$ to $\eta$ completes the proof. ∎
A.7 Proof of Lemma A.1
Proof.
We are ready to substitute the bounds for the parallel term and the orthogonal term into (A.5) to establish Lemma A.1. Recall from (A.12) and (A.13) that
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$
holds with probability $1-4\exp(-c_{2}m)$, and that
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|{f}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq C_{3}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$
holds with probability at least $1-4\exp(-c_{4}m)$ due to (A.19). We now observe that the stated sample complexity (A.1) implies (A.24) and (A.37), and hence we can apply Lemma A.6 and Lemma A.7 to obtain
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|{f}_{1}^{\|}(\mathbf{x},\mathbf{u})|=O\big{(}\eta\sqrt{\log(\eta^{-1})}\big{)},$
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big{(}\eta\sqrt{\log(\eta^{-1})}\big{)},$
that hold with the promised probability. Therefore, we arrive at
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big{(}\eta\sqrt{\log(\eta^{-1})}\big{)}.$
In view of (A.3) and (A.5), we derive $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}f(\mathbf{x},\mathbf{u},\mathbf{w})=O(\eta\sqrt{\log(\eta^{-1})})$, which is just the desired RIP with distortion $\delta_{\eta}$. The proof is complete. ∎
Appendix B Deferred Proofs (Proposition 3.1 & Lemma 4.1)
B.1 Proof of Proposition 3.1
Proof.
Note that $\varepsilon\geq(1+\delta_{1})\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=(1+\delta_{1})\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}$. As in the proof of Theorem 3.6, we can assume $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$ with the promised probability. Then we claim that all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{4}{5}\delta_{1}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2})$ satisfy the constraint $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}-\mathbf{e}_{1}\|_{2}\leq\varepsilon$. To see this, if $\mathbf{u}\in\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{4}{5}\delta_{1}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2})$, then we have
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}-\mathbf{e}_{1}\|_{2}\leq\|\mathbf{A}_{\breve{\mathbf{z}}}(\mathbf{u}-\mathbf{x}^{\star})\|_{2}+\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}\leq\sqrt{\frac{4}{3}}\frac{4\delta_{1}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}}{5}+\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}\leq\varepsilon.$
Next, we lower bound $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}$. We start with $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{2}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{1}{2}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}$, where in the inequality we use $\|\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{2}}\geq\frac{1}{2}$ that holds with the promised probability due to the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$. We denote the index set for the $\zeta_{0}m$ measurements with the largest $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ by $I_{\zeta_{0}}$. By recalling (1.7) and that $\bm{\zeta}$ changes the measurements in $I_{\zeta_{0}}$ from $z_{i}$ to $\mathbf{i}z_{i}$, we have
$\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}$ $\geq\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}=\frac{1}{\sqrt{m}}\left(\sum_{i\in I_{\zeta_{0}}}\big{[}\Im\big{(}\overline{(\mathbf{i}-1)\bm{\Phi}_{i}^{*}\mathbf{x}}\cdot\bm{\Phi}_{i}^{*}\mathbf{x}\big{)}\big{]}^{2}\right)^{1/2}$
$=\frac{1}{\sqrt{m}}\left(\sum_{i\in I_{\zeta_{0}}}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\right)^{1/2}\geq\frac{1}{\sqrt{m}}\cdot\frac{1}{\sqrt{\zeta_{0}m}}\sum_{i\in I_{\zeta_{0}}}|\bm{\Phi}_{i}^{*}\mathbf{x}|,$ (B.1)
where the last step follows from Cauchy-Schwarz inequality.
We further let $I_{\zeta_{0}}^{\prime}$ be the index set for the $\zeta_{0}m$ measurements with the largest $|(\bm{\Phi}^{\Re}_{i})^{\top}\mathbf{x}|$. Since $I_{\zeta_{0}}$ corresponds to the $\zeta_{0}m$ measurements with the largest $|\bm{\Phi}_{i}^{*}\mathbf{x}|$, continuing from (B.1) we have
$\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{1}{\sqrt{m}}\cdot\frac{1}{\sqrt{\zeta_{0}m}}\sum_{i\in I_{\zeta_{0}}^{\prime}}\big{|}(\bm{\Phi}_{i}^{\Re})^{\top}\mathbf{x}\big{|}.$
Now let us construct a set $\mathcal{V}:=\big{\{}\frac{1}{\sqrt{\zeta_{0}m}},-\frac{1}{\sqrt{\zeta_{0}m}},0\big{\}}^{m}\cap\Sigma^{m}_{\zeta_{0}m}$ whose elements are $m$-dimensional, $\zeta_{0}m$-sparse {±$\{\frac{\pm 1}{\sqrt{\zeta_{0}m}},0\}$-valued vectors. Combining with the definition of $I_{\zeta_{0}}^{\prime}$ we can write
$\frac{1}{\sqrt{\zeta_{0}m}}\sum_{i\in I_{\zeta_{0}}^{\prime}}\big{|}(\bm{\Phi}_{i}^{\Re})^{\top}\mathbf{x}\big{|}=\max_{\mathbf{v}\in\mathcal{V}}\mathbf{v}^{\top}\bm{\Phi}^{\Re}\mathbf{x}\stackrel{{\scriptstyle d}}{{=}}\max_{\mathbf{v}\in\mathcal{V}}\mathbf{g}^{\top}\mathbf{v},$
where we observe that $\bm{\Phi}^{\Re}\mathbf{x}\sim\mathcal{N}(0,\mathbf{I}_{m})$ for a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and further let $\mathbf{g}\sim\mathcal{N}(0,\mathbf{I}_{n})$ and assert that $\mathbf{v}^{\top}\bm{\Phi}^{\Re}\mathbf{x}$ and $\mathbf{g}^{\top}\mathbf{v}$ have the same distribution. Therefore, by Gaussian concentration (e.g., [49, Thm. 5.2.2]) we can show that $\max_{\mathbf{v}\in\mathcal{V}}\mathbf{v}^{\top}\bm{\Phi}^{\Re}\mathbf{x}\geq\frac{1}{2}\omega(\mathcal{V})$ with probability at least $1-2\exp(-c_{1}\omega^{2}(\mathcal{V}))$.
We now seek lower bound on $\omega(\mathcal{V})$. By the Sparse Varshamov-Gilbert construction (e.g., [45, Lem. 4.14]) there exist $(1+\frac{1}{2\zeta_{0}})^{\frac{\zeta_{0}m}{8}}$ distinct points contained in $\mathcal{V}$ with mutual $\ell_{2}$ distances greater than $\frac{1}{\sqrt{2}}$. This implies
$\mathscr{H}\Big{(}\mathcal{V},\frac{1}{2\sqrt{2}}\Big{)}=\log\mathscr{N}\Big{(}\mathcal{V},\frac{1}{2\sqrt{2}}\Big{)}\geq\frac{\zeta_{0}m}{8}\log\Big{(}1+\frac{1}{2\zeta_{0}}\Big{)},$
and Sudakov’s inequality (2.4) further gives $\omega^{2}(\mathcal{V})\geq c_{2}\zeta_{0}m\log(\frac{e}{\zeta_{0}})$. Combining these pieces, with the promised probability we have $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq c_{3}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$. Recall that all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{4\delta_{1}}{5}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2})$ satisfy the constraint $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}-\mathbf{e}_{1}\|_{2}\leq\varepsilon$. Since $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{c_{3}}{2}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$, all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$ satisfy the constraint of (1.8).
To conclude the proof, it remains to show $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\geq\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$. To do so, we proceed under the assumption
$\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})},$
and we seek to show that equality must hold (i.e., $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}=\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$). We first show that $\hat{\mathbf{x}}\in\Sigma^{n}_{s}$. In fact, if $\hat{\mathbf{x}}\notin\Sigma^{n}_{s}$, we construct $\hat{\mathbf{x}}^{\prime}$ from $\hat{\mathbf{x}}$ by setting all entries not in $\operatorname{supp}(\mathbf{x}^{\star})$ to zero; this gives $\|\hat{\mathbf{x}}^{\prime}\|_{1}<\|\hat{\mathbf{x}}\|_{1}$, since at least one nonzero entry becomes zero. Moreover,
$\|\hat{\mathbf{x}}^{\prime}-\mathbf{x}^{\star}\|_{2}\leq\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})},$
and hence $\hat{\mathbf{x}}^{\prime}$ satisfies the constraint of (1.8). This contradicts the optimality of $\hat{\mathbf{x}}$ to (1.8). Therefore, we obtain $\hat{\mathbf{x}}\in\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$. Because $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$ is a subset of the feasible domain of (1.8), we have $\hat{\mathbf{x}}=\hat{\mathbf{x}}_{c}$ where
$\hat{\mathbf{x}}_{c}=\mathrm{arg}\min~{}\|\mathbf{u}\|_{1},\qquad\text{subject to }\mathbf{u}\in\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}\Big{(}\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})}\Big{)}.$
Under small enough $\zeta_{0}$, we use $\|\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}$ to obtain $\frac{2c_{3}\delta_{1}\sqrt{\zeta_{0}\log(e/\zeta_{0})}}{5}<\frac{1}{2}\leq\|\mathbf{x}^{\star}\|_{2}$. Then, it is not hard to observe that $\hat{\mathbf{x}}_{c}$ must live in the boundary of $\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$. Hence, with the promised probability, we have $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}=\|\hat{\mathbf{x}}_{c}-\mathbf{x}^{\star}\|_{2}=\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$. The result follows. ∎
B.2 Proof of Lemma 4.1
We have the following refined statement.
Lemma B.1.
Suppose the entries of $\bm{\Phi}$ are drawn i.i.d. from $\mathcal{N}(0,1)+\mathcal{N}(0,1)\mathbf{i}$. Given some small enough $\eta\in[\frac{C_{1}}{m},1]$ and some $\mathcal{K}\subset\mathbb{S}^{n-1}$, we let $r=\frac{c_{1}\eta}{(\log(\eta^{-1}))^{1/2}}$ with sufficiently small $c_{1}$. If for sufficiently large $C_{2}$ we have
$m\geq C_{2}\left(\frac{\mathscr{H}(\mathcal{K},r)}{\eta}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{3}}\right),$
then with probability at least $1-3\exp(-c_{3}\eta m)$ we have $\sup_{\mathbf{x}\in\mathcal{K}}~{}|\mathcal{J}_{\mathbf{x},\eta}|\leq\eta m.$
Before proving this, we note that it immediately leads to Lemma 4.1: by $\omega^{2}(\mathcal{K}_{(r)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})\leq 4\omega^{2}(\mathcal{K})$ [49, Sec. 7.5.1] and $\mathscr{H}(\mathcal{K},r)\leq C_{1}\frac{\omega^{2}(\mathcal{K})}{r^{2}}=\Theta(\frac{\log(\eta^{-1})\omega^{2}(\mathcal{K})}{\eta^{2}})$ from (2.4), we find that $m=\Omega(\eta^{-3}\log(\eta^{-1})\omega^{2}(\mathcal{K}))$ in Lemma 4.1 suffices to imply the sample complexity in Lemma B.1.
Proof of Lemma B.1.
We use a covering approach to bound $\sup_{\mathbf{x}\in\mathcal{K}}|\mathcal{J}_{\mathbf{x},\eta}|=\sup_{\mathbf{x}\in\mathcal{K}}\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\big{)}$. We let $\mathcal{N}_{r}$ be a minimal $r$-net of $\mathcal{K}$ with $\log|\mathcal{N}_{r}|=\mathscr{H}(\mathcal{K},r)$, then for any $\mathbf{x}\in\mathcal{K}$ we let $\mathbf{x}^{\prime}=\mathrm{arg}\min_{\mathbf{u}\in\mathcal{N}_{r}}\|\mathbf{u}-\mathbf{x}\|_{2}$. Here, $\mathbf{x}^{\prime}$ depends on $\mathbf{x}$, but we drop such dependence to avoid cumbersome notation. Note that we have $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}\leq r$ and $\mathbf{x}-\mathbf{x}^{\prime}\in\mathcal{K}_{(r)}$. By the triangle inequality, we have
$\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\big{)}\leq\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}\mathbf{x}^{\prime}|-|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}^{\prime})|\leq\eta\big{)}$
$\leq\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}\mathbf{x}^{\prime}|\leq 1.1\eta\big{)}+\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}^{\prime})|>0.1\eta\big{)},$
which implies
$\sup_{\mathbf{x}\in\mathcal{K}}\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\big{)}\leq\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq 1.1\eta\big{)}+\sup_{\mathbf{u}\in\mathcal{K}_{(r)}}\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}\mathbf{u}|>0.1\eta\big{)}.$ (B.2)
We first bound $\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq 1.1\eta\big{)}=\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},1.1\eta}|$. For fixed $\mathbf{x}\in\mathbb{S}^{n-1}$, we have
$\mathbbm{P}(|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq 1.1\eta)\leq\mathbbm{P}(|\mathcal{N}(0,1)|\leq 1.1\eta)\leq 1.1\sqrt{\frac{2}{\pi}}\eta\leq 0.9\eta,$
and hence Chernoff bound gives $|\mathcal{J}_{\mathbf{x},1.1\eta}|\leq 0.95\eta m$ with probability at least $1-\exp(-c_{1}\eta m)$, where $c_{1}$ is some absolute constant. Therefore, when $m\geq\frac{C_{2}\mathscr{H}(\mathcal{K},r)}{\eta}$ with large enough $C_{2}$, we can take a union bound and obtain $\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},1.1\eta}|\leq 0.95\eta m$ with probability at least $1-\exp(-\frac{c_{1}\eta m}{2})$.
All that remains is to show
$\sup_{\mathbf{u}\in\mathcal{K}_{(r)}}\sum_{i=1}^{m}\mathbbm{1}\big{(}|\bm{\Phi}_{i}^{*}\mathbf{u}|>0.1\eta\big{)}\leq 0.05\eta m.$ (B.3)
For notational convenience, suppose that $0.05\eta m$ is a positive integer (we can round otherwise). We observe that a sufficient condition for (B.3) is
$\sup_{\mathbf{u}\in\mathcal{K}_{(r)}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=0.05\eta m\end{subarray}}\Big{(}\frac{1}{0.05\eta m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\Big{)}^{1/2}\leq 0.05\eta.$ (B.4)
Thus, it suffices to show (B.4). By Lemma 2.1, with probability at least $1-2\exp(-c_{2}\eta m\log(\eta^{-1}))$ it suffices to ensure
$\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\leq c_{3}\eta$
for some small enough absolute constant $c_{3}$, and hence it is sufficient to have $m\geq C_{4}\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{3}}$ and $r=\frac{c_{5}\eta}{(\log(\eta^{-1}))^{1/2}}$, where $C_{4}$ is sufficient large and $c_{5}$ is small enough. These assumptions are made in our statement, and hence the claim follows. ∎

### v3

@@
 In the noiseless case, this translates into exact reconstruction of sparse signals and accurate estimate of approximately sparse signals.
-As an example, if $\mathbf{A}$ satisfies RIP over the cone of sparse vectors and $\varepsilon\geq\|\bm{\epsilon}\|_{2}$, then basis pursuit $\Delta(\mathbf{A};\mathbf{y};\varepsilon)$
+As an example, if $\mathbf{A}$ satisfies RIP over the cone of sparse vectors and $\varepsilon\geq\|\bm{\epsilon}\|_{2}$, then basis pursuit
 $\hat{\mathbf{x}}=\mathrm{arg}\min\|\mathbf{u}\|_{1},\quad\text{subject to }\|\mathbf{A}\mathbf{u}-\mathbf{y}\|_{2}\leq\varepsilon$ (1.1)
@@
 for some absolute constants $C_{1},C_{2}$, where $\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s}):=\min_{\mathbf{u}\in\Sigma^{n}_{s}}\|\mathbf{x}-\mathbf{u}\|_{1}$ denotes the $\ell_{1}$ distance of $\mathbf{x}$ to $\Sigma^{n}_{s}.$ More details are given in Section 2.
-The focus of the present paper is on the nonlinear compressed sensing model of phase-only compressed sensing (PO-CS), which concerns the reconstruction of sparse signals in the standard sphere $\mathbb{S}^{n-1}=\{\mathbf{x}\in\mathbb{R}^{n}:\|\mathbf{x}\|_{2}=1\}$ from
+We write $\Delta(\mathbf{A};\mathbf{y};\varepsilon)$ for the estimator $\hat{\mathbf{x}}$ in (1.1) to emphasize the required inputs $(\mathbf{A},\mathbf{y},\varepsilon).$
+The focus of the present paper is on the nonlinear compressed sensing model of phase-only compressed sensing (PO-CS), which concerns the reconstruction of sparse signals in the standard sphere $\mathbb{S}^{n-1}=\{\mathbf{x}\in\mathbb{R}^{n}:\|\mathbf{x}\|_{2}=1\}$ from [^1] We focus on real-valued signals in this paper, but point out that our results generalize to complex-valued signals by slightly more work (see [10]).
 $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})=[\operatorname{sign}(\bm{\Phi}_{1}^{*}\mathbf{x}),\cdots,\operatorname{sign}(\bm{\Phi}_{m}^{*}\mathbf{x})]^{\top}$ (1.3)
 under a complex sensing matrix $\bm{\Phi}=[\bm{\Phi}_{1},\cdots,\bm{\Phi}_{m}]^{*}\in\mathbb{C}^{m\times n}$, with the phase function being given by $\operatorname{sign}(c)=\frac{c}{|c|}$ for $c\in\mathbb{C}\setminus\{0\}$ and $\operatorname{sign}(0)=1$ by convention.
-We assume that we observe $\breve{\mathbf{z}}\in\mathbb{C}^{m}$ that equaling $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ in the noiseless case or a perturbed version of $\mathbf{z}$ in the noisy case.
-We note that phase-only reconstruction of unstructured signals has been well studied [26, 38, 9, 25, 18, 32, 48, 33].
-PO-CS was initially considered as a natural extension of 1-bit compressed sensing [^1] This concerns the recovery of $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\mathbf{y}=\operatorname{sign}(\mathbf{A}\mathbf{x})$ with real sensing matrix $\mathbf{A}\in\mathbb{R}^{m\times n}.$ [3, 4] and was recently revisited in [19, 10, 27, 8].
+Let $\breve{\mathbf{z}}\in\mathbb{C}^{m}$ be our observation vector, which equals $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ in the noiseless case and is a perturbed version of $\mathbf{z}$ in the noisy case.
+[^2] Throughout this paper, $\mathbf{x}$ denotes the underlying true signal, $\bm{\Phi}\in\mathbb{C}^{m\times n}$ denotes the complex sensing matrix, $\mathbf{z}$ is reserved for the noiseless phase-only observations $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$, and $\breve{\mathbf{z}}$ is our observations and a possibly noisy version of $\mathbf{z}$.
+Therefore, $\mathbf{z}$ and $\breve{\mathbf{z}}$ implicitly depend on $\mathbf{x}$.
+Such a model may find applications in sensing systems subject to severe dynamic-range limitations or unknown gain fluctuations, and may also be useful under extreme quantization or multiplicative noise, since phase lies on a compact unit circle and is invariant under unknown scaling.
+We note that phase-only reconstruction of unstructured signals was studied extensively several decades ago [25, 39, 9, 26, 18, 33, 49, 34].
+PO-CS was initially considered as a natural extension of 1-bit compressed sensing [^3] This concerns the recovery of $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\mathbf{y}=\operatorname{sign}(\mathbf{A}\mathbf{x})$ with real sensing matrix $\mathbf{A}\in\mathbb{R}^{m\times n}.$ Note that the phase-only measurements (1.3) take continuous values (on the complex unit circle) and are therefore more informative than 1-bit measurements $\mathbf{y}=\operatorname{sign}(\mathbf{A}\mathbf{x})\in\{-1,1\}^{m}$.
+This distinction leads to a fundamental difference between the two problems: while 1-bit compressed sensing admits a strictly positive information-theoretic lower bound on the reconstruction error [28], exact recovery can be achieved in phase-only compressed sensing.
+[3, 4] and was recently revisited in [19, 10, 27, 8].
 Exact reconstruction was observed experimentally in [4] and theoretically proved very recently by Jacques and Feuillen [27] who proposed to recast PO-CS as a linear compressed sensing problem.
@@
 (The value of $\kappa$ here is non-essential but chosen to facilitate subsequent analysis.) Combining (1.4) and (1.5), we arrive at the linear compressed sensing problem
-[…]
+$\text{find sparse}~\mathbf{u},~~\text{such that }\begin{bmatrix}\frac{1}{\kappa m}\Re(\breve{\mathbf{z}}^{*}\bm{\Phi})\\ \frac{1}{\sqrt{m}}\Im(\operatorname{diag}(\breve{\mathbf{z}}^{*})\bm{\Phi})\end{bmatrix}\mathbf{u}=\mathbf{e}_{1}.$ (1.6)
 In the noisy case with $\breve{\mathbf{z}}\neq\mathbf{z}$, the linear equations in (1.6) become inexact and we instead encounter a noisy linear compressed sensing problem.
@@
 In fact, the new sensing matrix $\mathbf{A}_{\breve{\mathbf{z}}}$ depends on $\mathbf{x}$ through $\breve{\mathbf{z}}$, so proving the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$ for a fixed $\mathbf{x}$ only implies the nonuniform recovery of this specific $\mathbf{x}$.
-In [10], the authors used a covering argument to show that the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}:=\Sigma^{n}_{s}\cap\mathbb{S}^{n-1}\}$ simultaneously obey RIP under a near-optimal number of measurements.
+In [10], the authors used a covering argument to show that the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\Sigma^{n,*}_{s}:=\Sigma^{n}_{s}\cap\mathbb{S}^{n-1}\}$ simultaneously obey RIP under a near-optimal number of measurements.
 This leads to a uniform exact reconstruction guarantee over $\Sigma^{n,*}_{s},$ but provides no guarantee for $\mathbf{x}\notin\Sigma^{n,*}_{s}$.
@@
 $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{10\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}},\qquad\forall\mathbf{x}\in\mathbb{S}^{n-1}.$
-The main ingredient is to show the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$, where $\mathbb{B}_{1}^{n}(\sqrt{2s})=\{\mathbf{u}\in\mathbb{R}^{n}:\|\mathbf{u}\|_{1}\leq\sqrt{2s}\}$ is the scaled-$\ell_{1}$ ball, simultaneously satisfy RIP.
-Note that [10] proved that the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy RIP through a covering argument, yet their arguments are not sufficient to prove the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$.
+The main ingredient is to show the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$, where $\mathbb{B}_{1}^{n}(\sqrt{2s})=\{\mathbf{u}\in\mathbb{R}^{n}:\|\mathbf{u}\|_{1}\leq\sqrt{2s}\}$ is the scaled-$\ell_{1}$ ball, simultaneously satisfy RIP.
+Note that [10] proved that the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy RIP through a covering argument, yet their arguments are not sufficient to prove the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$.
 The main issue is that $\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}$ is essentially larger than $\Sigma^{n,*}_{s}$ in terms of metric entropy (or covering number) under an $o(1)$ covering radius.
@@
 More generally, we establish the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ over some cone $\mathcal{U}$ for an arbitrary set $\mathcal{K}\subset\mathbb{S}^{n-1}$.
-Here is an informal version.
+Here is an informal version, where $\omega(\mathcal{U})$ denotes the Gaussian width of a set $\mathcal{U}$ (see Section 2.2).
 Theorem (Informal).
-Given a cone $\mathcal{U}$ in $\mathbb{R}^{n}$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if $m\geq C_{1}\big{(}\omega^{2}(\mathcal{U}\cap\mathbb{S}^{n-1})+\omega^{2}(\mathcal{K})\big{)}$ with sufficiently large $C_{1}$, then with high probability on the complex Gaussian $\bm{\Phi}$, the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ satisfy RIP over $\mathcal{U}$ with small enough distortion.
+Given a cone $\mathcal{U}$ in $\mathbb{R}^{n}$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if
+$m\geq C_{1}\big(\omega^{2}(\mathcal{U}\cap\mathbb{S}^{n-1})+\omega^{2}(\mathcal{K})\big)$
+with sufficiently large $C_{1}$, then with high probability on the complex Gaussian $\bm{\Phi}$, the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\mathbf{A}\mathbf{x}),\,\mathbf{x}\in\mathcal{K}\}$ satisfy RIP over $\mathcal{U}$ with small enough distortion.
 Our second contribution is to understand the robustness of $\mathbf{x}^{\sharp}$ to different patterns of noise and corruption.
@@
 We expect that $\tilde{O}(\sqrt{\zeta_{0}})$ is tight for the specific estimator $\mathbf{x}^{\sharp}$, which we support by providing a partial analysis and numerical evidence.
-However, for general estimators such dependence is suboptimal and can be improved to “zero,” as in this regime we can still exactly recover $\mathbf{x}$, akin to existing studies of corrupted linear systems and corrupted sensing [21, 7, 24].
+However, for general estimators such dependence is suboptimal and can be improved to “zero,” as in this regime we can still exactly recover $\mathbf{x}$.
 In particular, we propose to reformulate PO-CS under sparse corruption to a noiseless linear compressed sensing problem with an extended new sensing matrix.
@@
 Bounded Dense Noise $\bm{\tau}$ $\|\bm{\tau}\|_{\infty}\leq\tau_{0}$ $O(\tau_{0})$ Thms. 3.3 & 3.4 Yes, up to log Thm. 3.5 Yes, up to log Thm. 3.5 Figs. 12
-Sparse Corruption $\bm{\zeta}$ $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ $O\Big{(}\sqrt{\zeta_{0}\log(1/\zeta_{0})}\Big{)}$ Thm. 3.6 Partially Yes Prop. 3.1 No, 00 is optimal Thm. 3.8 Fig. 3
+Sparse Corruption $\bm{\zeta}$ $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ $O\Big(\sqrt{\zeta_{0}\log(1/\zeta_{0})}\Big)$ Thm. 3.6 Partially Yes Prop. 3.1 No, 0 is optimal Thm. 3.8 Fig. 3
 Table 1: A summary of our results on robustness.
@@
 Then with high probability over the complex Gaussian $\bm{\Phi}$, we have
-$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq C_{1}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}+C_{2}\tau_{0}+C_{3}\sqrt{\zeta_{0}\log(1/\zeta_{0})}+C_{4}\sqrt{\frac{s\log(\frac{en}{s})}{m}},\qquad\forall\mathbf{x}\in\mathbb{S}^{n-1}.$ (1.9)
+$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq C_{1}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}+C_{2}\tau_{0}+C_{3}\sqrt{\zeta_{0}\log(1/\zeta_{0})}+C_{4}\sqrt{\frac{s\log(\frac{en}{s})}{m}}\mathbbm{1}(\zeta_{0}>0),\quad\forall\mathbf{x}\in\mathbb{S}^{n-1}.$ (1.9)
 To our knowledge, this type of result is novel in nonlinear sensing problems, for which most existing guarantees provide the same error rate to all signals of interest and hence are not instance optimal (e.g., [51, 11, 42, 43, 41, 45, 23]).
@@
 The standard Euclidean sphere, the $\ell_{2}$-ball and the $\ell_{1}$-ball in $\mathbb{R}^{n}$ are denoted by $\mathbb{S}^{n-1}$, $\mathbb{B}_{2}^{n}$ and $\mathbb{B}_{1}^{n}$, respectively.
-Given $\mathcal{K},\mathcal{K}^{\prime}\subset\mathbb{R}^{n}$ and some $\lambda\in\mathbb{R}$, we let $\mathcal{K}+\lambda\mathcal{K}^{\prime}:=\{\mathbf{u}+\lambda\mathbf{v}:\mathbf{u}\in\mathcal{K},\mathbf{v}\in\mathcal{K}^{\prime}\}$, $\mathcal{K}_{(\lambda)}=(\mathcal{K}-\mathcal{K})\cap(\lambda\mathbb{B}_{2}^{n})$ and $\mathcal{K}^{\rm(N)}=\mathcal{K}\cap\mathbb{S}^{n-1}$.
+Given $\mathcal{K},\mathcal{K}^{\prime}\subset\mathbb{R}^{n}$ and some $\lambda\in\mathbb{R}$, we let $\mathcal{K}+\lambda\mathcal{K}^{\prime}:=\{\mathbf{u}+\lambda\mathbf{v}:\mathbf{u}\in\mathcal{K},\mathbf{v}\in\mathcal{K}^{\prime}\}$, $\mathcal{K}_{(\lambda)}=(\mathcal{K}-\mathcal{K})\cap(\lambda\mathbb{B}_{2}^{n})$ and $\mathcal{K}^{\mathbb{S}}=\mathcal{K}\cap\mathbb{S}^{n-1}$.
 We write $\mathbb{B}_{2}^{n}(\mathbf{u};r):=\mathbf{u}+r\mathbb{B}_{2}^{n}$, $\mathbb{B}_{2}^{n}(r):=r\mathbb{B}_{2}^{n}$ and $\mathbb{B}_{1}^{n}(r):=r\mathbb{B}_{1}^{n}$.
 We also define $\operatorname{rad}(\mathcal{K})=\sup_{\mathbf{u}\in\mathcal{K}}\|\mathbf{u}\|_{2}$.
-Recall that $\Sigma^{n,*}_{s}:=(\Sigma^{n}_{s})^{\rm(N)}$ is the set of all $s$-sparse signals in $\mathbb{S}^{n-1}$.
+Recall that $\Sigma^{n,*}_{s}:=(\Sigma^{n}_{s})^{\mathbb{S}}$ is the set of all $s$-sparse signals in $\mathbb{S}^{n-1}$.
 We refer to complex numbers with absolute value $1$ as the phase-only values.
@@
 Further given $\mathbf{v}=[v_{i}]\in\mathbb{C}^{n}$, we work with the inner product $\langle\mathbf{u},\mathbf{v}\rangle=\mathbf{u}^{*}\mathbf{v}=\sum_{i=1}^{n}\overline{u_{i}}v_{i}$ and the Hadamard product $\mathbf{u}\odot\mathbf{v}=(u_{1}v_{1},u_{2}v_{2},\cdots,u_{n}v_{n})^{\top}$.
-For a complex matrix $\mathbf{A}=\mathbf{B}+\mathbf{C}\mathbf{i}$ with $\mathbf{i}$ reserved for the complex unit, we will use $\mathbf{A}^{\Re}$ or $\Re(\mathbf{A})$ to denote its real part $\mathbf{B}$, and $\mathbf{A}^{\Im}$ or $\Im(\mathbf{A})$ to denote its imaginary part $\mathbf{C}$.
-For a random variable $X$ we define the sub-Gaussian norm as $\|X\|_{\psi_{2}}=\inf\{t>0:\mathbbm{E}\exp(X^{2}/t^{2})<2\}$ and the sub-exponential norm as $\|X\|_{\psi_{1}}=\inf\{t>0:\mathbbm{E}\exp(|X|/t)<2\}$.
+For a complex matrix $\mathbf{A}=\mathbf{B}+\mathbf{C}\mathbf{i}$ with $\mathbf{i}$ reserved for the imaginary unit, we will use $\mathbf{A}^{\Re}$ or $\Re(\mathbf{A})$ to denote its real part $\mathbf{B}$, and $\mathbf{A}^{\Im}$ or $\Im(\mathbf{A})$ to denote its imaginary part $\mathbf{C}$.
+For a random variable $X$ we define the sub-Gaussian norm as
+$\|X\|_{\psi_{2}}=\inf\{t>0:\mathbbm{E}\exp(X^{2}/t^{2})<2\}$
+and the sub-exponential norm as
+$\|X\|_{\psi_{1}}=\inf\{t>0:\mathbbm{E}\exp(|X|/t)<2\}.$
 For independent zero-mean sub-Gaussian variables $\{X_{i}\}_{i=1}^{N}$, there exists absolute constant $C$ such that
@@
 The sub-Gaussian norm of a random vector $\mathbf{X}\in\mathbb{R}^{n}$ is defined as $\|\mathbf{X}\|_{\psi_{2}}=\sup_{\mathbf{v}\in\mathbb{S}^{n-1}}\|\mathbf{v}^{\top}\mathbf{X}\|_{\psi_{2}}$.
-We refer readers to [49, Sec.
+We refer readers to [50, Sec.
 2] for details on these definitions and properties.
-We use $\{C,C_{1},C_{2},\cdots\}$ and $\{c,c_{1},c_{2},\cdots\}$ to denote absolute constants whose values may vary from line to line.
-For two positive quantities $I_{1}$ and $I_{2}$, we write $I_{1}=O(I_{2})$ if $I_{1}\leq CI_{2}$ holds for some $C$, and write $I_{1}=\Omega(I_{2})$ if $I_{1}\geq cI_{2}$ for some $c>0$.
-We write $I_{1}=\Theta(I_{2})$ if $I_{1}=O(I_{2})$ and $I_{1}=\Omega(I_{2})$ simultaneously hold.
-We also use $\tilde{O}(\cdot),\tilde{\Omega}(\cdot),\tilde{\Theta}(\cdot)$ as the less precise versions of these that hide log factors.
-We use $o(1)$ to generically denote quantity that tends to zero when $m,n,s\to\infty$.
+[…]
 We say $I_{1}$ is small enough (or sufficiently small) if $I_{1}\leq c_{1}$ for some suitably small constant $c_{1}$.
 Conversely, it is large enough (or sufficiently large) if $I_{1}\geq C_{1}$ for some suitably large constant $C_{1}$.
-We say $I_{1}$ is bounded away from 00 if $I_{1}\geq c_{1}$ for some $c_{1}>0$.
-Linear Compressed Sensing: Let $\mathcal{U}$ be a cone in $\mathbb{R}^{n}$, we say $\mathbf{A}$ satisfies RIP over $\mathcal{U}$ with distortion $\delta>0$, denoted by $\mathbf{A}\sim\operatorname{RIP}(\mathcal{U},\delta)$, if
+We say $I_{1}$ is bounded away from 0 if $I_{1}\geq c_{1}$ for some $c_{1}>0$.
+2.1 Linear Compressed Sensing
+Let $\mathcal{U}$ be a cone in $\mathbb{R}^{n}$, we say $\mathbf{A}$ satisfies RIP over $\mathcal{U}$ with distortion $\delta>0$, denoted by $\mathbf{A}\sim\operatorname{RIP}(\mathcal{U},\delta)$, if
 $(1-\delta)\|\mathbf{u}\|_{2}^{2}\leq\|\mathbf{A}\mathbf{u}\|_{2}^{2}\leq(1+\delta)\|\mathbf{u}\|_{2}^{2},\qquad\forall\mathbf{u}\in\mathcal{U}.$
 By homogeneity, this is equivalent to
-$\sqrt{1-\delta}\leq\|\mathbf{A}\mathbf{u}\|_{2}\leq\sqrt{1+\delta},\qquad\forall\mathbf{u}\in\mathcal{U}^{\rm(N)}=\mathcal{U}\cap\mathbb{S}^{n-1}.$
-To be specific, we will focus on sparse recovery and the corresponding program of $\ell_{1}$-norm minimization (1.1).
-Therefore, we typically utilize the RIP over $\Sigma^{n}_{ts}$ for certain $t>0$ to imply the instance optimality.
-We shall work with RIP over $\Sigma^{n}_{2s}$ and utilize the following.
-(Note that RIP$(\Sigma^{n}_{2s},\delta)$ with $\delta<\frac{\sqrt{2}}{2}$ works, and we set $\delta=\frac{1}{3}$ just for concreteness.)
+$\sqrt{1-\delta}\leq\|\mathbf{A}\mathbf{u}\|_{2}\leq\sqrt{1+\delta},\qquad\forall\mathbf{u}\in\mathcal{U}^{\mathbb{S}}=\mathcal{U}\cap\mathbb{S}^{n-1}.$
+[…]
 Proposition 2.1 (see Thm. 2.1 in [5]).
@@
 Indeed, the central goal of this work is to prove an analog for an efficient PO-CS algorithm.
-[…]
+To recover block sparse signals $\mathbf{x}=(\mathbf{x}_{1}^{\top},\mathbf{x}_{2}^{\top})^{\top}\in\Sigma^{n_{1}}_{s_{1}}\times\Sigma^{n_{2}}_{s_{2}}\subset\mathbb{R}^{n}$ from $\mathbf{y}=\mathbf{A}\mathbf{x}+\bm{\epsilon}$, we can use a constrained weighted $\ell_{1}$ minimization,
 $\hat{\mathbf{x}}=(\hat{\mathbf{x}}_{1}^{\top},\hat{\mathbf{x}}_{2}^{\top})^{\top}=\mathrm{arg}\min_{\mathbf{u}=(\mathbf{u}_{1}^{\top},\mathbf{u}_{2}^{\top})^{\top}}~\frac{\|\mathbf{u}_{1}\|_{1}}{\sqrt{s_{1}}}+\frac{\|\mathbf{u}_{2}\|_{1}}{\sqrt{s_{2}}},\quad\text{subject to }\|\mathbf{A}\mathbf{u}-\mathbf{y}\|_{2}\leq\varepsilon.$ (2.3)
@@
 $\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}\leq C_{1}\varepsilon+C_{2}\left(\frac{\sigma_{\ell_{1}}(\mathbf{x}_{1},\Sigma^{n_{1}}_{s_{1}})}{\sqrt{s_{1}}}+\frac{\sigma_{\ell_{1}}(\mathbf{x}_{2},\Sigma^{n_{2}}_{s_{2}})}{\sqrt{s_{2}}}\right),\qquad\forall\mathbf{x}\in\mathbb{R}^{n}.$
-Gaussian Width & Metric Entropy: We need to work with two natural quantities that characterize the complexity of a set $\mathcal{K}$.
+2.2 Gaussian Width & Metric Entropy
+We need to work with two natural quantities that characterize the complexity of a set $\mathcal{K}$.
 The first one is the Gaussian width $\omega(\mathcal{K}):=\mathbbm{E}\sup_{\mathbf{u}\in\mathcal{K}}~\mathbf{g}^{\top}\mathbf{u}$, where $\mathbf{g}\sim\mathcal{N}(0,\mathbf{I}_{n})$.
 The second one is the metric entropy $\mathscr{H}(\mathcal{K},r)=\log\mathscr{N}(\mathcal{K},r)$ where $\mathscr{N}(\mathcal{K},r)$ denotes the covering number of $\mathcal{K}$ under radius $r$, defined as the minimal number of radius-$r$ $\ell_{2}$-balls needed to cover $\mathcal{K}$.
-Metric entropy can be bounded in terms of the Gaussian width via Sudakov’s inequality [49, Coro.
+Metric entropy can be bounded in terms of the Gaussian width via Sudakov’s inequality [50, Coro.
 7.4.3],
@@
 for some absolute constant $C$.
-We also have Dudley’s inequality [49, Sec.
+We also have Dudley’s inequality [50, Sec.
 8.1] for the converse purpose.
 A notable difference is that the Gaussian width remains invariant after taking the convex hull, while the metric entropy under $o(1)$ covering radius could change significantly.
-For instance, the set $\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{B}_{2}^{n}$ (whose elements are known as the approximately $s$-sparse signals in $\mathbb{B}_{2}^{n}$) can be essentially viewed as the convex hull of $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}$ [41, Lem.
+For instance, the set $\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{B}_{2}^{n}$ (whose elements are known as the approximately $s$-sparse signals in $\mathbb{B}_{2}^{n}$) can be essentially viewed as the convex hull of $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}$ [43, Lem.
 3.1].
-Their Gaussian widths are of the same order [40, Sec.
+Their Gaussian widths are of the same order [42, Sec.
 2],
@@
 $\mathscr{H}(\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{B}_{2}^{n},r)\leq C_{2}r^{-2}s\log\Big(\frac{en}{s}\Big).$ (2.7)
-The dependence on $r$ in (2.7) is tight in some regime; see [41, Sec.
-3] and [44, Sec.
+The dependence on $r$ in (2.7) is tight in some regime; see [43, Sec.
+3] and [41, Sec.
 4.3.3].
 In particular, the cardinality of an $r$-net for $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}$ logarithmically increases with $r^{-1}$, while that of $\sqrt{s}\mathbb{B}_{1}^{n}$ increases quadratically with $r^{-1}$.
+2.3 Concentration Bound
 Next, we introduce some useful sub-Gaussian concentration bounds that capture the Gaussian width of the relevant set.
@@
 holds with probability at least $1-2\exp(-C_{2}k\log(\frac{em}{k}))$, where $C_{1}$ and $C_{2}$ are absolute constants only depending on $L$.
-The following upper bound is a simple consequence of the matrix deviation inequality [49, Sec.
+The following upper bound is a simple consequence of the matrix deviation inequality [50, Sec.
 9.1] and will be of recurring use: if $\mathbf{A}$ has i.i.d.
@@
 $\sup_{\mathbf{u}\in\mathcal{T}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}\leq\sup_{\mathbf{u}\in\mathcal{T}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\mathcal{T}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\leq C_{2}.$ (2.10)
-Perturbation of Complex Phase: Under the convention $\frac{x}{0}=\infty$ for any $x\geq 0$, it holds for any $a,b\in\mathbb{C}$ that (e.g., [10, Lem. 8])
+2.4 Perturbation of Complex Phase
+Under the convention $\frac{x}{0}=\infty$ for any $x\geq 0$, it holds for any $a,b\in\mathbb{C}$ that (e.g., [10, Lem. 8])
 $|\operatorname{sign}(a)-\operatorname{sign}(b)|\leq\min\left\{\frac{2|a-b|}{\max\{|a|,|b|\}},2\right\}.$ (2.11)
@@
 3 Main Results
-We present our results for $\mathbf{x}^{\sharp}$ obtained by normalizing $\hat{\mathbf{x}}=\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\epsilon)$ in (1.8).
+We present our results for $\mathbf{x}^{\sharp}$ obtained by normalizing $\hat{\mathbf{x}}=\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\varepsilon)$ in (1.8).
 Throughout the paper, we assume complex Gaussian $\bm{\Phi}$ with i.i.d.
@@
 as it is easy to check $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$.
-(This means that $\mathbf{x}^{\star}$ is the point that satisfies the linear measurements in (1.8) in a noiseless case.) With $\kappa=\sqrt{\frac{\pi}{2}}$, $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ sharply concentrates about $1$ due to sub-Gaussian tail bounds, as we have $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=O(\frac{1}{\sqrt{m}})$ by (2.1).
-[^4] By (2.1), we have $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=\|\frac{1}{m}\sum_{i=1}^{m}(\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1)\|_{\psi_{2}}\leq\frac{C}{m}(\sum_{i=1}^{m}\|\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1\|_{\psi_{2}}^{2})^{1/2}\leq\frac{C_{1}}{\sqrt{m}}$.
-As a consequence, we have the sub-Gaussian tail bound $\mathbbm{P}(|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|\geq t)\leq 2\exp(-c_{2}mt)$ for any $t\geq 0$, hence $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ is small enough with probability at least $1-2\exp(-c_{3}m)$.
-We will use this observation in subsequent analysis.
-Hence, $\mathbf{x}^{\star}$ is in general very close to $\mathbf{x}$.
+(This means that $\mathbf{x}^{\star}$ is the point that satisfies the linear measurements in (1.8) in a noiseless case.) With $\kappa=\sqrt{\frac{\pi}{2}}$, $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ sharply concentrates about $1$: by (2.1), we have
+$\Big\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\Big\|_{\psi_{2}}=\Big\|\frac{1}{m}\sum_{i=1}^{m}(\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1)\Big\|_{\psi_{2}}\leq\frac{C}{m}\Big(\sum_{i=1}^{m}\|\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1\|_{\psi_{2}}^{2}\Big)^{1/2}\leq\frac{C_{1}}{\sqrt{m}},$
+which gives the sub-Gaussian tail bound
+$\mathbbm{P}\Big(\big|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\big|\geq t\Big)\leq 2\exp(-c_{2}mt)\,,\quad\forall t\geq 0.$ (3.2)
+Hence, $\mathbf{x}^{\star}$ is in general very close to $\mathbf{x}$. We will provide a uniform version of this fact in the subsequent Proposition 4.1.
+Before proceeding, we provide a deterministic proof framework of our upper bounds. It is adapted from the original work [27].
+Lemma 3.1 (Deterministic error bound).
+Fix $\mathbf{x}\in\mathbb{S}^{n-1}$, $\bm{\Phi}\in\mathbb{C}^{m\times n}$ such that $\|\bm{\Phi}\mathbf{x}\|_{1}>0$ and suppose that we observe $\breve{\mathbf{z}}\in\mathbb{C}^{m}$ that is a possibly noisy version of $\mathbf{z}:=\operatorname{sign}(\bm{\Phi}\mathbf{x})$.
+Suppose that we compute $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ with $\hat{\mathbf{x}}$ obtained by solving $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1},\varepsilon)$ in (1.8) with $\mathbf{A}_{\breve{\mathbf{z}}}$ defined in (1.7) and some $\varepsilon\geq 0$.
+If $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$ and
+$\varepsilon\geq\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\Big[\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\Big],$ (3.3)
+then
+$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{10\sigma_{\ell_{1}}(\mathbf{x};\Sigma^{n}_{s})}{\sqrt{s}}+\frac{14\|\bm{\Phi}\mathbf{x}\|_{1}\varepsilon}{\kappa m}.$
+Note that the condition (3.3) can be relaxed to $\varepsilon\geq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$.
+Proof.
+By (1.7), $\mathbf{x}^{\star}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\mathbf{x}$ and $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$, we have
+$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}\|_{2}=\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{x}\|_{2}$
+$\leq\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\Big[\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\Big]\leq\varepsilon.$
+(In view of the above display, the condition (3.3) can be relaxed to $\varepsilon\geq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$.) Combining with $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$, by Proposition 2.1 we obtain
+$\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 7\varepsilon+5\frac{\sigma_{\ell_{1}}(\mathbf{x}^{\star},\Sigma^{n}_{s})}{\sqrt{s}}.$
+By $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$, $\mathbf{x}=\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}$, (2.12) and $\|\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}$,
+$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ $=\Big\|\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}-\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}\Big\|_{2}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\max\{\|\hat{\mathbf{x}}\|_{2},\|\mathbf{x}^{\star}\|_{2}\}}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\|\mathbf{x}^{\star}\|_{2}}$
+$\leq\frac{14\varepsilon}{\|\mathbf{x}^{\star}\|_{2}}+\frac{10\sigma_{\ell_{1}}(\mathbf{x}^{\star},\Sigma^{n}_{s})}{\sqrt{s}\|\mathbf{x}^{\star}\|_{2}}=\frac{14\|\bm{\Phi}\mathbf{x}\|_{1}\varepsilon}{\kappa m}+\frac{10\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}},$
+as desired. ∎
 3.1 Instance Optimality
-Our first result concerns the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ for arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$ over a general cone $\mathcal{U}$.
+For any $\mathbf{w}\in\mathbb{C}^{m}$, recall that the matrix $\mathbf{A}_{\mathbf{w}}$ is defined in (1.7).
+Our first result concerns the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\mathcal{K}\}$ for arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$ over a general cone $\mathcal{U}$.
 Theorem 3.1 (RIP of a set of $\mathbf{A}_{\mathbf{z}}$).
@@
 If
-$m\geq C_{1}\left(\frac{\omega^{2}(\mathcal{U}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right)$ (3.2)
+$m\geq C_{1}\left(\frac{\omega^{2}(\mathcal{U}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right)$ (3.4)
 for some large enough $C_{1}$, then for some $C_{2}$ the event
-$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U},C_{2}\eta\log^{1/2}(\eta^{-1})),\quad\forall\mathbf{x}\in\mathcal{K}$
-holds with probability at least $1-C_{3}\exp(-c_{4}\omega^{2}(\mathcal{U}^{\rm(N)})-c_{4}\mathscr{H}(\mathcal{K},r))-C_{5}\exp(-c_{6}\eta^{2}m)$.
+$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U},C_{2}\eta\log^{1/2}(\eta^{-1})),\quad\forall\mathbf{x}\in\mathcal{K},\quad\text{where }\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$
+holds with probability at least $1-C_{3}\exp(-c_{4}\eta^{2}m)$.
 Remark 3.1 (Recovering $\mathcal{K}=\Sigma^{n,*}_{s}$ in [10]).
@@
 […]
-Further, setting $\eta=\frac{c\delta}{\sqrt{\log(\delta^{-1})}}$ with small enough $c$, Theorem 3.1 yields the following: The matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy $\operatorname{RIP}(\Sigma^{n}_{2s},\delta)$ (w.h.p.) as long as $m=\Omega(\delta^{-2}\log^{2}(\delta^{-1})s\log(\frac{en}{s}))$.
+Further, setting $\eta=\frac{c\delta}{\sqrt{\log(\delta^{-1})}}$ with small enough $c$, Theorem 3.1 yields the following: The matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy $\operatorname{RIP}(\Sigma^{n}_{2s},\delta)$ (w.h.p.) as long as $m=\Omega(\delta^{-2}\log^{2}(\delta^{-1})s\log(\frac{en}{s}))$.
 This improves on [10, Thm.
@@
 More importantly, Theorem 3.1 applies to arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$ with a number of measurements proportional to $\omega^{2}(\mathcal{K})$.
-To see this, by Sudakov’s inequality (2.4) and $\omega^{2}(\mathcal{K}_{(t)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})=4\omega^{2}(\mathcal{K})$ [49, Sec.
+To see this, by Sudakov’s inequality (2.4) and $\omega^{2}(\mathcal{K}_{(t)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})=4\omega^{2}(\mathcal{K})$ [50, Sec.
 7.5.1], we find that (3.4) can be implied by the following based only on the Gaussian width,
-$m\geq C_{1}^{\prime}\left(\frac{\omega^{2}(\mathcal{U}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K})}{\eta^{8}}\right)\quad\text{with large enough }C_{1}^{\prime}.$ (3.3)
+$m\geq C_{1}^{\prime}\left(\frac{\omega^{2}(\mathcal{U}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K})}{\eta^{8}}\right)\quad\text{with large enough }C_{1}^{\prime}.$ (3.5)
 The first informal theorem in introduction thus follows.
@@
 If $m\geq C_{1}s\log(\frac{en}{s})$, then with probability at least $1-C_{2}\exp(-c_{3}s\log(\frac{en}{s}))$ over the complex Gaussian $\bm{\Phi}$, we have
-$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3),\qquad\forall\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}.$
-The distortion $1/3$ can be replaced by any given positive constant $\delta$, up to changes in the values of $C_{1},C_{2},c_{3}$.
-The proof of Theorem 3.1 is based on covering and analogous to [10].
+$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3),\quad\forall\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1},$
+where $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$. The distortion $1/3$ can be replaced by any given positive constant $\delta$, up to changes in the values of $C_{1},C_{2},c_{3}$.
+The proof of Theorem 3.1 is based on covering techniques and is analogous to [10].
 Nonetheless, [10] is restricted to $\mathcal{K}=\Sigma^{n,*}_{s}$ or at most other $\mathcal{K}$ with metric entropy logarithmically depending on the covering radius, and the techniques therein do not suffice for proving Corollary 1.
@@
 We assume the event in Corollary 1 holds and consider any $\mathbf{x}\in\mathbb{S}^{n-1}$.
-If $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})$, then the event in Corollary 1 gives $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3)$, and further Proposition 2.1 implies
-$\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 5\frac{\sigma_{\ell_{1}}(\mathbf{x}^{\star},\Sigma^{n}_{s})}{\sqrt{s}}=5\|\mathbf{x}^{\star}\|_{2}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}.$
-Finally, we use (2.12) to obtain
-$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}=\left\|\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}-\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}\right\|_{2}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\|\mathbf{x}^{\star}\|_{2}}\leq 10\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}},$
-as claimed.
+If $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})$, then the event in Corollary 1 gives $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3)$, and further Lemma 3.1 implies
+$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 10\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}.$
 If $\mathbf{x}\notin\mathbb{B}_{1}^{n}(\sqrt{2s})$, meaning that $\|\mathbf{x}\|_{1}>\sqrt{2s}$, then we let $\mathbf{x}_{[s]}=\mathrm{arg}\min_{\mathbf{u}\in\Sigma^{n}_{s}}\|\mathbf{u}-\mathbf{x}\|_{2}$ and notice that $\|\mathbf{x}_{[s]}\|_{1}\leq\sqrt{s}$, and we have $\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})=\|\mathbf{x}-\mathbf{x}_{[s]}\|_{1}\geq\|\mathbf{x}\|_{1}-\|\mathbf{x}_{[s]}\|_{1}\geq(\sqrt{2}-1)\sqrt{s}$.
@@
 While we focus on sparse recovery via basis pursuit, the generality of Theorem 3.1 in terms of $(\mathcal{K},\mathcal{U})$ allows for a straightforward generalization to other signal structures, such as PO-CS of low-rank matrices.
-Our subsequent technical results (Theorem 3.7, Lemmas 4.1, 4.2) are also presented in a similar manner.
-Also note that Theorem 3.2 directly generalizes to other instance optimal algorithms such as iterative hard thresholding (see [20, Sec.
-6]).
+Our subsequent technical results (Theorem 3.7, Lemmas 4.2, 4.1) are also presented in a similar manner.
 3.2 Bounded Dense Noise
-Next, we proceed to explore the robustness to noise and corruption.
-We consider the nonuniform robustness concerning the reconstruction of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$, which restricts our attention to robustness without being distructed by the concerns of uniformity or instance optimality.
-We will discuss in Section 4 the work needed to improve these forthcoming nonuniform robustness results to uniform ones with instance optimality.
-We first consider pre-sign small dense noise $\bm{\tau}\in\mathbb{C}^{m}$ obeying $\|\bm{\tau}\|_{\infty}\leq\tau_{0}$, that is, observations given by $\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}$ [^5] We emphasize that the only constraint on $\bm{\tau}$ is a small enough max norm; under this constraint, it can be generated by an adversary having full knowledge of $(\bm{\Phi},\mathbf{x})$.
+Next, we study robustness to noise and corruption in a nonuniform setting.
+Specifically, we consider the reconstruction of a fixed sparse signal $\mathbf{x}\in\Sigma^{n,*}_{s},$ which allows us to decouple the desiderata of uniformity, instance optimality and robustness, and simply focus on the robustness for now.
+We discuss in Section 4 how these nonuniform results can be extended to uniform, instance-optimal guarantees.
+For convenience, we specialize Lemma 3.1 to our setting.
+Lemma 3.2 (Specialization of Lemma 3.1).
+Under complex Gaussian $\bm{\Phi}$ and a fixed $\mathbf{x}\in\Sigma^{n,*}_{s},$ let $\breve{\mathbf{z}}$ be a noisy version of $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$, and suppose that we compute $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ with $\hat{\mathbf{x}}$ obtained by solving $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1},\varepsilon)$ in (1.8).
+If (3.3) and
+$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}<\frac{1}{9}.$ (3.7)
+hold, then with probability at least $1-C\exp(-cs\log(\frac{en}{s}))$, we have
+$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{14\|\bm{\Phi}\mathbf{x}\|_{1}\varepsilon}{\kappa m}.$ (3.8)
+Proof.
+The statement follows from Lemma 3.1 after the validation of $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
+By $\mathbf{A}_{\breve{\mathbf{z}}}=\mathbf{A}_{\mathbf{z}}+\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}$, we have
+$\|\mathbf{A}_{\mathbf{z}}\mathbf{u}\|_{2}-\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{u}\|_{2}\leq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}\|_{2}\leq\|\mathbf{A}_{\mathbf{z}}\mathbf{u}\|_{2}+\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{u}\|_{2}\,,\,\forall\mathbf{u}\in\Sigma^{n,*}_{2s}$.
+By the RIP of $\mathbf{A}_{\mathbf{z}}$ in Corollary 1,
+$1-c^{\prime}\leq\|\mathbf{A}_{\mathbf{z}}\mathbf{u}\|_{2}\leq 1+c^{\prime}\,,\,\forall\mathbf{u}\in\Sigma^{n,*}_{2s}\,,\quad\text{where~}c^{\prime}:=\min\Big\{\sqrt{\frac{4}{3}}-\frac{10}{9},\frac{8}{9}-\sqrt{\frac{2}{3}}\Big\}$ (3.9)
+holds with the promised probability. By (1.7), triangle inequality and (3.7), we obtain
+$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{u}\|_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}<\frac{1}{9}.$
+Combining the preceding three displays yields $\sqrt{\frac{2}{3}}\leq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}\|_{2}\leq\sqrt{\frac{4}{3}}\,,\,\forall\mathbf{u}\in\Sigma^{n,*}_{2s},$ which is exactly $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
+∎
+We first consider post-sign small dense noise $\bm{\tau}\in\mathbb{C}^{m}$ obeying $\|\bm{\tau}\|_{\infty}\leq\tau_{0}$, that is, observations given by $\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}$.
+[^6] We emphasize that the only constraint on $\bm{\tau}$ is a small enough max norm; under this constraint, it can be generated by an adversary having full knowledge of $(\bm{\Phi},\mathbf{x})$.
 This has been treated in [27, Sec.
@@
 3].
-We reproduce the result and proof here to demonstrate the two-step analysis for noisy PO-CS from [27]:
-  • (1)
-    Show the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$—based on the RIP of $\mathbf{A}_{\mathbf{z}}$ (Corollary 1) we only need to control the impact of noise;
-  • (2)
-    Estimate $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$, which is the noise level in the resulting linear compressed sensing problem, to indicate the choice of $\epsilon$ in (1.8).
+We reproduce the result here as an illustration of the applicability of Lemma 3.2.
 Theorem 3.3 (Post-sign noise).
@@
 Proof.
-By the linearity of $\mathbf{A}_{\mathbf{u}}$ on $\mathbf{u}$ and $\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}$, we have $\mathbf{A}_{\breve{\mathbf{z}}}=\mathbf{A}_{\mathbf{z}}+\mathbf{A}_{\bm{\tau}}$.
-We proceed in several steps.
-Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: We first show that $\mathbf{A}_{\bm{\tau}}$ cannot have great affect on the RIP of $\mathbf{A}_{\mathbf{z}}$ under small enough $\tau_{0}$.
-To that end, we need to bound $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}}\mathbf{u}\|_{2}$.
-By (1.7) and triangle inequality, we have
-$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}}\mathbf{u}\|_{2}$ $\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}$ (3.5)
-$\leq\tau_{0}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}+\tau_{0}\left(\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\right).$
+By Lemma 3.2, we only need to establish (3.7) and (3.3) with $\varepsilon=\frac{5}{2}\tau_{0}$ with the promised probability: once these two conditions hold true, then by Lemma 3.2 we have $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{\|\bm{\Phi}\mathbf{x}\|_{1}35\tau_{0}}{\kappa m}$, and further using $\mathbb{P}(\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}\leq\frac{36}{35})\geq 1-\exp(-cm)$ yields the claim (recall that $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=O(1/\sqrt{m})$ due to (2.1)).
+Establish (3.7): Note that $\breve{\mathbf{z}}-\mathbf{z}=\bm{\tau}$ satisfies $\|\bm{\tau}\|_{\infty}\leq\tau_{0}\leq\frac{1}{36}$, we use triangle inequality to obtain
+$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}$
+$\leq\tau_{0}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}+\tau_{0}\left(\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\right).$ (3.10)
 We use a concentration bound from prior works in the area: by [10, Lem. 6] (or [19, Thm. 6]), if $m\geq C_{1}s\log(\frac{en}{s})$ for large enough $C_{1}$, then we have
-$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}\leq 1+c_{1}$
+$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}\leq 1+c_{1}$ (3.11)
 for some $c_{1}$ that can be set sufficiently small with probability at least $1-2\exp(-c_{2}s\log(\frac{en}{s}))$.
@@
 $\mathcal{N}(0,1)$ entries, then (2.9) gives that if $m\geq C_{1}s\log(\frac{en}{s})$ for large enough $C_{1}$, we have
-$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\leq 2+c_{3}$
-for some $c_{3}$ that can be set sufficiently small, with probability at least $1-\exp(-s\log(\frac{en}{s}))$.
-Setting $c_{1}=c_{3}=\frac{1}{2}$, we obtain $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}}\mathbf{u}\|_{2}\leq 4\tau_{0}\leq\frac{1}{9}$ with the promised probability.
-We now invoke Corollary 1 to ensure small enough $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}|\|\mathbf{A}_{\mathbf{z}}\mathbf{u}\|_{2}-1|$ when $m\geq C_{1}s\log(\frac{en}{s})$ for large enough $C_{1}$, with the desired probability.
-Taken collectively, we find that $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
-Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: Next, we bound
-$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\|\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}+\mathbf{A}_{\bm{\tau}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\|\mathbf{A}_{\bm{\tau}}\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\bm{\tau}}\mathbf{x}\|_{2}$ (3.6)
-for fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ and the corresponding $\mathbf{x}^{\star}=\frac{\kappa m\cdot\mathbf{x}}{\|\bm{\Phi}\mathbf{x}\|_{1}}$. Again using (1.7) we have
-$\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\tau}\mathbf{x}\|_{2}$ $\leq\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\left[\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right]$
-$\leq\tau_{0}+\tau_{0}\cdot\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\cdot\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\leq\frac{5\tau_{0}}{2},$ (3.7)
-where in the last inequality we use sub-Gaussian tail bounds to ensure sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ (recall that $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=O(1/\sqrt{m})$ due to (2.1)), and Bernstein’s inequality [49, Thm.
-2.8.1] to ensure sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{2}^{2}}{m}-2|$, with the promised probability.
-Now by Proposition 2.1 we have $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 7\cdot\frac{5\tau_{0}}{2}=\frac{35}{2}\tau_{0}$.
-Then by (2.12) and the condition of $\|\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}$ being sufficiently close to $1$ we have
-$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}=\left\|\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}-\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}\right\|_{2}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\|\mathbf{x}^{\star}\|_{2}}\leq 36\tau_{0},$ (3.8)
-as claimed. ∎
+$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\leq 2+c_{3}$ (3.12)
+for some $c_{3}$ that can be set sufficiently small, with probability at least $1-\exp(-s\log(\frac{en}{s}))$. Setting $c_{1}=c_{3}=\frac{1}{2}$, we obtain (3.7) with the promised probability.
+Establish (3.3) with $\varepsilon=\frac{5}{2}\tau_{0}$: This can be seen by
+$\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\left[\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right]\leq\tau_{0}+\tau_{0}\cdot\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\cdot\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\leq\frac{5\tau_{0}}{2},$ (3.13)
+where in the first inequality we use $|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{x})|\leq\tau_{0}\|\bm{\Phi}\mathbf{x}\|_{1}$ and $\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{x})\|_{2}\leq\tau_{0}\|\bm{\Phi}\mathbf{x}\|_{2}$, in the second inequality we use $\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\cdot\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}<\frac{3}{2}$ which holds because sub-Gaussian tail bound in (3.2) guarantees sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$, and Bernstein’s inequality [50, Thm.
+2.8.1] yields $\mathbb{P}(|\frac{\|\bm{\Phi}\mathbf{x}\|_{2}^{2}}{m}-2|\geq t)\leq 2\exp(-c_{4}mt^{2})$ for any $t\in(0,1)$ and thus guarantees sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{2}^{2}}{m}-2|$ with the promised probability.
+∎
 While explicit constants are provided in some of our results, no attempts have been made to optimize them.
-We move to post-sign small dense noise.
+We move to pre-sign small dense noise.
 We again denote the dense noise by $\bm{\tau}$, but now the noisy observations are $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})$.
@@
 The reason is that a small enough pre-sign perturbation $\tau$ can still greatly affect $\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau)$ if $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ is small.
-This makes the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$ less evident, for which we have to separately treat a small fraction of measurements with small $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ and the majority with $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ bounded away from 00.
+This makes the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$ less evident, for which we have to separately treat a small fraction of measurements with small $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ and the majority with $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ bounded away from 0.
 On the other hand, it comes a bit surprising that the error horizon remains at $O(\tau_{0})$, since some algebra with (2.12) finds $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=O(\tau_{0})$.
@@
 Proof.
-We first transfer $\bm{\tau}$ to post-sign noise by writing $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\tilde{\bm{\tau}}$, where
+In light of Lemma 3.2, we only need to establish (3.7) and (3.3) with $\varepsilon=4\tau_{0}$ with the promised probability, then by Lemma 3.2 we have $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{\|\bm{\Phi}\mathbf{x}\|_{1}\cdot 56\tau_{0}}{\kappa m}$, which leads to the claim on the event $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}\leq\frac{57}{56}$ that holds with probability at least $1-2\exp(-cm)$.
+To get started, we first transfer $\bm{\tau}$ to post-sign noise by writing $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\tilde{\bm{\tau}}$, where
 $\tilde{\bm{\tau}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})-\operatorname{sign}(\bm{\Phi}\mathbf{x}).$
 The entries of $\tilde{\bm{\tau}}$ may not be uniformly small, but we can establish a decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ where $\|\tilde{\bm{\tau}}_{1}\|_{\infty}$ is small and $\tilde{\bm{\tau}}_{2}$ is sparse.
-For the fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ and for some $\eta_{0}>0$ to be chosen, we have
+We let $\eta_{0}:=144c_{0}$.
+For the fixed $\mathbf{x}\in\Sigma^{n,*}_{s},$ we have
 $\mathbbm{P}(|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta_{0})\leq\mathbbm{P}\big(|\Re(\bm{\Phi}_{i}^{*}\mathbf{x})|\leq\eta_{0}\big)\leq\sqrt{\frac{2}{\pi}}\eta_{0}.$
 Letting $\mathcal{J}_{\mathbf{x},\eta_{0}}=\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta_{0}\}$, the Chernoff bound gives
-$\mathbbm{P}\big{(}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\big{)}\geq\mathbbm{P}\Big{(}{\rm Bin}(m,\sqrt{2/\pi}\eta_{0})\leq\eta_{0}m\Big{)}\geq 1-\exp(-c_{1}\eta_{0}m)$ (3.9)
+$\mathbbm{P}\big(|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\big)\geq\mathbbm{P}\Big({\rm Bin}(m,\sqrt{2/\pi}\cdot\eta_{0})\leq\eta_{0}m\Big)\geq 1-\exp(-c_{1}\eta_{0}m)$ (3.14)
 for some $c_{1}>0$.
@@
 Moreover, the entries of $\tilde{\bm{\tau}}_{1}$ take the form
-$[\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})]\mathbbm{1}(|\bm{\Phi}_{i}^{*}\mathbf{x}|>\eta_{0}),$
-and hence by (2.11) we obtain $\|\tilde{\bm{\tau}}_{1}\|_{\infty}\leq\frac{2\tau_{0}}{\eta_{0}}$.
-The observations can now be expressed as $\breve{\mathbf{z}}=\mathbf{z}+\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}.$
-Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: We now want to show $\mathbf{A}_{\breve{\mathbf{z}}}=\mathbf{A}_{\mathbf{z}}+\mathbf{A}_{\tilde{\bm{\tau}}_{1}}+\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
-Similarly to the proof of Theorem 3.3, we can use Corollary 1 to obtain $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\bm{\Sigma}^{n}_{2s},c_{2})$ with sufficiently small $c_{2}$, and all that remains is to show $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{1}}\mathbf{u}\|_{2}$ and $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}$ are both small enough.
-Now let us fix $\eta_{0}$ to be some small enough positive constant.
-With the promised probability, the arguments from the proof of Theorem 3.3 imply
-$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{1}}\mathbf{u}\|_{2}\leq 4\|\tilde{\bm{\tau}}_{1}\|_{\infty}\leq\frac{8\tau_{0}}{\eta_{0}},$
-which is small enough because of sufficiently small $\tau_{0}$ (and a specified $\eta_{0}>0$).
-Next, we need to make sure $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}$ is sufficiently small.
-We let $\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\in\{0,1\}^{m}$ be the vector whose $1$’s indicate the support of $\tilde{\bm{\tau}}_{2}$.
-Starting as in (3.5) while proceeding with $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$, $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$, and Cauchy-Schwarz inequality, we obtain
-$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\tilde{\bm{\tau}}_{2}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}_{2}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}$
-$\leq\|\tilde{\bm{\tau}}_{2}\|_{2}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\kappa m}+\|\tilde{\bm{\tau}}_{2}\|_{\infty}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\sqrt{m}}$
+$\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})]\mathbbm{1}(|\bm{\Phi}_{i}^{*}\mathbf{x}|>\eta_{0}),$
+and hence by (2.11) we obtain $\|\tilde{\bm{\tau}}_{1}\|_{\infty}\leq\frac{2\tau_{0}}{\eta_{0}}$. The observations can now be expressed as
+$\breve{\mathbf{z}}=\mathbf{z}+\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}.$ (3.15)
+Establish (3.7): By (3.15) and triangle inequality, to guarantee (3.7), it is sufficient to establish
+$T_{1}:=\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\tilde{\bm{\tau}}_{1}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}_{1}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}\leq\frac{1}{18},$ (3.16)
+$T_{2}:=\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\tilde{\bm{\tau}}_{2}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}_{2}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}\leq\frac{1}{18}.$ (3.17)
+[…]
+$T_{2}$ $\leq\|\tilde{\bm{\tau}}_{2}\|_{2}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\kappa m}+\|\tilde{\bm{\tau}}_{2}\|_{\infty}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\sqrt{m}}$
 $\leq\Big(2+\frac{\sqrt{\eta_{0}}}{\kappa}\Big)\sqrt{\eta_{0}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\sqrt{\eta_{0}m}}.$ (3.18)
@@
 for some absolute constant $C_{3}$ with the probability $2\exp(-c_{3}\eta_{0}\log(\frac{e}{\eta_{0}})m)$. Under $m=\Omega(s\log(\frac{en}{s}))$, substituting (3.19) into (3.18) yields
-$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\eta_{0}\log(\frac{e}{\eta_{0}})}\right),$ (3.12)
-which is small enough due to the scaling of $m$ and sufficiently small $\eta_{0}$.
-Overall, we have arrived at the desired $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
-Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: Next, as with (3.6), we bound the $\ell_{2}$ measurement error at $\mathbf{x}^{\star}$,
-$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}\|_{2}\leq(1+c_{4})\left[\frac{|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right],$ (3.13)
-where $c_{4}>0$ can be small enough due to the sub-Gaussian concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$.
+$T_{2}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\eta_{0}\log(\frac{e}{\eta_{0}})}\right),$ (3.20)
+which is small enough due to the scaling of $m$, $\eta_{0}=144c_{0}$ and the small enough $c_{0}$.
+Establish (3.3) with $\varepsilon=4\tau_{0}$: Next, we seek to show
+$\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\Big[\frac{|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\breve{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\Big]\leq 4\tau_{0}.$ (3.21)
 Letting $\tilde{\tau}_{i}=\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})$ be the $i$-th entry of $\tilde{\bm{\tau}}$, we have
@@
 $|\tilde{\tau}_{i}^{*}\bm{\Phi}_{i}^{*}\mathbf{x}|=|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})||\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\frac{2|\tau_{i}|}{|\bm{\Phi}_{i}^{*}\mathbf{x}|}|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq 2|\tau_{i}|\leq 2\tau_{0}.$ (3.24)
-Thus, we obtain $|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|\leq 2m\tau_{0}$ and $\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}\leq 2\sqrt{m}\tau_{0}$, and hence by (3.13) with small enough $c_{4}$ we obtain $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}\leq 4\tau_{0}$.
-Therefore, Proposition 2.1 implies $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 28\tau_{0}$.
-By using an argument analogous to (3.8), we arrive at $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 57\tau_{0}$.
+Thus, we obtain $|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|\leq 2m\tau_{0}$ and $\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}\leq 2\sqrt{m}\tau_{0}$, which, together with the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$ in Equation (3.2), yield (3.21).
+The proof is complete.
 ∎
@@
 The idea is to identify another signal $\mathbf{x}^{\prime}\in\Sigma^{n,*}_{s}$ that is indistinguishable from $\mathbf{x}$ and satisfies $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}=\tilde{\Omega}(\tau_{0})$.
-We leave the removal of log factors to future work.
 Theorem 3.5 ($O(\tau_{0})$ is nearly sharp).
@@
 $\bm{\Phi}_{i}^{*}\tilde{\bm{\delta}}=0,\qquad\forall i\in\mathcal{J}_{\mathbf{x},s/4m}.$ (3.27)
-On the event $\{|\mathcal{J}_{\mathbf{x},s/4m}|\leq\frac{s}{4}\}$, (3.19) translates into $\frac{s}{2}$ real linear equations, so such $\tilde{\bm{\delta}}$ exists when $s\geq 4$.
+On the event $\{|\mathcal{J}_{\mathbf{x},s/4m}|\leq\frac{s}{4}\}$, (3.27) translates into (no more than) $\frac{s}{2}$ real linear equations, so such $\tilde{\bm{\delta}}$ exists when $s\geq 4$.
 Now we consider $\mathbf{x}^{\prime}=\frac{\mathbf{x}+\tilde{\bm{\delta}}}{\|\mathbf{x}+\tilde{\bm{\delta}}\|_{2}}\in\Sigma^{n,*}_{s},$ and similarly to (3.26) we have $\frac{\tau_{*}}{2}\leq\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}\leq\frac{3\tau_{*}}{2}$, and $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ holds with probability at least $1-\frac{4}{m}$.
-We further bound $\|\operatorname{sign}(\bm{\Phi}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}\mathbf{x})\|_{\infty}$.
+We further bound $\|\operatorname{sign}(\bm{\Phi}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}\mathbf{x})\|_{\infty}$ by discussing two cases:
   • For $i\in\mathcal{J}_{\mathbf{x},s/4m}$, we have $\bm{\Phi}_{i}^{*}\tilde{\bm{\delta}}=0$ and hence $\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}^{\prime})=\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})$.
@@
     $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})|\leq\frac{2|\bm{\Phi}_{i}^{*}(\mathbf{x}^{\prime}-\mathbf{x})|}{s/(4m)}\leq 48C_{0}\log\big(\frac{en}{s}\big)\sqrt{\log m}\cdot\tau_{*},$
-where the last inequality we use $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ and substitute $m=C_{0}s\log(\frac{en}{s})$. It follows that
+    where the last inequality we use $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ and substitute $m=C_{0}s\log(\frac{en}{s})$.
+Taken collectively, it follows that
 $\|\operatorname{sign}(\bm{\Phi}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}\mathbf{x})\|_{\infty}\leq 48C_{0}\log\big(\frac{en}{s}\big)\sqrt{\log m}\cdot\tau_{*}.$
@@
 Therefore, no algorithm can distinguish $\mathbf{x}$ and $\mathbf{x}^{\prime}$ in the regime of Theorem 3.3 where an adversary can add post-sign noise bounded by $\tau_{0}$. ∎
-Simulation: [^6] The MATLAB codes for generating the figures in this paper are available in https://junrenchen58.github.io/.
+Simulation: [^7] The MATLAB code used to generate the figures in this paper is available at https://github.com/junrenchen58/instance-optimal-pocs.
 We pause to use experimental results to provide evidence of the achievability and tightness of $O(\tau_{0})$.
-In all of our experiments, the data points are averaged over $50$ independent trials, each of which concerns the recovery of $\mathbf{x}$ uniformly drawn from $\Sigma^{500,*}_{5}$ from $300$ phase-only measurements.
+In all of our experiments, the data points are averaged over $50$ independent trials, each of which concerns the recovery of $\mathbf{x}$ uniformly drawn from $\Sigma^{500,*}_{5}$ using $300$ phase-only measurements.
 We provide the optimally tuned $\varepsilon$ to basis pursuit (1.8), namely $\varepsilon=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$.
@@
 Figure 2: Reconstruction errors under pre-sign bounded by $\tau_{0}$
+Remark 3.4 (The choice of noises in simulations).
+Note that $O(\tau_{0})$ in Theorems 3.3, 3.4 is a worst-case upper bound that applies to arbitrary $\bm{\tau}$ obeying $\|\bm{\tau}\|_{\infty}\leq\tau_{0}$, but it may not be tight for some $\bm{\tau}$: for instance, under the pre-sign noise $\bm{\tau}=\tau_{0}\operatorname{sign}(\bm{\Phi}\mathbf{x})$, the observations $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ are identical to the noiseless ones $\mathbf{z}$, and therefore perfect recovery is achieved and $O(\tau_{0})$ is not tight.
+In view of Lemma 3.1, the final error is at the same order of $\varepsilon$ obeying $\varepsilon\geq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$.
+To numerically observe the tightness of $O(\tau_{0})$, the noises in Figures 12 are chosen such that $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\Omega(\tau_{0})$ with high probability.
 3.3 Sparse Phase Corruption
@@
 Theorem 3.6 (Sparse corruption).
-Consider PO-CS of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\zeta}$ with $\bm{\zeta}$ obeying $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ for some small enough $\zeta_{0}$ and $\|\bm{\zeta}\|_{\infty}\leq 2$.
-If $m\geq C_{1}s\log(\frac{en}{s})$ with sufficiently large $C_{1}$, then the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$, with $\hat{\mathbf{x}}$ being solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};11\zeta_{0}\log(\frac{e}{\zeta_{0}}))$ in (1.8), satisfies
+Consider PO-CS of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\zeta}$ with $\bm{\zeta}$ obeying $\|\bm{\zeta}\|_{\infty}\leq 2$ and $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ for some $\zeta_{0}\leq c_{0}$ where $c_{0}$ is a small enough absolute constant.
+If $m\geq C_{1}s\log(\frac{en}{s})$ with sufficiently large $C_{1}$, then the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$, with $\hat{\mathbf{x}}$ being solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};5\sqrt{\zeta_{0}\log(\frac{e}{\zeta_{0}})})$ in (1.8), satisfies
 $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 71\sqrt{\zeta_{0}\log(e/{\zeta_{0}})}$
@@
 Proof.
-Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: The first step is to show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
-As we can use Corollary 1 to show $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},c_{1})$ for sufficiently small $c_{1}$ with the promised probability, we only need to ensure small enough $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$.
-In the proof of Theorem 3.4 we show (3.12) for $\tilde{\bm{\tau}}_{2}$ satisfying $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$ and $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$.
-An identical argument yields
-$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\zeta_{0}\log(e/\zeta_{0})}\right)$
-with the probability we promise. Thus, $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ is small enough by the scaling of $m$ and $\zeta_{0}$ being small enough.
-Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: The second step is to bound the $\ell_{2}$ measurement error at $\mathbf{x}^{\star}=\frac{\kappa m\cdot\mathbf{x}}{\|\bm{\Phi}\mathbf{x}\|_{1}}$,
-$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\leq(1+c_{2})\left[\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right]$ (3.21)
-where $c_{2}>0$ is small enough due to the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$.
-We let $\mathbf{1}_{\operatorname{supp}(\bm{\zeta})}\in\{-1,1\}^{m}$ whose $1$’s indicate the support set of $\bm{\zeta}$.
+By Lemma 3.2, it suffices to establish (3.7) and (3.3) with $\varepsilon=5\sqrt{\zeta_{0}\log(\frac{e}{\zeta_{0}})}$, then we have (3.8) that reads $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{\|\bm{\Phi}\mathbf{x}\|_{1}\cdot 70\sqrt{\zeta_{0}\log(e/\zeta_{0})}}{\kappa m}$.
+This yields the claim due to $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}<\frac{71}{70}$ that holds with the promised probability (see Equation (3.2)).
+Establish (3.7): In the proof of Theorem 3.4 we show (3.20) for $\tilde{\bm{\tau}}_{2}$ satisfying $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$ and $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$.
+For $\bm{\zeta}=\breve{\mathbf{z}}-\mathbf{z}$ obeying $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ and $\|\bm{\zeta}\|_{\infty}\leq 2$ in the current setting, an identical argument yields
+$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\zeta_{0}\log(e/\zeta_{0})}\right)$
+with the probability we promise. This yields (3.7) by the scaling of $m$ and $\zeta_{0}\leq c_{0}$ for some small enough $c_{0}$.
+Establish (3.3) with $\varepsilon=5\sqrt{\zeta_{0}\log(\frac{e}{\zeta_{0}})}$: We let $\mathbf{1}_{\operatorname{supp}(\bm{\zeta})}\in\{-1,1\}^{m}$ whose $1$’s indicate the support set of $\bm{\zeta}$.
 Then we have
@@
 For fixed $I$ with cardinality $\zeta_{0}m$, $\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}$ follows Chi-squared distribution with $2\zeta_{0}m$ degrees of freedom.
-Then by a standard concentration bound [30, Lem.
+Then by a standard concentration bound [31, Lem.
 1], we obtain that for any $t\geq 0$ and any $I\subset[m]$ with $|I|=\zeta_{0}m$,
@@
 with probability at least $1-\exp(-\zeta_{0}m\log(\frac{e}{\zeta_{0}}))$.
-Combining (3.21), (3.22) and (3.23), using small enough $\zeta_{0},c_{2}$, we obtain $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}\leq 5\sqrt{\zeta_{0}\log(e/\zeta_{0})}$.
-Combining the two prior steps, we invoke Proposition 2.1 to obtain $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 37\sqrt{\zeta_{0}\log(e/{\zeta_{0}})}$.
-This leads to $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 71\sqrt{\zeta_{0}\log(e/\zeta_{0})}$ by re-iterating (3.8).
+Substituting (3.30) into (3.29), using small enough $\zeta_{0}$ and $\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\leq 1+c$ for small enough $c$ due to the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$ in Equation (3.2), we establish (3.3) with $\varepsilon=5\sqrt{\zeta_{0}\log(\frac{e}{\zeta_{0}})}$.
 ∎
@@
 Simulation: We pause to provide numerical evidence on the sharpness of $\tilde{O}(\sqrt{\zeta_{0}})$ for $\mathbf{x}^{\sharp}$, even under the optimally tuned noise level $\varepsilon=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$.
-We adopt the same settings as in earlier simulations but replace the dense noise $\bm{\tau}$ by $\zeta_{0}m$ adversarial phase corruptions.
-We test $\zeta_{0}m=\{1,2,3,5,7,9,11,13\}$ and corrupt the measurements through the mechanism described in Proposition 3.1.
+We adopt the same settings as in earlier simulations (i.e., $m=300$, $n=500$, $s=5$) but replace the dense noise $\bm{\tau}$ by $\zeta_{0}m$ adversarial phase corruptions.
+We test $\zeta_{0}m\in\{1,2,3,5,7,9,11,13\}$ and corrupt the measurements through the mechanism described in Proposition 3.1.
 The log-log curve in Figure 3(Left) roughly has a slope of $\frac{1}{2}$ over small $\zeta_{0}$.
@@
 Figure 3: The impact of sparse phase corruption on $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$
-As related context, in 1-bit compressed sensing, $\zeta_{0}m$ adversarial bit flips increment the $\ell_{2}$ error of the convex relaxation approach [40] by $\tilde{O}(\sqrt{\zeta_{0}})$, which was then improved to $\tilde{O}(\zeta_{0})$ using different algorithms [12, 11, 34, 1], and this (almost) linear increment is near-optimal under Gaussian designs (e.g., see [39, Thm.
+As related context, in 1-bit compressed sensing, $\zeta_{0}m$ adversarial bit flips increment the $\ell_{2}$ error of the convex relaxation approach [42] by $\tilde{O}(\sqrt{\zeta_{0}})$, which was then improved to $\tilde{O}(\zeta_{0})$ using different algorithms [12, 11, 35, 1], and this (almost) linear increment is near-optimal under Gaussian designs (e.g., see [40, Thm.
 2.4]).
@@
 Consider $\mathcal{U}_{c}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ for some cones $\mathcal{U}_{1}\subset\mathbb{R}^{n}$ and $\mathcal{U}_{2}\subset\mathbb{R}^{m}$, fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and given $\delta\in(0,1)$.
-For some absolute constants $C_{1}$ and $c_{2}$, if $m\geq C_{1}\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\rm(N)})$, then $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},\delta)$ with probability at least $1-\exp(-c_{2}\delta^{2}m)$.
+For some absolute constants $C_{1}$ and $c_{2}$, if $m\geq C_{1}\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})$, then $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},\delta)$ with probability at least $1-\exp(-c_{2}\delta^{2}m)$.
 The proof of Theorem 3.7 can be found in Appendix A.4. We can improve it to be a uniform statement over $\mathbf{x}\in\mathcal{K}$ for some $\mathcal{K}\subset\mathbb{S}^{n-1}$; see Lemma A.1.
@@
 We first show $\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\mathcal{U}_{c},\frac{1}{3})$ where $\mathcal{U}_{c}=\Sigma^{n}_{2s}\times\Sigma^{m}_{2\zeta_{0}m}$.
-By $\mathcal{U}_{c}^{\rm(N)}\subset(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})\times(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{n})$, we have $\omega(\mathcal{U}_{c}^{{\rm(N)}})\leq\omega(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+\omega(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})$ and hence
-$\omega^{2}(\mathcal{U}_{c}^{\rm(N)})\leq 2\omega^{2}(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+2\omega^{2}(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})\leq C_{0}\Big{(}s\log\big{(}\frac{en}{s}\big{)}+\zeta_{0}\log\big{(}\frac{e}{\zeta_{0}}\big{)}m\Big{)}$
+By $\mathcal{U}_{c}^{\mathbb{S}}\subset(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})\times(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{n})$, we have $\omega(\mathcal{U}_{c}^{\mathbb{S}})\leq\omega(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+\omega(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})$ and hence
+$\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})\leq 2\omega^{2}(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+2\omega^{2}(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})\leq C_{0}\Big(s\log\big(\frac{en}{s}\big)+\zeta_{0}\log\big(\frac{e}{\zeta_{0}}\big)m\Big)$
 for some absolute constant $C_{0}$.
-Therefore, when $\zeta_{0}$ is sufficiently small, $m\geq C_{1}s\log(\frac{en}{s})$ with large enough $C_{1}$ implies $m\geq C_{2}\omega^{2}(\mathcal{U}_{c}^{\rm(N)})$ with large enough $C_{2}$.
+Therefore, when $\zeta_{0}$ is sufficiently small, $m\geq C_{1}s\log(\frac{en}{s})$ with large enough $C_{1}$ implies $m\geq C_{2}\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})$ with large enough $C_{2}$.
 Then, by Theorem 3.7 we have $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},c_{3})$ with some small enough $c_{3}\leq\frac{1}{12}$ with probability at least $1-\exp(-c_{4}m)$.
-All that remains is to ensure $\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\rm(N)}}\|(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\|_{2}$ to be sufficiently small.
+All that remains is to ensure $\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\mathbb{S}}}\|(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\|_{2}$ to be sufficiently small.
 Comparing $\mathbf{A}_{\mathbf{w}}$ in (1.7) and $\widetilde{\mathbf{A}}_{\mathbf{w}}$ in (3.33), we have
-$\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\rm(N)}}\big{\|}(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\big{\|}_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\big{\|}(\mathbf{A}_{\breve{\mathbf{z}}}-\mathbf{A}_{\mathbf{z}})\mathbf{u}\big{\|}_{2}=\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}.$
+$\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\mathbb{S}}}\big\|(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\big\|_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\big\|(\mathbf{A}_{\breve{\mathbf{z}}}-\mathbf{A}_{\mathbf{z}})\mathbf{u}\big\|_{2}=\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}.$
 In the proofs of Theorems 3.4 and 3.6 we have shown the bound
 $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}=O\left(\sqrt{\frac{s\log(en/s)}{m}}+\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})}\right)$
-with probability at least $1-2\exp(-c_{5}\zeta_{0}\log(\frac{e}{\zeta_{0}})m)$; see (3.10)–(3.12).
+with probability at least $1-2\exp(-c_{5}\zeta_{0}\log(\frac{e}{\zeta_{0}})m)$; see Equations (3.18)–(3.20).
 Thus, $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ is small enough by the scaling of $m$ and small enough $\zeta_{0}.$
@@
 $\frac{\Re(\breve{\mathbf{z}}^{*}\bm{\Phi}\mathbf{x})}{\kappa m}=\frac{\Re(\mathbf{z}^{*}\bm{\Phi}\mathbf{x})+\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})}{\kappa m}=\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{1}{2}-\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}\geq\frac{1}{4},$
-where the last two inequalities hold with the promised probability due to the sub-Gaussian concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$ and the proven sufficiently small $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$.
+where the last two inequalities hold with the promised probability due to the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$ in Equation (3.2) and the sufficiently small $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$.
 ∎
@@
 We first announce our result, which is novel for nonlinear compressed sensing and closely resembles the standard guarantee in linear case (e.g., Proposition 2.1).
-It is also a formal version of the second informal statement provided in introduction.
+It is also the formal version of the second informal statement provided in introduction.
 Theorem 4.1 (Robust, instance optimal & uniform).
-Consider the above setting with sufficiently small $\tau_{0}$ and $\zeta_{0}$, and the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ where $\hat{\mathbf{x}}$ is solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\varepsilon)$ in (1.8) with
-$\varepsilon=C_{1}\tau_{0}+C_{2}\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})}+C_{3}\sqrt{\frac{s\log({en}/{s})}{m}}$ (4.1)
+Consider the above setting with $\max\{\tau_{0},\zeta_{0}\}\leq c_{0}$ for some sufficiently small $c_{0}$, and the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ where $\hat{\mathbf{x}}$ is solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\varepsilon)$ in (1.8) with
+$\varepsilon=C_{1}\tau_{0}+C_{2}\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})}+C_{3}\sqrt{\frac{s\log({en}/{s})}{m}}\mathbbm{1}(\zeta_{0}>0)$ (4.1)
 for some large enough absolute constants $C_{1},C_{2},C_{3}$.
@@
 4.1 Proof Strategy
-To prove Theorem 4.1, we essentially need to run the arguments in Theorems 3.2, 3.3, 3.4 and 3.6 again, along with some additional steps to make certain arguments uniform.
-To avoid repetition, we will only outline the proof strategies and emphasize the additional techniques.
-Without loss of generality, we can assume that the supports of $\bm{\zeta}_{1}$ and $\bm{\zeta}_{2}$ are disjoint.
-By Remark 3.4 we can consolidate $\bm{\zeta}_{1}$ and $\bm{\zeta}_{2}$ to rewrite $\breve{\mathbf{z}}$ as
+Without loss of generality, we can assume that the supports of $\bm{\zeta}_{(1)}$ and $\bm{\zeta}_{(2)}$ are disjoint.
+By Remark 3.5 we can consolidate $\bm{\zeta}_{(1)}$ and $\bm{\zeta}_{(2)}$ to rewrite $\breve{\mathbf{z}}$ as
 $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})+\bm{\tau}_{(2)}+\bm{\zeta},$
 where $\bm{\zeta}\in\mathbb{C}^{m}$ satisfies $\|\bm{\zeta}\|_{0}\leq 2\zeta_{0}m$ and $\|\bm{\zeta}\|_{\infty}\leq 2$.
-Recall also that we can focus on $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}:=\mathcal{X}$ (see the proof of Theorem 3.2).
 Further, we let $\tilde{\bm{\tau}}:=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$ and write
 $\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}_{(2)}+\tilde{\bm{\tau}}+\bm{\zeta}.$ (4.3)
-We then perform the two-step analysis.
-As we shall see, extra care is needed to ensure that every piece of the prior arguments is uniform for all $\mathbf{x}\in\mathcal{X}$, and the major additional work is on a uniform upper bound on $|\mathcal{J}_{\mathbf{x},\eta_{0}}|=|\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta_{0}\}|$ and a uniform concentration bound on $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$.
-We proceed in several steps.
-Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: Corollary 1 yields $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},c_{1})$ for some sufficiently small $c_{1}$, and one can check that the relevant arguments in the proofs of Theorems 3.3 and 3.6 can show $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{u}\|_{2}$ and $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ are sufficiently small.
-Therefore, it remains to show $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{u}\|_{2}$ to be small enough.
-In the proof of Theorem 3.4, the central idea is the decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ where $\tilde{\bm{\tau}}_{1}$ is near-dense and has entries bounded by $\frac{2\tau_{0}}{\eta_{0}}$, and $\tilde{\bm{\tau}}_{2}$ is $(\eta_{0}m)$-sparse with support set $\mathcal{J}_{\mathbf{x},\eta_{0}}$ and has entries bounded by $2$.
-We obtain the desired result by setting $\eta_{0}$ as a small absolute constant because we can then re-iterate the arguments for small dense noise to treat $\tilde{\bm{\tau}}_{1}$, and these for sparse phase corruption to treat $\tilde{\bm{\tau}}_{2}$.
-Such idea and most arguments remain valid, while the only notable issue is that the proof of Theorem 3.4 only shows $\|\tilde{\bm{\tau}}_{2}\|_{0}=|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m$ for a fixed $\mathbf{x}$ and hence only ensures the existence of the decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ for this $\mathbf{x}$.
-We need to strengthen this step and show that such a decomposition exists for all $\mathbf{x}\in\mathcal{X}$; to that end, we need to show $\sup_{\mathbf{x}\in\mathcal{X}}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m$.
-This is still true with high probability under $m=\Omega(s\log(\frac{en}{s}))$, as guaranteed by the following lemma.
-See a more refined statement and the proof in Appendix B.2.
-Lemma 4.1.
-Given some sufficiently small $\eta\in[\frac{C_{1}}{m},1)$ for some absolute constant $C_{1}$ and some $\mathcal{K}\subset\mathbb{S}^{n-1}$, if for sufficiently large $C_{2}$ we have $m\geq C_{2}\eta^{-3}\log(\eta^{-1})\omega^{2}(\mathcal{K})$, then we have
-$\mathbbm{P}\left(\sup_{\mathbf{x}\in\mathcal{K}}~{}|\mathcal{J}_{\mathbf{x},\eta}|\leq\eta m\right)\geq 1-3\exp(-c_{3}\eta m).$
-Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: By (4.3) and $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$, we seek to bound
-$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$ $\leq\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{x}^{\star}\|_{2}+\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}^{\star}\|_{2}+\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}$
-$=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\Big{(}\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{x}\|_{2}+\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}\|_{2}+\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\Big{)}.$
-Recall that we repeatedly use the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$; e.g., see (3.7), (3.13), (3.21)).
-While this follows from sub-Gaussian concentration for a fixed $\mathbf{x}$, we need to strengthen it to uniform concentration over all $\mathbf{x}\in\mathcal{X}$.
+Recall also that we can focus on $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}:=\mathcal{X}$ (see the proof of Theorem 3.2).
+We now make the following claim.
+Claim 1.
+To prove Theorem 4.1, all that remains is to establish
+[…]
+$\sup_{\mathbf{x}\in\mathcal{X}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}<\frac{1}{9},$ (4.5)
+$\frac{13}{14}\leq\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}\leq\frac{15}{14}\,,\quad\forall\mathbf{x}\in\mathcal{X}.$ (4.6)
+Proof.
+To see this claim, we note that (4.4)–(4.5) strengthen the conditions in Lemma 3.2 by making them uniform over $\mathbf{x}\in\mathcal{X}$.
+Since the RIP of $\mathbf{A}_{\mathbf{z}}$ in Corollary 1 used in Equation (3.9) is uniform over all $\mathbf{x}\in\mathcal{X}$ with the promised probability, the argument of Lemma 3.2 applies to all $\mathbf{x}\in\mathcal{X}$ and yields
+$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{10\sigma_{\ell_{1}}(\mathbf{x};\Sigma^{n}_{s})}{\sqrt{s}}+\frac{14\|\bm{\Phi}\mathbf{x}\|_{1}\varepsilon}{\kappa m}\,,\quad\forall\mathbf{x}\in\mathcal{X}.$
+(Note that the term $\frac{10\sigma_{\ell_{1}}(\mathbf{x};\Sigma^{n}_{s})}{\sqrt{s}}$ appears since $\mathbf{x}$ may not be in $\Sigma^{n,*}_{s}$.) Combining with (4.6) yields (4.2). ∎
+All that remains is to show (4.4)–(4.6) separately.
+To this end, the arguments in Theorems 3.2, 3.3, 3.4, and 3.6 must be revisited, together with additional techniques to ensure uniformity over all $\mathbf{x}\in\mathcal{X}$.
+To avoid repetition, we only outline the proof strategy and highlight the additional techniques.
+Show (4.6).
+Previously, we use (3.2) to achieve this for a fixed $\mathbf{x}$.
+Here, we need to strengthen it to uniform concentration over all $\mathbf{x}\in\mathcal{X}$.
 The following lemma shows that $\sup_{\mathbf{x}\in\mathcal{X}}|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ is small enough under $m=\Omega(s\log(\frac{en}{s}))$.
-Up to some simple modifications (from $\mathbb{R}$ to $\mathbb{C}$), the proof is identical to that of [42, Lem.
+Up to some simple modifications (from $\mathbb{R}$ to $\mathbb{C}$), the proof is identical to that of [44, Lem.
 2.1], and is hence omitted.
@@
 $\mathbbm{P}\left(\sup_{\mathbf{u}\in\mathcal{T}}\Big|\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}-1\Big|\leq\frac{C_{1}(\omega(\mathcal{T})+t)}{\sqrt{m}}\right)\geq 1-2\exp(-c_{2}t).$
-Now, for $\mathbf{w}=\bm{\tau}_{(2)},\tilde{\bm{\tau}},\bm{\zeta}$, it remains to bound
-$\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\mathbf{w}}\mathbf{x}\|_{2}\leq\sup_{\mathbf{x}\in\mathcal{X}}\frac{|\Re(\mathbf{w}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\Im(\operatorname{diag}(\mathbf{w}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}.$
-We have
-$\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{x}\|_{2}\leq\tau_{0}\left[\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\right]=O(\tau_{0})$
-due to (2.10).
-For $\tilde{\bm{\tau}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$, we also have $\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}\|_{2}=O(\tau_{0})$ as our original arguments in (3.14)–(3.16) are already uniform over $\mathbf{x}\in\mathcal{X}$.
-The bound on $\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}$ is a bit more tricky. From (3.22) we have
-$\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\leq C_{1}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=\zeta_{0}m\end{subarray}}\left(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\right)^{1/2},$
-and for a fixed $\mathbf{x}$ we achieve the bound $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}=O(\sqrt{\zeta_{0}\log(1/\zeta_{0})})$ through (3.23).
-In contrast, here we need a uniform bound over $\mathbf{x}\in\mathcal{X}$, and we use Lemma 2.1 to obtain
-$\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\leq C_{1}\sup_{\mathbf{x}\in\mathcal{X}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=\zeta_{0}m\end{subarray}}\left(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\right)^{1/2}=O\left(\sqrt{\zeta_{0}\log(e/\zeta_{0})}+\sqrt{\frac{s\log(en/s)}{m}}\right).$
-Therefore, a new term $O(\sqrt{\frac{s}{m}\log(\frac{en}{s})})$ arises in the error horizon, and consequently it also appears in our choice of the noise level (4.1).
+Show (4.5). By (4.3) and triangle inequality, it suffices to show, for $\mathbf{w}=\bm{\tau}_{(2)},\tilde{\bm{\tau}},\bm{\zeta}$, that
+$\sup_{\mathbf{x}\in\mathcal{X}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\mathbf{w}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\mathbf{w}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}<\frac{1}{27}.$ (4.7)
+For $\mathbf{w}=\bm{\tau}_{(2)}$ satisfying $\|\bm{\tau}_{(2)}\|_{\infty}\leq\tau_{0}$, re-iterating the arguments in Equations (3.10)–(3.12) shows that the left-hand side of (4.7) is bounded by $O(\tau_{0})$, and thus smaller than $\frac{1}{27}$ under small enough $\tau_{0}$.
+For $\mathbf{w}=\bm{\zeta}$ obeying $\|\bm{\zeta}\|_{0}\leq 2\zeta_{0}m$ and $\|\bm{\zeta}\|_{\infty}\leq 2$, the arguments in Equations (3.18)–(3.20) are sufficient to show that the left-hand side of (4.7) is bounded by $O(\sqrt{\frac{s}{m}\log(\frac{en}{s})})+\sqrt{\zeta_{0}\log(1/\zeta_{0})})$, and therefore is smaller than $\frac{1}{27}$ under small enough $\zeta_{0}$ and $m\gtrsim s\log(\frac{en}{s})$.
+It is more tricky to establish (4.7) for $\mathbf{w}=\tilde{\bm{\tau}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$ for $\bm{\tau}_{(1)}\leq\tau_{0}$.
+To this end, we first revisit the argument in Theorem 3.4: Define $\eta_{0}=C^{*}c_{0}$ for some suitably large absolute constant $C^{*}$ and the index set $\mathcal{J}_{\mathbf{x},\eta_{0}}:=\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta_{0}\}$, the key ingredient is a decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ where $\tilde{\bm{\tau}}_{1}$ and $\tilde{\bm{\tau}}_{2}$ accommodate the entries in $\mathcal{J}_{\mathbf{x},\eta_{0}}^{c}$ and $\mathcal{J}_{\mathbf{x},\eta_{0}}$, respectively; note that the entries of $\tilde{\bm{\tau}}_{1}$ is bounded by $\frac{2\tau_{0}}{\eta_{0}}\leq\frac{2c_{0}}{C^{*}c_{0}}=\frac{2}{C^{*}}$, which is suitably small under a large $C^{*}$, and thus $\tilde{\bm{\tau}}_{1}$ can be treated similarly to $\bm{\tau}_{(2)}$ by the arguments in Equations (3.10)–(3.12); under the high-probability event $\{|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\}$ (due to Chernoff bound), $\tilde{\bm{\tau}}_{2}$ satisfies $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$ and $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$; since $\eta_{0}=C^{*}c_{0}$ is small enough (under absolute constant $C^{*}$ and small enough $c_{0}$), $\tilde{\bm{\tau}}_{2}$ can be treated similarly to $\bm{\zeta}$ via the arguments in Equations (3.18)–(3.20).
+The proof strategy remains valid, with the only gap being that the event $\{|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\}$ for a fixed $\mathbf{x}$ needs to be strengthened to $\{\sup_{\mathbf{x}\in\mathcal{X}}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\}$, in order to guarantee that the decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ for some $(\eta_{0}m)$-sparse $\tilde{\bm{\tau}}_{2}$ exists for all $\mathbf{x}\in\mathcal{X}$.
+The following lemma shows that $\sup_{\mathbf{x}\in\mathcal{X}}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m$ holds with high probability under $m=\Omega(s\log(\frac{en}{s}))$.
+See a more refined statement and the proof in Appendix B.2.
+Lemma 4.2.
+Given some sufficiently small $\eta\in[\frac{C_{1}}{m},1)$ for some absolute constant $C_{1}$ and some $\mathcal{K}\subset\mathbb{S}^{n-1}$, if for sufficiently large $C_{2}$ we have $m\geq C_{2}\eta^{-3}\log(\eta^{-1})\omega^{2}(\mathcal{K})$, then we have
+$\mathbbm{P}\left(\sup_{\mathbf{x}\in\mathcal{K}}~|\mathcal{J}_{\mathbf{x},\eta}|\leq\eta m\right)\geq 1-3\exp(-c_{3}\eta m).$
+Show (4.4). By (4.3), triangle inequality, $\sup_{\mathbf{x}\in\mathcal{X}}\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\leq\frac{14}{13}$ from (4.6) and $\varepsilon$ in (4.1), it suffices to show
+$\sup_{\mathbf{x}\in\mathcal{X}}\frac{|\Re(\mathbf{w}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\Im(\operatorname{diag}(\mathbf{w}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}$
+$=O\left(\tau_{0}+\sqrt{\zeta_{0}\log(e/\zeta_{0})}+\sqrt{\frac{s\log(en/s)}{m}}\mathbbm{1}(\zeta_{0}>0)\right)$ (4.8)
+For $\mathbf{w}=\bm{\tau}_{(2)}$, $\tilde{\bm{\tau}}$ and $\bm{\zeta}$. For $\mathbf{w}=\bm{\tau}_{(2)}$ obeying $\|\bm{\tau}_{(2)}\|_{\infty}\leq\tau_{0}$, we use (2.10) and (4.6) to achieve
+$\sup_{\mathbf{x}\in\mathcal{X}}\frac{|\Re(\bm{\tau}_{(2)}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\Im(\operatorname{diag}(\bm{\tau}_{(2)}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\leq\tau_{0}\left(\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\right)=O(\tau_{0}).$
+For $\mathbf{w}=\tilde{\bm{\tau}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$, our original calculations in Equations (3.22)–(3.24) suffice to show that the left-hand side of (4.8) is bounded by $O(\tau_{0})$.
+When showing (4.8) with $\mathbf{w}=\bm{\zeta}$, however, an additional term $\sqrt{\frac{s}{m}\log(\frac{en}{s})}\mathbbm{1}(\zeta_{0}>0)$ arises compared to the non-uniform analysis in Theorem 3.6: For a fixed $\mathbf{x}$, Equations (3.29)–(3.30) establish
+$\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\leq C_{1}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=2\zeta_{0}m\end{subarray}}\Big(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\Big)^{1/2}\leq C_{2}\sqrt{\zeta_{0}\log(e/\zeta_{0})}\,;$
+in contrast, here we establish a uniform bound on $\mathbf{x}\in\mathcal{X}$ by using Lemma 2.1 to obtain
+$\sup_{\mathbf{x}\in\mathcal{X}}\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}$
+$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{X}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=2\zeta_{0}m\end{subarray}}\Big(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\Big)^{1/2}\leq C_{2}\sqrt{\zeta_{0}\log(e/\zeta_{0})}+C_{3}\sqrt{\frac{s\log(en/s)}{m}}\mathbbm{1}(\zeta_{0}>0),$
+yielding (4.8) with $\mathbf{w}=\bm{\zeta}$.
 5 Conclusion
@@
 We conjectured that this is tight for our specific estimator and provided some evidence.
-Yet we showed that it can be improved to 00 by proposing an extended linearization approach which perfectly recovers sparse signal under sparse corruption.
+Yet we showed that it can be improved to 0 by proposing an extended linearization approach which perfectly recovers sparse signal under sparse corruption.
 We believe the following questions are interesting for future study:
@@
 Acknowledgement
-The authors are thankful to Prof.
-Zhaoqiang Liu for his helps and discussions in the numerical simulations.
-Junren Chen is also thankful to Prof.
-Laurent Jacques for some stimulating discussions on instance optimality and PO-CS.
+The authors would like to thank Zhaoqiang Liu for his helps and discussions concerning the numerical simulations.
+J.C.
+also thanks Laurent Jacques for insightful discussions on instance optimality and phase-only compressive sensing.
 Appendix
@@
 Suppose $\mathcal{U}_{c}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ for some cones $\mathcal{U}_{1}\subset\mathbb{R}^{n}$ and $\mathcal{U}_{2}\subset\mathbb{R}^{m}$.
-There exist some absolute constants $c_{1},C_{1},C_{2},C_{3},c_{4},C_{5},c_{6}$ such that for any $\eta\in(0,c_{1})$, if we let $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$ and $\delta_{\eta}=C_{1}\eta(\log(\eta^{-1}))^{1/2}$, if
-$m\geq C_{2}\left[\frac{\omega^{2}(\mathcal{U}_{c}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right],$ (A.1)
+There exist some absolute constants $c_{1},C_{1},C_{2},C_{3},c_{4}$ such that for any $\eta\in(0,c_{1})$, if we let $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$ and $\delta_{\eta}=C_{1}\eta(\log(\eta^{-1}))^{1/2}$, if
+$m\geq C_{2}\left[\frac{\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right],$ (A.1)
 then the event
 $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},\delta_{\eta}),\quad\forall\mathbf{x}\in\mathcal{K}$ (A.2)
-holds with probability at least $1-C_{3}\exp(-c_{4}\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-c_{4}\mathscr{H}(\mathcal{K},r))-C_{5}\exp(-c_{6}\eta^{2}m)$.
+holds with probability at least $1-C_{3}\exp(-c_{4}\eta^{2}m)$.
 We first show that this statement immediately leads to Theorem 3.1.
@@
 By homogeneity and some algebra, we find that (A.2) is equivalent to
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\underbrace{\frac{[\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}]^{2}}{\kappa^{2}m^{2}}+\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}^{2}-1}_{:=f(\mathbf{x},\mathbf{u},\mathbf{w})}\Big{|}\leq\delta_{\eta}.$ (A.3)
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\underbrace{\frac{[\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}]^{2}}{\kappa^{2}m^{2}}+\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}^{2}-1}_{:=f(\mathbf{x},\mathbf{u},\mathbf{w})}\Big|\leq\delta_{\eta}.$ (A.3)
 Given the underlying signal $\mathbf{x}\in\mathbb{S}^{n-1}$ and some $\mathbf{u}\in\mathbb{R}^{n}$, we have the decomposition
 $\mathbf{u}=\underbrace{\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x}}_{:=\mathbf{u}_{\mathbf{x}}^{\|}}+\underbrace{(\mathbf{u}-\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x})}_{:=\mathbf{u}_{\mathbf{x}}^{\bot}}$ (A.4)
-where $\|\mathbf{u}_{\mathbf{x}}^{\|}\|_{2}^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}=\|\mathbf{u}\|_{2}^{2}.$ We recall that $\kappa:=\mathbbm{E}|\Phi_{i,j}|=\sqrt{\frac{\pi}{2}}$ and $\mathbf{u}_{\mathbf{x}}^{\bot}=\mathbf{u}-\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x}$, and observe that $1=\|\mathbf{u}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}=|\langle\mathbf{u},\mathbf{x}\rangle|^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}$ for $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$.
-Hence, we decompose $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|$ into
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.5)
+where $\|\mathbf{u}_{\mathbf{x}}^{\|}\|_{2}^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}=\|\mathbf{u}\|_{2}^{2}.$ We recall that $\kappa:=\mathbbm{E}|\Phi_{i,j}|=\sqrt{\frac{\pi}{2}}$ and $\mathbf{u}_{\mathbf{x}}^{\bot}=\mathbf{u}-\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x}$, and observe that $1=\|\mathbf{u}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}=|\langle\mathbf{u},\mathbf{x}\rangle|^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}$ for $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}$.
+Hence, we decompose $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|$ into
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.5)
 where
@@
 We refer to (A.6) as the parallel term and (A.7) as the orthogonal term.
-We first control $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}f^{\|}(\mathbf{x},\mathbf{u})$ in Lemma A.3 and $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}f^{\bot}(\mathbf{x},\mathbf{u})$ in Lemma A.5, and then strengthen them to uniform bounds over $\mathbf{x}\in\mathcal{K}$ by covering arguments in Lemma A.6 and Lemma A.7, respectively.
+We first control $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}f^{\|}(\mathbf{x},\mathbf{u})$ in Lemma A.3 and $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}f^{\bot}(\mathbf{x},\mathbf{u})$ in Lemma A.5, and then strengthen them to uniform bounds over $\mathbf{x}\in\mathcal{K}$ by covering arguments in Lemma A.6 and Lemma A.7, respectively.
 A.1 Technical Contributions
@@
 In the analysis, we need to bound
-$\frac{1}{\sqrt{m}}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{\|}\Im\big{[}\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big{]}\big{\|}_{2}$
-$=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i=1}^{m}\Big{[}\Im\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{]}^{2}\right)^{1/2}$
+$\frac{1}{\sqrt{m}}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big\|\Im\big[\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big]\big\|_{2}$
+$=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i=1}^{m}\Big[\Im\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big]^{2}\right)^{1/2}$
 where $\mathbf{x}_{r}:=\mathrm{arg}\min_{\mathbf{u}\in\mathcal{N}_{r}}\|\mathbf{u}-\mathbf{x}\|_{2}$ is the point in $\mathcal{N}_{r}$ (which is a minimal $r$-net of $\mathcal{K}$) closest to $\mathbf{x}$.
@@
 Thus, with respect to (A.9), they have to show the following is small enough:
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i\notin\mathcal{J}_{\mathbf{x},\eta}}\Big{[}\Im\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{]}^{2}\right)^{1/2}$
-$\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
-$\leq\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{v}|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}.$ (A.10)
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i\notin\mathcal{J}_{\mathbf{x},\eta}}\Big[\Im\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big]^{2}\right)^{1/2}$
+$\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
+$\leq\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{v}|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}.$ (A.10)
 This is a heavy-tailed random process that is in general hard to control.
 The argument in [10] is to use extremely small $r$ with $o(1)$ scaling to ensure this term to be small enough. Note that their equations (III.63)–(III.64) essentially bound (A.10) as
-$O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\|\bm{\Phi}\mathbf{v}\|_{\infty}\cdot\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\eta\sqrt{m}}\right)=O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\frac{\|\bm{\Phi}\mathbf{v}\|_{\infty}}{\eta}\right).$
+$O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\|\bm{\Phi}\mathbf{v}\|_{\infty}\cdot\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\eta\sqrt{m}}\right)=O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\frac{\|\bm{\Phi}\mathbf{v}\|_{\infty}}{\eta}\right).$
 By Gaussian concentration, it is easy to show that, with high probability,
@@
 Other Refinements: We also briefly note that we have refined or simplified some steps.
-As an example, in contrast to the covering approach taken in [10], we directly use known concentration bounds to establish the uniformity over $\mathcal{U}_{c}^{\rm(N)}$.
-Consequently, the sample complexity (A.1) is only based on the Gaussian width of $\mathcal{U}_{c}^{\rm(N)}$ and is free of its metric entropy.
+As an example, in contrast to the covering approach taken in [10], we directly use known concentration bounds to establish the uniformity over $\mathcal{U}_{c}^{\mathbb{S}}$.
+Consequently, the sample complexity (A.1) is only based on the Gaussian width of $\mathcal{U}_{c}^{\mathbb{S}}$ and is free of its metric entropy.
 As another example, while [10, Lem.
@@
 A.2 Fixed-$\mathbf{x}$ Bound on the Parallel Term
-Observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$ implies $\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}$, and that $f^{\|}(\mathbf{x},\mathbf{u})$ does not depend on $\mathbf{w}$.
+Observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}$ implies $\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}$, and that $f^{\|}(\mathbf{x},\mathbf{u})$ does not depend on $\mathbf{w}$.
 We start by utilizing
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}}|f^{\|}(\mathbf{x},\mathbf{u})|=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|,$ (A.12)
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}}|f^{\|}(\mathbf{x},\mathbf{u})|=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|,$ (A.12)
 where the last equality follows from $|f^{\|}(\mathbf{x},t\mathbf{u})|=t^{2}|f^{\|}(\mathbf{x},\mathbf{u})|$ for $t>0$. By $a^{2}-b^{2}=(a-b)(a+b)$, we have
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|$ $\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{\rm(N)}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big{|}\right)\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big{|}\right)$
-$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big{|},$ (A.13)
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|$ $\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big|\right)\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big|\right)$
+$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big|,$ (A.13)
 where (A.13) follows from
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big{|}\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\|\mathbf{z}\|_{2}\cdot\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa m}\Big{|}+1\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa\sqrt{m}}+1=O(1),$ (A.14)
-which holds with probability at least $1-4\exp(-cm)$ provided that $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see (2.10). Therefore, we only need to bound the term in (A.13), which reads as
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}-\langle\mathbf{x},\mathbf{u}\rangle\Big{|}:=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ (A.15)
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\|\mathbf{z}\|_{2}\cdot\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa m}\Big|+1\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa\sqrt{m}}+1=O(1),$ (A.14)
+which holds with probability at least $1-4\exp(-cm)$ provided that $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}}))$; see (2.10). Therefore, we only need to bound the term in (A.13), which reads as
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big(\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}\mathbf{u}\big)-\langle\mathbf{x},\mathbf{u}\rangle\Big|:=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ (A.15)
 We note that $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ is zero-mean.
@@
 $\mathbbm{E}\big[\kappa^{-1}\Re\big(\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}\mathbf{u}\big)\big]=\mathbbm{E}\big[\kappa^{-1}\Re\big(\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}\mathbf{u}_{\mathbf{x}}^{\|}\big)\big]=\langle\mathbf{x},\mathbf{u}\rangle\kappa^{-1}\mathbbm{E}|\bm{\Phi}_{i}^{*}\mathbf{x}|=\langle\mathbf{x},\mathbf{u}\rangle.$ (A.16)
-Because $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ is linear in $\mathbf{u}$, $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ can be treated as the supremum of a standard random process and hence be directly controlled by the following lemma.
+Because $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ is linear in $\mathbf{u}$, $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ can be treated as the supremum of a standard random process and hence be directly controlled by the following lemma.
 Lemma A.2 (See Sec. 8.6 in [50]).
@@
 Lemma A.3.
-Consider $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}f_{1}^{\|}(\mathbf{x},\mathbf{u})$ as in (A.15) with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and $\mathcal{U}_{1}^{\rm(N)}\subset\mathbb{S}^{n-1}$.
+Consider $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}f_{1}^{\|}(\mathbf{x},\mathbf{u})$ as in (A.15) with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and $\mathcal{U}_{1}^{\mathbb{S}}\subset\mathbb{S}^{n-1}$.
 Then for some absolute constant $C$ and any $t\geq 0$, we have
-$\mathbbm{P}\left(\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C[\omega(\mathcal{U}_{1}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2}).$ (A.17)
+$\mathbbm{P}\left(\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C[\omega(\mathcal{U}_{1}^{\mathbb{S}})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2}).$ (A.17)
 Proof.
-For any $\mathbf{u}_{1},\mathbf{u}_{2}\in\mathcal{U}_{1}^{\rm(N)}\cup\{0\}$, we have
+For any $\mathbf{u}_{1},\mathbf{u}_{2}\in\mathcal{U}_{1}^{\mathbb{S}}\cup\{0\}$, we have
 $\big\|f_{1}^{\|}(\mathbf{x},\mathbf{u}_{1})-f_{1}^{\|}(\mathbf{x},\mathbf{u}_{2})\big\|_{\psi_{2}}$
@@
 To deal with the orthogonal term $f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ defined in (A.7), we begin with
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ $\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}\right)$
-$\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}+\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}\right)$
-$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\underbrace{\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}}_{:=f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})}\Big{|},$ (A.19)
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$
+$\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big|\right)$
+$\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}+\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big|\right)$
+$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\underbrace{\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}}_{:=f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})}\Big|,$ (A.19)
 where the last inequality follows from $\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}\leq 1$ and
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}$ $\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{*}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}+1=O(1),$
-which holds with probability at least $1-4\exp(-cm)$ under $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see (2.10).
-It remains to bound $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$, where $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ is defined in (A.19).
-We first bound $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ for a fixed $\mathbf{x}$ by using the following extended matrix deviation inequality, along with the rotational invariance of $\bm{\Phi}$.
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}$ $\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{*}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}+1=O(1),$
+which holds with probability at least $1-4\exp(-cm)$ under $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}}))$; see (2.10).
+It remains to bound $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$, where $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ is defined in (A.19).
+We first bound $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ for a fixed $\mathbf{x}$ by using the following extended matrix deviation inequality, along with the rotational invariance of $\bm{\Phi}$.
 Lemma A.4 (see [7]).
@@
 Let $\mathbf{x}\in\mathbb{S}^{n-1}$ be fixed and $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ be given in (A.19). Then, for any $t\geq 0$, we have
-$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C[\omega(\mathcal{U}_{c}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2}).$
+$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C[\omega(\mathcal{U}_{c}^{\mathbb{S}})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2}).$
 Proof.
@@
 $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})=\left\|\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}\mathbf{e}_{1}))^{*})\tilde{\bm{\Phi}})\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}}{\sqrt{m}}+\mathbf{w}\right\|_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}~.$
-By $\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}\in\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathcal{U}^{\rm(N)}=\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathbf{P}^{\top}\mathbf{P}\mathcal{U}^{\rm(N)}=(\mathbf{I}_{n}-\mathbf{e}_{1}\mathbf{e}_{1}^{\top})\mathbf{P}\mathcal{U}^{\rm(N)}$, we know the first entry of $\tilde{\mathbf{u}}:=\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}$ is always 00.
-We further observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$ implies
-$\begin{bmatrix}\tilde{\mathbf{u}}\\ \mathbf{w}\end{bmatrix}=\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ \mathbf{w}\end{bmatrix}\in\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\mathcal{U}_{c}^{\rm(N)}:=\tilde{\mathcal{U}}_{0}.$
-We let $\tilde{\mathbf{u}}_{1}\in\mathbb{R}^{n-1}$ be the restriction of $\tilde{\mathbf{u}}\in\mathbb{R}^{n}$ to the last $n-1$ entries, and let $\tilde{\mathcal{U}}\subset\mathbb{R}^{m+n-1}$ be the restriction of $\tilde{\mathcal{U}}_{0}\subset\mathbb{R}^{m+n}$ to the last $m+n-1$ entries, so that we have $(\tilde{\mathbf{u}}_{1}^{\top},\mathbf{w}^{\top})^{\top}\in\tilde{\mathcal{U}}$.
+By $\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}\in\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathcal{U}^{\mathbb{S}}=\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})\mathbf{P}^{\top}\mathbf{P}\mathcal{U}^{\mathbb{S}}=(\mathbf{I}_{n}-\mathbf{e}_{1}\mathbf{e}_{1}^{\top})\mathbf{P}\mathcal{U}^{\mathbb{S}}$, we know the first entry of $\tilde{\mathbf{u}}:=\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}$ is always 0.
+We further observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}$ implies
+$\begin{bmatrix}\tilde{\mathbf{u}}\\ \mathbf{w}\end{bmatrix}=\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ \mathbf{w}\end{bmatrix}\in\begin{bmatrix}\mathbf{P}(\mathbf{I}_{n}-\mathbf{x}\mathbf{x}^{\top})&0\\ 0&\mathbf{I}_{m}\end{bmatrix}\mathcal{U}_{c}^{\mathbb{S}}:=\tilde{\mathcal{U}}_{0}.$
+We let $\tilde{\mathbf{u}}_{1}\in\mathbb{R}^{n-1}$ be the restriction of $\tilde{\mathbf{u}}\in\mathbb{R}^{n}$ to the last n1n-1 entries, and let $\tilde{\mathcal{U}}\subset\mathbb{R}^{m+n-1}$ be the restriction of $\tilde{\mathcal{U}}_{0}\subset\mathbb{R}^{m+n}$ to the last m+n1m+n-1 entries, so that we have $(\tilde{\mathbf{u}}_{1}^{\top},\mathbf{w}^{\top})^{\top}\in\tilde{\mathcal{U}}$.
 This can be more precisely formulated as
 $\tilde{\mathcal{U}}=\mathbf{R}\tilde{\mathcal{U}}_{0},\qquad\text{where }\mathbf{R}=[0,\mathbf{I}_{m+n-1}]\in\mathbb{R}^{(m+n-1)\times(m+n)}.$
-In light of [49, Exercise 7.5.4] we have $\omega(\tilde{\mathcal{U}})\leq\omega(\tilde{\mathcal{U}}_{0})\leq\omega(\mathcal{U}_{c}^{\rm(N)})$.
-With these preparations, we let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}^{[1]},\tilde{\bm{\Phi}}^{[2:n]}]$ where $\tilde{\bm{\Phi}}^{[1]}$ is the first column of $\tilde{\bm{\Phi}}$ and $\tilde{\bm{\Phi}}^{[2:n]}\in\mathbb{R}^{m\times(n-1)}$ is composed by the last $n-1$ columns of $\tilde{\bm{\Phi}}$, and can proceed as follows:
-$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.20)
-$\qquad=\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}\mathbf{e}_{1}))^{*}\tilde{\bm{\Phi}})\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big{|}$
+In light of [50, Exercise 7.5.4] we have $\omega(\tilde{\mathcal{U}})\leq\omega(\tilde{\mathcal{U}}_{0})\leq\omega(\mathcal{U}_{c}^{\mathbb{S}})$.
+With these preparations, we let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}^{[1]},\tilde{\bm{\Phi}}^{[2:n]}]$ where $\tilde{\bm{\Phi}}^{[1]}$ is the first column of $\tilde{\bm{\Phi}}$ and $\tilde{\bm{\Phi}}^{[2:n]}\in\mathbb{R}^{m\times(n-1)}$ is composed by the last n1n-1 columns of $\tilde{\bm{\Phi}}$, and can proceed as follows:
+$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.20)
+$\qquad=\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}\mathbf{e}_{1}))^{*}\tilde{\bm{\Phi}})\mathbf{P}\mathbf{u}_{\mathbf{x}}^{\bot}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}-\sqrt{\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big|$
 $\qquad\leq\sup_{(\tilde{\mathbf{u}},\mathbf{w})\in\tilde{\mathcal{U}}_{0}}\Big|\Big\|\frac{\Im(\operatorname{diag}(\operatorname{sign}(\tilde{\bm{\Phi}}\mathbf{e}_{1}))^{*}\tilde{\bm{\Phi}})\tilde{\mathbf{u}}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}-\sqrt{\|\tilde{\mathbf{u}}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}}\Big|$
@@
 $\mathcal{N}(0,1)$ entries almost surely.
-Recalling $\omega(\tilde{\mathcal{U}})\leq\omega(\mathcal{U}_{c}^{\rm(N)})$ and observing $\operatorname{rad}(\tilde{\mathcal{U}})\leq\operatorname{rad}(\tilde{\mathcal{U}}_{0})\leq\operatorname{rad}(\mathcal{U}_{c}^{\rm(N)})\leq 1$, a straightforward application of Lemma A.4 yields the desired claim.
+Recalling $\omega(\tilde{\mathcal{U}})\leq\omega(\mathcal{U}_{c}^{\mathbb{S}})$ and observing $\operatorname{rad}(\tilde{\mathcal{U}})\leq\operatorname{rad}(\tilde{\mathcal{U}}_{0})\leq\operatorname{rad}(\mathcal{U}_{c}^{\mathbb{S}})\leq 1$, a straightforward application of Lemma A.4 yields the desired claim.
 ∎
@@
 We let $\mathcal{K}=\{\mathbf{x}\}$ for a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$. Substituting the fixed-$\mathbf{x}$ bound in Lemma A.3 into (A.12)–(A.13) yields
-$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C_{1}[\omega(\mathcal{U}_{1}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2})-4\exp(-c_{2}m).$ (A.22)
+$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C_{1}[\omega(\mathcal{U}_{1}^{\mathbb{S}})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2})-4\exp(-c_{2}m).$ (A.22)
 Similarly, we substitute the fixed-$\mathbf{x}$ bound in Lemma A.5 into (A.19) to obtain
-$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{3}[\omega(\mathcal{U}_{c}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2})-4\exp(-c_{4}m).$ (A.23)
-By (A.5), $m=\Omega(\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\rm(N)}))$, and letting $t=c_{5}\sqrt{m}\delta$ with sufficiently small $c_{5}$ in (A.22)–(A.23), we obtain
-$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\delta\right)\geq 1-11\exp(-c_{6}\delta^{2}m).$
+$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{3}[\omega(\mathcal{U}_{c}^{\mathbb{S}})+t]}{\sqrt{m}}\right)\geq 1-\exp(-t^{2})-4\exp(-c_{4}m).$ (A.23)
+By (A.5), $m=\Omega(\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}}))$, and letting $t=c_{5}\sqrt{m}\delta$ with sufficiently small $c_{5}$ in (A.22)–(A.23), we obtain
+$\mathbbm{P}\left(\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\delta\right)\geq 1-11\exp(-c_{6}\delta^{2}m).$
 Note that this event is the same as $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},\delta)$, so the proof is complete. ∎
@@
 Under the setting of Lemma A.3, let $\mathcal{K}$ be an arbitrary subset of $\mathbb{S}^{n-1}$.
-There exist absolute constants $c_{1},C_{2},C_{3},C_{4}$ such that for any $\eta\in(0,c_{1})$ and $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$, if
-$m\geq\frac{C_{2}}{\log(\eta^{-1})}\left(\frac{\omega^{2}(\mathcal{U}_{1}^{\rm(N)})}{\eta^{2}}+\frac{\mathscr{H}(\mathcal{K},r)}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}}\right),$ (A.24)
-then with probability at least $1-C_{3}\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$, we have
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq C_{4}\eta\sqrt{\log(\eta^{-1})}.$
+There exist absolute constants $c_{1},C_{2},C_{3},c_{4},C_{5}$ such that for any $\eta\in(0,c_{1})$ and $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$, if
+$m\geq\frac{C_{2}}{\log(\eta^{-1})}\left(\frac{\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}})}{\eta^{2}}+\frac{\mathscr{H}(\mathcal{K},r)}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}}\right),$ (A.24)
+then with probability at least $1-C_{3}\exp(-c_{4}\eta^{2}\log(\eta^{-1})m)$, we have
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|\leq C_{5}\eta\sqrt{\log(\eta^{-1})}.$
 Proof.
@@
 Then by Lemma A.3 and a union bound, we obtain
-$\mathbbm{P}\left(\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq\frac{C(\omega(\mathcal{U}_{1}^{\rm(N)})+t)}{\sqrt{m}}\right)\geq 1-2\exp\big{(}\mathscr{H}(\mathcal{K},r)-t^{2}\big{)}$
-for any $t\geq 0$. Therefore, setting $t=\Theta(\sqrt{\mathscr{H}(\mathcal{K},r)}+\omega(\mathcal{U}_{1}^{\rm(N)}))$ yields that the event
-$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq\frac{C_{1}[\omega(\mathcal{U}_{1}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},r)}]}{\sqrt{m}}$ (A.25)
-holds with probability at least $1-2\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$.
+$\mathbbm{P}\left(\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|\leq\frac{C(\omega(\mathcal{U}_{1}^{\mathbb{S}})+t)}{\sqrt{m}}\right)\geq 1-2\exp\big(\mathscr{H}(\mathcal{K},r)-t^{2}\big)$
+for any $t\geq 0$.
+Under the sample complexity $m=\Omega(\frac{1}{\eta^{2}\log(\eta^{-1})}[\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}})+\mathscr{H}(\mathcal{K},r)])$, setting $t=\eta\sqrt{\log(\eta^{-1})}\sqrt{m}$ yields that the event
+$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|\leq C_{1}\eta\sqrt{\log(\eta^{-1})}$ (A.25)
+holds with probability at least $1-2\exp(-\frac{1}{2}\eta^{2}\log(\eta^{-1})m)$.
 Bounding the Number of Small Measurements: For small enough $\eta>0$, we recall $\mathcal{J}_{\mathbf{x},\eta}:=\big\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\big\}.$ We now bound $|\mathcal{J}_{\mathbf{x},\eta}|$ over $\mathbf{x}\in\mathcal{N}_{r}$.
-For a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$, by $\Re(\bm{\Phi}_{i}^{*}\mathbf{x})\sim\mathcal{N}(0,1)$, we have $p_{0}:=\mathbbm{P}\big{(}|\bm{\Phi}_{i}\mathbf{x}|\leq\eta\big{)}\leq\mathbbm{P}\big{(}|\Re(\bm{\Phi}_{i}^{*}\mathbf{x})|\leq\eta\big{)}\leq\sqrt{\frac{2}{\pi}}\eta.$ Note that $|\mathcal{J}_{\mathbf{x},\eta}|\sim\text{Bin}(m,p_{0})$, so the Chernoff bound (e.g., [36, Sec.
+For a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$, by $\Re(\bm{\Phi}_{i}^{*}\mathbf{x})\sim\mathcal{N}(0,1)$, we have $p_{0}:=\mathbbm{P}\big(|\bm{\Phi}_{i}\mathbf{x}|\leq\eta\big)\leq\mathbbm{P}\big(|\Re(\bm{\Phi}_{i}^{*}\mathbf{x})|\leq\eta\big)\leq\sqrt{\frac{2}{\pi}}\eta.$ Note that $|\mathcal{J}_{\mathbf{x},\eta}|\sim\text{Bin}(m,p_{0})$, so the Chernoff bound (e.g., [37, Sec.
 4.1]) gives $\mathbbm{P}\big(|\mathcal{J}_{\mathbf{x},\eta}|\geq\eta m\big)\leq\exp(-c_{1}\eta m)$ for some absolute constant $c_{1}$.
@@
 The remainder of the proof proceeds on the events (A.25) and (A.26).
-Bounding the Approximation Error: We seek to understand the gap between $\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ and $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ For any $\mathbf{x}\in\mathcal{K}$ we let $\mathbf{x}_{r}=\mathrm{arg}\min_{\mathbf{u}\in\mathcal{N}_{r}}\|\mathbf{u}-\mathbf{x}\|_{2}$.
+Bounding the Approximation Error: We seek to understand the gap between $\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ and $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ For any $\mathbf{x}\in\mathcal{K}$ we let $\mathbf{x}_{r}=\mathrm{arg}\min_{\mathbf{u}\in\mathcal{N}_{r}}\|\mathbf{u}-\mathbf{x}\|_{2}$.
 Here, $\mathbf{x}_{r}$ depends on $\mathbf{x}$, but we drop such dependence to avoid cumbersome notation.
@@
 Now by $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ defined in (A.15), we calculate that
-$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}-\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big{|}\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})-f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big{|}$ (A.27)
-$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}\langle\mathbf{x}_{0}-\mathbf{x}_{r},\mathbf{u}\rangle\big{|}$
-$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}+r.$ (A.28)
+$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|-\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big|\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})-f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big|$ (A.27)
+$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|\langle\mathbf{x}_{0}-\mathbf{x}_{r},\mathbf{u}\rangle\big|$
+$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|+r.$ (A.28)
 To bound the first term in (A.28), we first divide it into two terms according to $\mathcal{J}_{\mathbf{x}_{r},\eta}$:
-$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}$
-$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}$
-$+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}:=I_{1}+I_{2}.$ (A.29)
+$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|$
+$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|$
+$+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|:=I_{1}+I_{2}.$ (A.29)
 Bounding $I_{1}$: On the event of (A.26) we have $|\mathcal{J}_{\mathbf{x}_{r},\eta}|<\eta m$.
 Thus, we simply use the universal bound $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|\leq 2$ to get
-$I_{1}$ $\leq\frac{1}{\kappa m}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\big{|}\bm{\Phi}_{i}^{*}\mathbf{u}\big{|}^{2}\right)^{1/2}$
-$\leq\frac{2\eta}{\kappa}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq\eta m\end{subarray}}\left(\frac{1}{\eta m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}\leq C_{4}\eta\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{\eta m}}+\sqrt{\log(\eta^{-1})}\right),$ (A.30)
+$I_{1}$ $\leq\frac{1}{\kappa m}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\big|\bm{\Phi}_{i}^{*}\mathbf{u}\big|^{2}\right)^{1/2}$
+$\leq\frac{2\eta}{\kappa}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq\eta m\end{subarray}}\left(\frac{1}{\eta m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}\leq C_{4}\eta\left(\frac{\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{\eta m}}+\sqrt{\log(\eta^{-1})}\right),$ (A.30)
 where the last inequality follows from Lemma 2.1 and holds with probability at least $1-2\exp(-c_{5}\eta m\log(\eta^{-1}))$.
@@
 This allows us to proceed as follows:
-$I_{2}\leq\frac{2}{\kappa\eta}\frac{1}{m}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\big{|}\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})\big{|}\big{|}\bm{\Phi}_{i}^{*}\mathbf{u}\big{|}$ (A.31)
-$\leq\frac{2}{\kappa\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$ (A.32)
+$I_{2}\leq\frac{2}{\kappa\eta}\frac{1}{m}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\big|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})\big|\big|\bm{\Phi}_{i}^{*}\mathbf{u}\big|$ (A.31)
+$\leq\frac{2}{\kappa\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$ (A.32)
 $\leq\frac{C_{5}}{\eta}\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{v}|^{2}\right)^{1/2}\leq\frac{C_{6}}{\eta}\left(\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{m}}+r\right),$ (A.33)
@@
 Substituting (A.30) and (A.33) into (A.29) we can bound the first term in (A.28) as
-$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\right|\leq C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.34)
+$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\right|$
+$\leq C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.34)
 Completing the Proof: Note that $\mathbf{x}$ in (A.27) can be arbitrary point in $\mathcal{K}$, and our bound (A.34) is uniform for all $\mathbf{x}\in\mathcal{K}$.
 Substituting (A.34) into (A.27)–(A.28) and taking the supremum over $\mathcal{K}$, we obtain
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}$ $\leq\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}+r$
-$+C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right),$
-and then we further apply (A.25), as well as $\eta<1$, to arrive at
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq C_{8}\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{\frac{\mathscr{H}(\mathcal{K},r)}{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.35)
-In summary, this bound holds with probability at least $1-C_{9}\exp(-c_{10}\eta m)-C_{11}\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathcal{H}(\mathcal{K},r))$ under the sample complexity $m=\Omega(\frac{\mathscr{H}(\mathcal{K},r)}{\eta}+\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see such sample complexity and probability term from (A.26), (A.25), (A.30) and (A.29).
-We now set $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$ as in our statement, and the bound in (A.35) reads as
-$O\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{\frac{\mathscr{H}(\mathcal{K},r)}{m}}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}\right).$ (A.36)
-Under the assumed sample complexity (A.24), the bound (A.36) scales as $O(\eta\sqrt{\log(\eta^{-1})})$, and we can promise a probability at least $1-C_{12}\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$ for the claim to hold.
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|$ $\leq\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|+r+C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right),$
+and then we further apply (A.25), as well as $\eta<1$ and $\frac{\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}\leq\eta\sqrt{\log(\eta^{-1})}$ that holds under $m=\Omega(\frac{\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})})$, to arrive at
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|\leq C_{8}\left(\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.35)
+In summary, this bound holds with probability at least $1-C_{9}\exp(-c_{10}\eta^{2}\log(\eta^{-1})m)$ under the sample complexity $m=\Omega(\frac{1}{\eta^{2}\log(\eta^{-1})}[\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}})+\mathscr{H}(\mathcal{K},r)])$; see such sample complexity and probability term from (A.25), (A.26) and (A.30).
+We now set $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$, and using the assumed sample complexity (A.24), then the bound in (A.35) reads as $O(\eta\sqrt{\log(\eta^{-1})})$.
 This completes the proof.
@@
 Lemma A.7.
-In the setting of Lemma A.5, there exist some absolute constants $c_{1},C_{2},C_{3},c_{4},C_{5},c_{6},C_{7}$ such that given any $\eta\in(0,c_{1})$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if
-$m\geq C_{2}\left(\frac{\omega^{2}(\mathcal{U}_{c}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}\right),$ (A.37)
-then with probability at least $1-C_{3}\exp(-c_{4}[\omega^{2}(\mathcal{U}_{c}^{\rm(N)})+\mathscr{H}(\mathcal{K},\eta^{3})])-C_{5}\exp(-c_{6}\eta^{2}m)$, we have
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}\leq C_{7}\eta\sqrt{\log(\eta^{-1})}.$
+In the setting of Lemma A.5, there exist some absolute constants $c_{1},C_{2},C_{3},c_{4},C_{5}$ such that given any $\eta\in(0,c_{1})$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if
+$m\geq C_{2}\left(\frac{\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}\right),$ (A.36)
+then with probability at least $1-C_{3}\exp(-c_{4}\eta^{2}m)$, we have
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big|\leq C_{5}\eta\sqrt{\log(\eta^{-1})}.$
 Proof.
@@
 We apply Lemma A.5 to every $\mathbf{x}\in\mathcal{N}_{r}$, along with a union bound, to obtain that for any $t\geq 0$, the event
-$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{1}(\omega(\mathcal{U}_{c}^{\rm(N)})+t)}{\sqrt{m}}$
-holds with probability at least $1-2\exp(\mathscr{H}(\mathcal{K},r)-t^{2})$. Therefore, setting $t^{2}=\Theta(\omega^{2}(\mathcal{U}_{c}^{\rm(N)})+\mathscr{H}(\mathcal{K},r))$ yields that the event
-$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}\leq\frac{C_{2}(\omega(\mathcal{U}_{c}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},r)})}{\sqrt{m}}$ (A.38)
-holds with probability at least $1-\exp(-\omega^{2}(\mathcal{U}_{c}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$.
+$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{1}(\omega(\mathcal{U}_{c}^{\mathbb{S}})+t)}{\sqrt{m}}$
+holds with probability at least $1-2\exp(\mathscr{H}(\mathcal{K},r)-t^{2})$.
+Therefore, under the sample complexity $m=\Omega(\frac{1}{\eta\log(\eta^{-1})}[\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})+\mathscr{H}(\mathcal{K},r)])$, setting $t=\sqrt{\eta\log(\eta^{-1})m}$ yields that the event
+$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big|\leq C_{2}\sqrt{\eta\log(\eta^{-1})}$ (A.37)
+holds with probability at least $1-2\exp(-\frac{1}{2}\eta\log(\eta^{-1})m)$.
 Bounding the Number of Small Measurements: As shown in the proof of Lemma A.6, under the sample complexity $m\geq\frac{C_{3}\mathscr{H}(\mathcal{K},r)}{\eta}$ we have $\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|<\eta m$ with probability at least $1-\exp(-c_{4}\eta m)$, where $\mathcal{J}_{\mathbf{x},\eta}=\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta\}$.
@@
 For clarity we consider a given $\mathbf{x}\in\mathcal{K}$, but we note that the forthcoming arguments hold uniformly for all $\mathbf{x}\in\mathcal{K}.$ By $f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})$ defined in (A.19) and $\|\mathbf{u}-\mathbf{v}\|_{2}\leq|\|\mathbf{u}\|_{2}-\|\mathbf{v}\|_{2}|$, we can utilize the decomposition
-$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}-\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big{|}$ (A.39)
-$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})-f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big{|}$
-$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}\Big{|}$ (A.40)
-$\quad\quad\quad\quad+\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\begin{bmatrix}\mathbf{u}_{\mathbf{x}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big{\|}_{2}-\Big{\|}\begin{bmatrix}\mathbf{u}_{\mathbf{x}_{r}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big{\|}_{2}\Big{|}.$ (A.41)
+$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big|-\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big|$ (A.38)
+$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})-f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big|$
+$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}-\Big\|\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}\Big|$ (A.39)
+$\quad\quad\quad\quad+\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\begin{bmatrix}\mathbf{u}_{\mathbf{x}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big\|_{2}-\Big\|\begin{bmatrix}\mathbf{u}_{\mathbf{x}_{r}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big\|_{2}\Big|.$ (A.40)
 For any $\mathbf{u}\in\mathbb{S}^{n-1}$, $\mathbf{x}\in\mathcal{K}$ and its associated $\mathbf{x}_{r}\in\mathcal{N}_{r}$, by the triangle inequality we have
@@
 so the term in (A.40) is bounded by $2r$ uniformly for all $\mathbf{x}\in\mathcal{K}$. It remains to bound the term in (A.39).
-By the triangle inequality and the observation that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$ gives $\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}$, the term in (A.40) is bounded by
+By the triangle inequality and the observation that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}$ gives $\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}$, the term in (A.39) is bounded by
 (the term in (A.39)) $\leq\frac{1}{\sqrt{m}}\sup_{\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}}\big\|\Im\big[\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big]\big\|_{2}$
-$=\frac{1}{\sqrt{m}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{\|}\Im\big{[}\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big{]}\big{\|}_{2}.$ (A.42)
+$=\frac{1}{\sqrt{m}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big\|\Im\big[\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big]\big\|_{2}.$ (A.41)
 We divide the $m$ measurements into two parts according to certain index sets.
@@
 and then bound the term in (A.41) as $I_{3}+I_{4}$, where
-$I_{3}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}\Big{[}\Im\Big{(}\big{[}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big{]}\bm{\Phi}_{i}^{*}\mathbf{u}\Big{)}\Big{]}^{2}\right)^{1/2},$
-$I_{4}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i\notin\mathcal{E}_{\mathbf{x}}}\Big{[}\Im\Big{(}\big{[}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big{]}\bm{\Phi}_{i}^{*}\mathbf{u}\Big{)}\Big{]}^{2}\right)^{1/2}.$
+$I_{3}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}\Big[\Im\Big(\big[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big]\bm{\Phi}_{i}^{*}\mathbf{u}\Big)\Big]^{2}\right)^{1/2},$
+$I_{4}=\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i\notin\mathcal{E}_{\mathbf{x}}}\Big[\Im\Big(\big[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}\big]\bm{\Phi}_{i}^{*}\mathbf{u}\Big)\Big]^{2}\right)^{1/2}.$
 Bounding $I_{3}$: The issue for measurements in $\mathcal{E}_{\mathbf{x}}$ is the lack of a good bound on $|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|$.
@@
 Combining with $|\Im([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u})|\leq 2|\bm{\Phi}_{i}^{*}\mathbf{u}|$, we proceed as follows:
-$I_{3}$ $\leq 2\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{(}\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\Big{)}^{1/2}$
-$\leq 2\sqrt{\eta+\beta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq(\eta+\beta)m\end{subarray}}\left(\frac{1}{(\eta+\beta)m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
-$\leq C_{7}\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\frac{e}{\eta+\beta}}\right),$ (A.47)
+$I_{3}$ $\leq 2\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big(\frac{1}{m}\sum_{i\in\mathcal{E}_{\mathbf{x}}}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\Big)^{1/2}$
+$\leq 2\sqrt{\eta+\beta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq(\eta+\beta)m\end{subarray}}\left(\frac{1}{(\eta+\beta)m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
+$\leq C_{7}\left(\frac{\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\big(\frac{e}{\eta+\beta}\big)}\right),$ (A.46)
 where (A.46) follows from Lemma 2.1 and holds with probability at least $1-4\exp(-c_{8}(\eta+\beta)m\log(\frac{e}{\eta+\beta}))$.
@@
 Therefore, by $|\Im([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u})|\leq\frac{2\eta^{\prime}}{\eta}|\bm{\Phi}_{i}^{*}\mathbf{u}|$ we can bound $I_{4}$ as
-$I_{4}\leq\frac{2\eta^{\prime}}{\eta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}\leq\frac{C_{9}\eta^{\prime}}{\eta},$ (A.49)
-where the second inequality holds with probability at least $1-4\exp(-c_{10}m)$ if $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see (2.10).
+$I_{4}\leq\frac{2\eta^{\prime}}{\eta}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\sqrt{m}}\leq\frac{C_{9}\eta^{\prime}}{\eta},$ (A.48)
+where the second inequality holds with probability at least $1-4\exp(-c_{10}m)$ if $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}}))$; see (2.10).
 Combining (A.46) and (A.48) and recalling (A.41), we arrive at
@@
 Substituting them into (A.38)–(A.40), along with a supremum over $\mathbf{x}\in\mathcal{K}$, yields
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big{|}$
-$\qquad+C_{11}\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\Big{(}\frac{e}{\eta+\beta}\Big{)}}+\frac{\eta^{\prime}}{\eta}\right)+2r.$
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big|$
+$\qquad+C_{11}\left(\frac{\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}+\sqrt{(\eta+\beta)\log\Big(\frac{e}{\eta+\beta}\Big)}+\frac{\eta^{\prime}}{\eta}\right)+2r.$
 Combining with the bound in (A.37), taking $\beta=\Theta(\eta)$, and also summarizing the sample complexity and probability terms, we arrive at the following conclusion: Suppose
-$m\geq C_{12}\left[\omega^{2}(\mathcal{U}_{1}^{\rm(N)})+\frac{\mathscr{H}(\mathcal{K},r)}{\eta}\right]\qquad\text{with large enough }C_{12},$ (A.51)
+$m\geq C_{12}\left[\frac{\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})}{\eta\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},r)}{\eta}\right]\quad\text{with large enough }C_{12},$ (A.50)
 and
-$\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\leq c_{13}\eta^{\prime}\qquad\text{with small enough }c_{13},$ (A.52)
+$\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\leq c_{13}\eta^{\prime}\quad\text{with small enough }c_{13},$ (A.51)
 the event
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{14}\left(\frac{\omega(\mathcal{U}^{\rm(N)}_{c})+\sqrt{\mathscr{H}(\mathcal{K},r)}}{\sqrt{m}}+\sqrt{\eta\log(\eta^{-1})}+\frac{\eta^{\prime}}{\eta}+r\right)$
-holds with probability at least $1-C_{15}\exp(-c_{16}(\omega^{2}(\mathcal{U}^{\rm(N)}_{c})+\mathscr{H}(\mathcal{K},r)))-C_{17}\exp(-c_{18}\eta m)$.
-We mention that the condition (A.51) is needed for ensuring $\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|<\eta m$ and the second inequality of (A.49), and the condition (A.52) is needed in (A.44).
-Further Simplification: We now take the tightest choice for $\eta^{\prime}$ that satisfies the required (A.52), namely $\eta^{\prime}=\Theta\big{(}\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\big{)}.$ We further set $\eta=\hat{\eta}^{2}$ and $r=\eta^{3/2}=\hat{\eta}^{3}$, and then the above statement simplifies to the following: if
-$m\geq C_{19}\left(\omega^{2}(\mathcal{U}^{\rm(N)}_{c})+\frac{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}{\hat{\eta}^{2}}\right),$
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{14}\left(\sqrt{\eta\log(\eta^{-1})}+\frac{\eta^{\prime}}{\eta}+r\right)$
+holds with probability at least $1-C_{15}\exp(-c_{16}\eta m)$.
+We mention that the condition (A.50) is needed for ensuring (A.37), $\sup_{\mathbf{x}\in\mathcal{N}_{r}}|\mathcal{J}_{\mathbf{x},\eta}|<\eta m$ and the second inequality of (A.48), and the condition (A.51) is needed in (A.43).
+Further Simplification: We now take the tightest choice for $\eta^{\prime}$ that satisfies the required (A.51), namely
+$\eta^{\prime}=\Theta\Big(\frac{\omega(\mathcal{K}_{(r)})}{\sqrt{\eta m}}+r\sqrt{\log(\eta^{-1})}\Big).$
+We further set $\eta=\hat{\eta}^{2}$ and $r=\eta^{3/2}=\hat{\eta}^{3}$ for some $\hat{\eta}>0$, and then the above statement simplifies to the following: if
+$m\geq C_{17}\left(\frac{\omega^{2}(\mathcal{U}^{\mathbb{S}}_{c})}{\hat{\eta}^{2}\log(\hat{\eta}^{-1})}+\frac{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}{\hat{\eta}^{2}}\right),$
 then the event
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{20}\left(\frac{\omega(\mathcal{U}_{c}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}+\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}+\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})}\right)$ (A.53)
-holds with probability at least $1-C_{21}\exp(-c_{22}(\omega^{2}(\mathcal{U}_{c}^{\rm(N)})+\mathscr{H}(\mathcal{K},\hat{\eta}^{3})))-C_{23}\exp(-c_{24}\hat{\eta}^{2}m)$.
-Under the sample complexity in (A.37) but with $\eta$ replaced by $\hat{\eta}$, it is easy to see
-$\frac{\omega(\mathcal{U}_{c}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}+\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}=O(\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})}).$
-Further renaming $\hat{\eta}$ to $\eta$ completes the proof. ∎
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u})|\leq C_{18}\left(\frac{\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}+\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})}\right)$ (A.52)
+holds with probability at least $1-C_{19}\exp(-c_{20}\hat{\eta}^{2}m)$. Under the sample complexity
+$m=\Omega\left(\frac{\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})}{\hat{\eta}^{2}\log(\hat{\eta}^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\hat{\eta}^{3})})}{\hat{\eta}^{8}\log(\hat{\eta}^{-1})}+\frac{\mathscr{H}(\mathcal{K},\hat{\eta}^{3})}{\hat{\eta}^{2}}\right),$
+which is identical to (A.36) except for except for the notation $\eta$ versus $\hat{\eta}$, we have $\frac{\hat{\eta}^{-3}\omega(\mathcal{K}_{(\hat{\eta}^{3})})}{\sqrt{m}}=O(\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})})$ and hence the right-hand side of (A.52) is bounded by $O(\hat{\eta}\sqrt{\log(\hat{\eta}^{-1})})$.
+Further renaming $\hat{\eta}$ to $\eta$ completes the proof.
+∎
 A.7 Proof of Lemma A.1
@@
 We are ready to substitute the bounds for the parallel term and the orthogonal term into (A.5) to establish Lemma A.1. Recall from (A.12) and (A.13) that
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$
 holds with probability $1-4\exp(-c_{2}m)$, and that
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|{f}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq C_{3}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|{f}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq C_{3}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$
 holds with probability at least $1-4\exp(-c_{4}m)$ due to (A.19).
 We now observe that the stated sample complexity (A.1) implies (A.24) and (A.36), and hence we can apply Lemma A.6 and Lemma A.7 to obtain
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|{f}_{1}^{\|}(\mathbf{x},\mathbf{u})|=O\big{(}\eta\sqrt{\log(\eta^{-1})}\big{)},$
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big{(}\eta\sqrt{\log(\eta^{-1})}\big{)},$
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|{f}_{1}^{\|}(\mathbf{x},\mathbf{u})|=O\big(\eta\sqrt{\log(\eta^{-1})}\big),$
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big(\eta\sqrt{\log(\eta^{-1})}\big),$
 that hold with the promised probability. Therefore, we arrive at
-$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big{(}\eta\sqrt{\log(\eta^{-1})}\big{)}.$
-In view of (A.3) and (A.5), we derive $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}f(\mathbf{x},\mathbf{u},\mathbf{w})=O(\eta\sqrt{\log(\eta^{-1})})$, which is just the desired RIP with distortion $\delta_{\eta}$.
+$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u},\mathbf{w})|=O\big(\eta\sqrt{\log(\eta^{-1})}\big).$
+In view of (A.3) and (A.5), we derive $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}f(\mathbf{x},\mathbf{u},\mathbf{w})=O(\eta\sqrt{\log(\eta^{-1})})$, which is just the desired RIP with distortion $\delta_{\eta}$.
 The proof is complete.
@@
 $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{1}{\sqrt{m}}\cdot\frac{1}{\sqrt{\zeta_{0}m}}\sum_{i\in I_{\zeta_{0}}^{\prime}}\big|(\bm{\Phi}_{i}^{\Re})^{\top}\mathbf{x}\big|.$
-Now let us construct a set $\mathcal{V}:=\big{\{}\frac{1}{\sqrt{\zeta_{0}m}},-\frac{1}{\sqrt{\zeta_{0}m}},0\big{\}}^{m}\cap\Sigma^{m}_{\zeta_{0}m}$ whose elements are $m$-dimensional, $\zeta_{0}m$-sparse {±$\{\frac{\pm 1}{\sqrt{\zeta_{0}m}},0\}$-valued vectors.
+Now let us construct a set $\mathcal{V}:=\big\{\frac{1}{\sqrt{\zeta_{0}m}},-\frac{1}{\sqrt{\zeta_{0}m}},0\big\}^{m}\cap\Sigma^{m}_{\zeta_{0}m}$ whose elements are $m$-dimensional, $\zeta_{0}m$-sparse $\{\frac{\pm 1}{\sqrt{\zeta_{0}m}},0\}$-valued vectors.
 Combining with the definition of $I_{\zeta_{0}}^{\prime}$ we can write
@@
 where we observe that $\bm{\Phi}^{\Re}\mathbf{x}\sim\mathcal{N}(0,\mathbf{I}_{m})$ for a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and further let $\mathbf{g}\sim\mathcal{N}(0,\mathbf{I}_{n})$ and assert that $\mathbf{v}^{\top}\bm{\Phi}^{\Re}\mathbf{x}$ and $\mathbf{g}^{\top}\mathbf{v}$ have the same distribution.
-Therefore, by Gaussian concentration (e.g., [49, Thm.
+Therefore, by Gaussian concentration (e.g., [50, Thm.
 5.2.2]) we can show that $\max_{\mathbf{v}\in\mathcal{V}}\mathbf{v}^{\top}\bm{\Phi}^{\Re}\mathbf{x}\geq\frac{1}{2}\omega(\mathcal{V})$ with probability at least $1-2\exp(-c_{1}\omega^{2}(\mathcal{V}))$.
 We now seek lower bound on $\omega(\mathcal{V})$.
-By the Sparse Varshamov-Gilbert construction (e.g., [45, Lem.
+By the Sparse Varshamov-Gilbert construction (e.g., [46, Lem.
 4.14]) there exist $(1+\frac{1}{2\zeta_{0}})^{\frac{\zeta_{0}m}{8}}$ distinct points contained in $\mathcal{V}$ with mutual $\ell_{2}$ distances greater than $\frac{1}{\sqrt{2}}$.
@@
 Recall that all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{4\delta_{1}}{5}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2})$ satisfy the constraint $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}-\mathbf{e}_{1}\|_{2}\leq\varepsilon$.
-Since $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{c_{3}}{2}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$, all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$ satisfy the constraint of (1.8).
+Since
+$\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\geq\frac{c_{3}}{2}\sqrt{\zeta_{0}\log(e/\zeta_{0})},$
+all points in $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$ satisfy the constraint of (1.8).
 To conclude the proof, it remains to show $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\geq\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$. To do so, we proceed under the assumption
@@
 $\hat{\mathbf{x}}_{c}=\mathrm{arg}\min~\|\mathbf{u}\|_{1},\qquad\text{subject to }\mathbf{u}\in\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}\Big(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})}\Big).$
-Under small enough $\zeta_{0}$, we use $\|\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}$ to obtain $\frac{2c_{3}\delta_{1}\sqrt{\zeta_{0}\log(e/\zeta_{0})}}{5}<\frac{1}{2}\leq\|\mathbf{x}^{\star}\|_{2}$.
+Under small enough $\zeta_{0}$, we use $\|\mathbf{x}^{\star}\|_{2}\geq\frac{1}{2}$ to obtain
+$\frac{2c_{3}\delta_{1}\sqrt{\zeta_{0}\log(e/\zeta_{0})}}{5}<\frac{1}{2}\leq\|\mathbf{x}^{\star}\|_{2}.$
 Then, it is not hard to observe that $\hat{\mathbf{x}}_{c}$ must live in the boundary of $\mathbb{B}_{2}^{n}(\mathbf{x}^{\star},\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})})$.
-Hence, with the promised probability, we have $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}=\|\hat{\mathbf{x}}_{c}-\mathbf{x}^{\star}\|_{2}=\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})}$.
-The result follows.
-∎
+Hence, with the promised probability, we have
+$\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}=\|\hat{\mathbf{x}}_{c}-\mathbf{x}^{\star}\|_{2}=\frac{2c_{3}\delta_{1}}{5}\sqrt{\zeta_{0}\log(e/\zeta_{0})}.$
+The result follows. ∎
 B.2 Proof of Lemma 4.2
@@
 then with probability at least $1-3\exp(-c_{3}\eta m)$ we have $\sup_{\mathbf{x}\in\mathcal{K}}~|\mathcal{J}_{\mathbf{x},\eta}|\leq\eta m.$
-Before proving this, we note that it immediately leads to Lemma 4.1: by $\omega^{2}(\mathcal{K}_{(r)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})\leq 4\omega^{2}(\mathcal{K})$ [49, Sec.
+Before proving this, we note that it immediately leads to Lemma 4.2: by $\omega^{2}(\mathcal{K}_{(r)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})\leq 4\omega^{2}(\mathcal{K})$ [50, Sec.
 7.5.1] and $\mathscr{H}(\mathcal{K},r)\leq C_{1}\frac{\omega^{2}(\mathcal{K})}{r^{2}}=\Theta(\frac{\log(\eta^{-1})\omega^{2}(\mathcal{K})}{\eta^{2}})$ from (2.4), we find that $m=\Omega(\eta^{-3}\log(\eta^{-1})\omega^{2}(\mathcal{K}))$ in Lemma 4.2 suffices to imply the sample complexity in Lemma B.1.